### v1

# Is Quadratic Gravity ghost free because the ghost is only virtual?

Abstract
In [1], inspired by the works [7]-[30], a prescription for calculating the correlation functions in Quadratic Gravity [31]-[32] was presented and further exploited in [2]-[5]. This procedure does not enter in conflict with unitarity. The corresponding Hamiltonian quantization is based on a covariant and contra-variant distinction in the non positive definite metric in the space of states [13]. The Gauss-Ostrogradsky method for higher order theories defines two momentum densities $P_{1}$ and $P_{2}$ and two coordinate densities $Q_{1}$ and $Q_{2}$, one pair is standard, the other ghost like. The approach in [1] involves the continuation $P_{2}\to iP_{2}$ and $Q_{2}\to iQ_{2}$ of the ghost variables acting on kets $|>$ after taking mean values. In the present work, following [42], the LSZ rules are derived, with a formalism adapted to full quartic or higher order theories. The derivation of these rules is possible only after the second quantization of the free theory. The hypothesis for quantization are that $Q_{1},Q_{2}]=\text{"gauge terms"}$, $P_{1},P_{2}]=\text{"gauge terms"}$ and $P_{1},Q_{1}]=iI+\text{"gauge terms"}$. This alone leads to the conclusion that $P_{2},Q_{2}]=-iI+\text{"gauge terms"}$, thus this last pair of variables is ghost like. The graviton contains a massless mode, which is the standard graviton, plus two massive modes with masses $m_{2}$ and $m_{3}$. The third mode is usually interpreted as a ghost in the literature [32]. Here it is shown that, even after making the continuation $P_{2}\to iP_{2}$ and $Q_{2}\to iQ_{2}$, the creation and annihilation operators for this mode commute, thus this particle becomes classical, therefore are not allowed as external lines in a quantum context due to the uncertainty principle. More precisely, the third mode does not appear as a free wave, although it may affect spherically symmetric solutions [32]-[35]. This does not invalidate the model. The effective action $\Gamma$ can be calculated following [31], and can be constrained by the Slavnov-Taylor identities [37]-[38]. The surprising outcome is that the LSZ rules do not contain the third mode as external state, although the correlation functions are conditioned by this mode since it depends on its mass scale. The second mode however, becomes a ghost and is taken care with the methods of [1].
1. Introduction
Stelle’s gravity [31]-[32] is a renormalizable gravity model, which contains a massless mode and two massive ones. The renormalization program for this model was presented by using the Slavnov-Taylor quantization method [37]-[38]. However, one of the massive modes contributes to the kinetic energy with a wrong sign, and was interpreted as a ghost which spoils unitarity. The theory is of higher order, its equations of motion are of fourth order and its Cauchy problem is well posed [40]-[41]. The unitarity problem was already pointed out as the main issue which obstructs it as a quantum gravity model, at least when the quantization is done around a flat background.
On the other hand, starting from the seminal works of Dirac and Pauli [6]-[7], several authors [8]-[30] and more recently [36] started different programs for quantization of higher order theories and avoiding ghosts. These works are not only related to Quantum Gravity, instead they present different motivations, some of them [8]-[9] are even related to Statistical Physics. All of them bring interesting features which are not evident at first sight. An interesting approach was due to Boulware and Gross [13], who made a formal description of the non positive defined metric in the Hilbert space. These authors consider states such that their inner product $n|m>=\eta_{mn}$ is not positive definite. This reference describes a covariant $\eta_{nm}$, contra-variant $\eta^{nm}$ and mixed metric $\delta_{n}^{m}$. The mixed one is positive definite and is the one to be employed for calculating probabilities. The Hilbert space itself has covariant and contra variant components. The mean values have to be adapted to the covariant or contra variant Hilbert space, but usually negative and non bounded from below quantities become bounded by this covariant and contra-variant distinction. These authors also present several useful formulas for the path integral for these theories, together with some interesting insights. In particular, that depending on the use of the covariant, contra-variant or mixed version, a self adjoint observable may be presented in terms of non hermitian matrices. Despite these interesting findings, those authors open the possibility that the path integral they define may not exist. The main point is the possible presence of real and unbounded exponentials in the path integral, which make the full result divergent.
On the other hand, the authors [4]-[29] made several efforts to physically interpret the results of these references. In particular, in [1] the author made the remarkable observation that the quantization scheme of Boulware-Gross applied to the Stelle model leads, in the euclidean setting, to a perfectly defined path integral. This was further exploited in [2]-[5]. The potentially dangerous exponentials become imaginary and the full result is well defined (in a path integral sense). The main observation is that, in a Hamiltonian setting, there are variables which behave normally under time reversal, and there are ones that not. It is not possible to state in brief the results of these detailed works, but it is worthy to mention that the quantization scheme with these variables, denoted as $P_{2}$ and $Q_{2}$ is consistent after making continuation to imaginary values $P_{2}\to iP_{2}$ and $Q_{2}\to iQ_{2}$. Heuristically, if the commutator $P_{2},Q_{2}]$ has a wrong sign, this continuation removes the unwanted minus sign.
Furthermore, both the results of [13] and [1], strongly point out that this quantization does not enter in any conflict with unitarity. As the final result is the euclidean version of Stelle gravity, one may calculate the physical quantities in the euclidean setting, and then go to the Lorenzian one by continuation to real times. The results will be renormalizable, and is possible that unitarity will be preserved.
The results of all the references mentioned above are of significant interest. However, a full presentation requires an analysis of the LSZ formulas which, in this context, have to be taken with care. The point is that, in this quantization method, the creation and annihilation algebra may have some particularities not present in an standard quantum theory. A program about this topic was initiated in [42]. The LSZ rules require a second quantization of the free model consistent with the principles employed in all these references. In the present work, the idea is to present an LSZ deduction fully adapted to a fourth order theory, regardless how the modes are decomposed. The outcome is very interesting. The usually mode denominated as ghost is a ghost, in the sense that is absent as an external line. This conclusion follows using mild assumptions, which can be relaxed. If they are relaxed a bit, this particle emerges in the physical spectrum, but as a decent state with good properties. In any case, all these features will be discussed along the text.
It is important to remark that the bulk of this text presents a quantization scheme which the author believes to be equivalent to the one in [1]-[2]. A full proof will be desirable. The advantage is its compact form, which avoids the above mentioned covariant and contra-variant distinction. However, the main result presented here is related to the LSZ rules, the quantization program is a result that has to be attributed to all the references.
Furthermore, there is nothing wrong with the covariant and contra-variant formalism of [13]. The presentation here is not a critic against this formalism, it just a way for presenting the results in more concise manner. At the end, the references [1]-[5], [13] and [28]-[29] contain the full formalism, and can be consulted if some clarifications are needed.
The present work is organized as follows. Section 2 presents the arguments showing that the third massive graviton mode in Stelle gravity is either confined or purely virtual at classical and quantum level. The Gupta-Bleuler quantization is taken into account. Section 3 reviews the Gauss-Ostrogradsky quantization for the Pais-Uhlenbeck oscillator, and how the ghost emerges. A simple prescription for avoiding the effect of these ghosts is presented. I believe that this prescription is equivalent to the one presented in [1], and is heavily inspired in that reference and also by [13]. The presentation is intended to be more compact, and this has a price of containing less details. In Section 4 the LSZ rules for the Pais-Uhlenbeck oscillator are derived, in a formalism fully adapted for higher order theories. Section 5 reviews the notable result of reference [1] about the finiteness of the path integral when the above described quantization is presented, which presents a quantization of the model which does not allow divergences due to exponential integrations in the path integral and, at the end, results in the euclidean version of Stelle gravity. In Section 6 the LSZ formulas adapted to this fourth order model is derived, the universal form of the effective action is specified and the effect of the non standard creation annihilation operators is clarified in the last part of this section. Section 7 contains possible research programs related to these results.
2. The classical and quantum disappearance of the third massive graviton mode
2.1 The free Stelle action in the De Donder gauge
Consider the Stelle action [31]-[32], which will be taken as
$S_{s}=\int\sqrt{-g}[-\frac{\gamma}{\kappa^{2}}R+\beta R^{2}-\alpha R_{\mu\nu}R^{\mu\nu}]d^{4}x,$
Here $\gamma=2$ and $\kappa^{2}=32\pi G_{N}$. The last action may be linearized in $h^{\mu\nu}=g^{\mu\nu}-\eta^{\mu\nu}$ by taking into account that $\sqrt{g}=1+\frac{h}{2}+O(h^{2})$ and the curvature expansions
$R_{\mu\nu}=\frac{1}{2}[\partial^{\rho}\partial_{\mu}h_{\nu\rho}+\partial^{\rho}\partial_{\nu}h_{\mu\rho}-\square h_{\mu\nu}-\partial_{\mu}\partial_{\nu}h],\qquad R=\partial_{\mu}\partial_{\nu}h^{\mu\nu}-\square h+O(h^{2}),$
$\beta R^{2}=\beta h\square^{2}h+\beta h^{\mu\nu}\partial_{\mu}\partial_{\nu}\partial_{\alpha}\partial_{\beta}h^{\alpha\beta}-2\beta h\square\partial_{\mu}\partial_{\nu}h^{\mu\nu}+\text{total derivatives},$
$\alpha R_{\mu\nu}R^{\mu\nu}=-\frac{\alpha}{4}h\square^{2}h-\frac{\alpha}{4}h_{\mu\nu}\square^{2}h^{\mu\nu}+\frac{\alpha}{2}h_{\mu}^{\sigma}\square\partial_{\sigma}\partial_{\nu}h^{\nu\mu}$
$\frac{\alpha}{2}h\square\partial_{\mu}\partial_{\nu}h^{\nu\mu}-\frac{\alpha}{2}h^{\mu\nu}\partial_{\mu}\partial_{\nu}\partial_{\alpha}\partial_{\beta}h^{\alpha\beta}.$
$\text{total derivatives}$
$\beta R^{2}-\alpha R_{\mu\nu}R^{\mu\nu}=-\frac{\alpha}{4}h^{\mu\nu}\square[\square h_{\mu\nu}-2\partial_{\sigma}\partial^{\nu}h_{\nu\mu}]+(\beta-\frac{\alpha}{4})h\square[\square h-2\partial_{\mu}\partial_{\nu}h^{\nu\mu}]$
$(\beta-\frac{\alpha}{2})h^{\mu\nu}\partial_{\mu}\partial_{\nu}\partial_{\alpha}\partial_{\beta}h^{\alpha\beta}+\text{total derivatives}.$
At this order, the Fierz-Pauli action corresponding to GR is
$\frac{1}{16\pi G}\int\sqrt{g}Rd^{4}x\simeq-\frac{1}{\kappa^{2}}\int h^{\mu\nu}[\square h_{\mu\nu}-2\partial^{\rho}\partial_{(\mu}h_{\nu)\rho}]+\frac{h}{\kappa^{2}}[\square h-2\partial_{\mu}\partial_{\nu}h^{\mu\nu}]d^{3}x,$
this result follows only after expanding the curvature $R$ up to second order in $h^{\mu\nu}$.
The free Stelle action corresponds to the terms quadratic in $h^{\mu\nu}$. With these formulas, it is straightforward to find it, which will be the generalization of the Fierz-Pauli lagrangian adapted to the present context. The result is
$S_{s}=-\frac{2}{\kappa^{2}}\int\bigg[\frac{1}{2}h^{\mu\nu}\bigg(1+\frac{\kappa^{2}\alpha}{4}\Box\bigg)[\square h_{\mu\nu}-2\partial^{\gamma}\partial_{(\nu}h_{\mu)\gamma}]$
$\frac{1}{2}h\bigg(1+\frac{\kappa^{2}(4\beta-\alpha)}{4}\square\bigg)[2\partial_{\alpha}\partial_{\beta}h^{\alpha\beta}-\square h]-\frac{\kappa^{2}}{2}(\beta-\frac{\alpha}{2})h^{\mu\nu}\partial_{\mu}\partial_{\nu}\partial_{\alpha}\partial_{\beta}h^{\alpha\beta}\bigg]d^{4}x.$ (2.1)
If the parameters $\alpha$ and $\beta$ vanish, this reduces to the well known Fierz-Pauli action. The above deduction requires to take into account the $R$ terms up to second order and the remaining ones to lower order, since they will be quadratic in $h_{\mu\nu}$ automatically. The non gauged Stelle action may be expressed in an alternative form
$S_{s}=-\frac{2}{\kappa^{2}}\int\bigg[\frac{1}{2}h^{\mu\nu}\bigg(1+\frac{\kappa^{2}\alpha}{4}\Box\bigg)[\square h_{\mu\nu}-2\partial^{\gamma}\partial_{(\nu}h_{\mu)\gamma}]$
$\frac{1}{2}h\bigg[\frac{2}{3}\bigg(1+\frac{\kappa^{2}(3\beta-\alpha)}{2}\square\bigg)+\frac{1}{3}\bigg(1+\frac{\kappa^{2}\alpha}{4}\square\bigg)\bigg][2\partial_{\alpha}\partial_{\beta}h^{\alpha\beta}-\square h]$
$h^{\mu\nu}\bigg[\bigg(1+\frac{\kappa^{2}\alpha}{4}\partial_{\mu}\partial_{\nu}\bigg)-\bigg(1+\frac{\kappa^{2}(3\beta-\alpha)}{2}\partial_{\mu}\partial_{\nu}\bigg)\bigg]\partial_{\alpha}\partial_{\beta}h^{\alpha\beta}\bigg]d^{4}x.$ (2.2)
The reason for writing all in this form is that it is expressed in terms of the mass scales that will be found below in (2.25) namely
$m_{1}^{2}=0,\qquad m_{2}^{2}=\frac{4}{\kappa^{2}\alpha},\qquad m_{3}^{2}=\frac{2}{\kappa^{2}(3\beta-\alpha)}.$
However, one simply may consider this as a valid way to write the action. In any case, the equations of motion without fixing the gauge are
$\bigg(1+\frac{\kappa^{2}\alpha}{4}\Box\bigg)\square h_{\mu\nu}-\eta_{\mu\nu}\bigg[\frac{2}{3}\bigg(1+\frac{\kappa^{2}(3\beta-\alpha)}{2}\square\bigg)+\frac{1}{3}\bigg(1+\frac{\kappa^{2}\alpha}{4}\square\bigg)\bigg]\square h$
$2\bigg(1+\frac{\kappa^{2}\alpha}{4}\Box\bigg)\partial^{\gamma}\partial_{(\nu}h_{\mu)\gamma}+\eta_{\mu\nu}\bigg[\frac{2}{3}\bigg(1-\frac{\kappa^{2}(3\beta-\alpha)}{2}\square\bigg)+\frac{1}{3}\bigg(1-\frac{\kappa^{2}\alpha}{4}\square\bigg)\bigg]\partial_{\alpha}\partial_{\beta}h^{\alpha\beta}$
$\bigg[\frac{2}{3}\bigg(1+\frac{\kappa^{2}(3\beta-\alpha)}{2}\square\bigg)+\frac{1}{3}\bigg(1+\frac{\kappa^{2}\alpha}{4}\square\bigg)\bigg]\partial_{\mu}\partial_{\nu}h$
$\kappa^{2}(\beta-\frac{\alpha}{2})\partial_{\mu}\partial_{\nu}\partial_{\alpha}\partial_{\beta}h^{\alpha\beta}=0.$ (2.3)
The classical graviton field can be Fourier expanded as in (2.12). Now comes the problem of gauge fixing the model. A customary gauge employed in cosmology is the De Donder gauge
$\partial_{\mu}h^{\mu\nu}=\frac{1}{2}\eta^{\mu\nu}\partial_{\mu}h.$ (2.4)
This gauge will be considered below. But for an easy visualization of results, consider first the Lorenz gauge $\partial_{\mu}h^{\mu\nu}=0$ applied to the equations of motions. The De Donder gauge will be applied later on. The equations in the Lorenz gauge are
$\bigg(1+\frac{\kappa^{2}\alpha}{4}\Box\bigg)\square h_{\mu\nu}-\eta_{\mu\nu}\bigg[\frac{2}{3}\bigg(1+\frac{\kappa^{2}(3\beta-\alpha)}{2}\square\bigg)+\frac{1}{3}\bigg(1+\frac{\kappa^{2}\alpha}{4}\square\bigg)\bigg]\square h$
$\bigg[\frac{2}{3}\bigg(1+\frac{\kappa^{2}(3\beta-\alpha)}{2}\square\bigg)+\frac{1}{3}\bigg(1+\frac{\kappa^{2}\alpha}{4}\square\bigg)\bigg]\partial_{\mu}\partial_{\nu}h=0.$
Contracting the last differential equation with $\eta^{\mu\nu}$ gives
$\bigg(1+\frac{\kappa^{2}(3\beta-\alpha)}{2}\square\bigg)\square h=0.$
Therefore the trace of the metric $h$ is composed by masses $m^{2}_{1}=0$ and $m_{3}^{2}$. The mode with mass $m_{2}^{2}$ does not give contribution $h_{2}$ to the trace. The equations of motion reduce to
$\bigg(1+\frac{\kappa^{2}\alpha}{4}\Box\bigg)\square h_{\mu\nu}-\frac{1}{3}\eta_{\mu\nu}\bigg(1+\frac{\kappa^{2}\alpha}{4}\square\bigg)\square h$
$\bigg[\frac{2}{3}\bigg(1+\frac{\kappa^{2}(3\beta-\alpha)}{2}\square\bigg)+\frac{1}{3}\bigg(1+\frac{\kappa^{2}\alpha}{4}\square\bigg)\bigg]\partial_{\mu}\partial_{\nu}h=0.$ (2.5)
Decompose the trace $h$ into its components $h_{1}$ and $h_{3}$. The last equation is then
$\bigg(1-\frac{\kappa^{2}\alpha}{4}m_{3}^{2}\bigg)m_{3}^{2}h^{3}_{\mu\nu}+\frac{1}{3}\eta_{\mu\nu}m_{3}^{2}\bigg(1-\frac{\kappa^{2}\alpha}{4}m_{3}^{2}\bigg)h_{3}$
$\partial_{\mu}\partial_{\nu}h_{1}+\frac{1}{3}\bigg(1-\frac{\kappa^{2}\alpha}{4}m_{3}^{2}\bigg)\partial_{\mu}\partial_{\nu}h_{3}=0.$ (2.6)
The task is now to constraint the polarizations in (2.12). For this, note that the Lorenz gauge condition implies that
$k^{i}_{\mu}\epsilon^{i\mu\nu}=0,$
where the index $i$ is indicating that the four momenta for every mode are different, since they satisfy a different mass shell relation. All three polarizations $\epsilon^{i}_{\mu\nu}$ with $i=1,2,3$ satisfy the last condition. Also, the fact that $h_{2}$ is zero is equivalent to
$\epsilon_{2}^{p\mu\nu}\eta_{\mu\nu}=0.$
The remaining condition follows from (2.6) and is
$\frac{1}{\omega_{3}(k)}\bigg(1-\frac{\kappa^{2}\alpha}{4}m_{3}^{2}\bigg)m_{3}^{2}\epsilon^{3}_{\mu\nu}+\frac{1}{3\omega_{3}(k)}\eta_{\mu\nu}m_{3}^{2}\bigg(1-\frac{\kappa^{2}\alpha}{4}m_{3}^{2}\bigg)\epsilon_{3}$
$\frac{1}{3\omega_{3}(k)}\bigg(1-\frac{\kappa^{2}\alpha}{4}m_{3}^{2}\bigg)k^{3}_{\mu}k^{3}_{\nu}\epsilon_{3}=0,$
$k_{\mu}k_{\nu}\epsilon_{1}=0.$
The second one follows since $h_{1}$ and $h_{3}$ have different Fourier expansions and have to be considered as independent. But this implies that $\epsilon_{1}=0$. Contracting the first with $k^{3\mu}$ and taking into account that $\epsilon^{3}_{\mu\nu}k^{3\nu}=0$ shows that
$\frac{2}{3\omega_{3}(k)}k_{\mu}m_{3}^{2}\bigg(1-\frac{\kappa^{2}\alpha}{4}m_{3}^{2}\bigg)\epsilon_{3}=0,$ (2.7)
thus $\epsilon_{3}=0$ as well. Therefore $h=0$ and the polarization $\epsilon^{i}_{\mu\nu}$ is orthogonal to $k_{i}^{\mu}$ (this is a bit loosely speaking, as the definition of orthogonality in four dimensions is subtle). If the direction of propagation is $\hat{z}$ it is known that the polarizations for the massless mode are
$\epsilon^{+}_{1\mu\nu}=\left(\begin{array}[]{cccc}0&0&0&0\\ 0&1&0&0\\ 0&0&-1&0\\ 0&0&0&0\end{array}\right),\qquad\epsilon^{x}_{1\mu\nu}=\left(\begin{array}[]{cccc}0&0&0&0\\ 0&0&1&0\\ 0&1&0&0\\ 0&0&0&0\end{array}\right)$ (2.8)
For the massive mode there more possible polarizations. A massive graviton with mass $m$ is not relativistic and in the rest frame its 4-momentum is $k^{\mu}=(m,0,0,0)$. A linearly independent set satisfying the transversality and trace zero conditions are
$\epsilon^{+}_{\mu\nu}=\left(\begin{array}[]{cccc}0&0&0&0\\ 0&1&0&0\\ 0&0&-1&0\\ 0&0&0&0\end{array}\right),\qquad\epsilon^{1}_{\mu\nu}=\left(\begin{array}[]{cccc}0&0&0&0\\ 0&1&0&0\\ 0&0&0&0\\ 0&0&0&-1\end{array}\right),$
$\epsilon^{x}_{\mu\nu}=\left(\begin{array}[]{cccc}0&0&0&0\\ 0&0&1&0\\ 0&1&0&0\\ 0&0&0&0\end{array}\right),\qquad\epsilon^{2}_{\mu\nu}=\left(\begin{array}[]{cccc}0&0&0&0\\ 0&0&0&1\\ 0&0&0&0\\ 0&1&0&0\end{array}\right),\qquad\epsilon^{3}_{\mu\nu}=\left(\begin{array}[]{cccc}0&0&0&0\\ 0&0&0&0\\ 0&0&0&1\\ 0&0&1&0\end{array}\right).$ (2.9)
Defining the sum over polarizations as
$P_{\mu\nu,\alpha\beta}=\sum_{\lambda}\epsilon^{\lambda}_{1\mu\nu}\epsilon^{\lambda}_{1\alpha\beta},$
it follows that the non zero components are
$P_{00,00}=0,\qquad P_{11,11}=2,\qquad P_{22,22}=P_{33,33}=1,$
$P_{12,12}=P_{21,21}=P_{13,13}=P_{31,31}=P_{23,23}=P_{32,32}=1,$
$P_{12,21}=P_{21,12}=P_{13,31}=P_{31,13}=P_{23,32}=P_{32,23}=1.$
By making a Lorenz transformation to a moving frame $k^{\mu}=(\omega(k),0,0,k)$
$\Lambda^{\mu}_{\nu}=\frac{1}{m}\left(\begin{array}[]{cccc}\omega&0&0&k\\ 0&1&0&0\\ 0&0&1&0\\ k&0&0&\omega\end{array}\right),$ (2.10)
the sum of polarizations can be calculated as
$P^{\prime}_{\mu\nu,\alpha\beta}=\Lambda^{\gamma}_{\mu}\Lambda^{\delta}_{\nu}\Lambda^{\epsilon}_{\alpha}\Lambda^{\kappa}_{\beta}P_{\gamma\delta,\epsilon\kappa}$
The sum depends on the masses. The polarization tensors in the new frame can also be calculated by making Lorentz transformations to the above ones (2.9).
There is one further important point. All the above discussion has shown that $h=0$, therefore (2.5) becomes
$\bigg(1-\frac{\kappa^{2}\alpha}{4}\Box\bigg)\square h_{\mu\nu}=0.$ (2.11)
But this implies that $h^{3}_{\mu\nu}=0$, since its corresponding mass $m_{3}$ is not present in this equation. Therefore (2.12) becomes
$h^{\mu\nu}(x,t)=\sum_{p}\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}\omega_{1}(k)}[\epsilon_{1}^{p\mu\nu}(k)e^{i\omega_{1}(k)t-ik\cdot x}+\epsilon_{1}^{\ast p\mu\nu}(k)e^{-i\omega_{1}(k)t+ik\cdot x}]$
$\sum_{p}\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}\omega_{2}(k)}[\epsilon_{2}^{p\mu\nu}(k)e^{i\omega_{2}(k)t-ik\cdot x}+\epsilon_{2}^{\ast p\mu\nu}(k)e^{-i\omega_{2}(k)t+ik\cdot x}].$ (2.12)
This leads to the strange conclusion that the third mode $h^{3}_{\mu\nu}$ is gone as a plane wave.
Assume that the above gauge fixing procedure is the problem for finding this unusual conclusion. Choose instead a healthier way to impose gauge conditions such as the De Donder gauge (2.4). One has to be sure that the above conclusion is not arising due to a bad fixing procedure. The equations of motion (2.3) in this gauge are
$\bigg(1+\frac{\kappa^{2}\alpha}{4}\Box\bigg)\square h_{\mu\nu}-\eta_{\mu\nu}\bigg[\frac{2}{3}\bigg(1+\frac{\kappa^{2}(3\beta-\alpha)}{2}\square\bigg)+\frac{1}{3}\bigg(1+\frac{\kappa^{2}\alpha}{4}\square\bigg)\bigg]\square h$
$2\bigg(1+\frac{\kappa^{2}\alpha}{4}\Box\bigg)\eta_{\gamma(\mu}\partial^{\gamma}\partial_{\nu)}h+\frac{1}{2}\eta_{\mu\nu}\bigg[\frac{2}{3}\bigg(1-\frac{\kappa^{2}(3\beta-\alpha)}{2}\square\bigg)+\frac{1}{3}\bigg(1-\frac{\kappa^{2}\alpha}{4}\square\bigg)\bigg]\square h$
$\bigg[\frac{2}{3}\bigg(1+\frac{\kappa^{2}(3\beta-\alpha)}{2}\square\bigg)+\frac{1}{3}\bigg(1+\frac{\kappa^{2}\alpha}{4}\square\bigg)\bigg]\partial_{\mu}\partial_{\nu}h-\kappa^{2}(\beta-\frac{\alpha}{2})\square\partial_{\mu}\partial_{\nu}h=0.$ (2.13)
Contracting the last differential equation with $\eta^{\mu\nu}$ gives
$\bigg(1+\frac{\kappa^{2}\alpha}{4}\Box\bigg)\square h-4\bigg[\frac{2}{3}\bigg(1+\frac{\kappa^{2}(3\beta-\alpha)}{2}\square\bigg)+\frac{1}{3}\bigg(1+\frac{\kappa^{2}\alpha}{4}\square\bigg)\bigg]\square h$
$2\bigg(1+\frac{\kappa^{2}\alpha}{4}\Box\bigg)\square h+2\bigg[\frac{2}{3}\bigg(1-\frac{\kappa^{2}(3\beta-\alpha)}{2}\square\bigg)+\frac{1}{3}\bigg(1-\frac{\kappa^{2}\alpha}{4}\square\bigg)\bigg]\square h$
$\bigg[\frac{2}{3}\bigg(1+\frac{\kappa^{2}(3\beta-\alpha)}{2}\square\bigg)+\frac{1}{3}\bigg(1+\frac{\kappa^{2}\alpha}{4}\square\bigg)\bigg]\square h-\kappa^{2}(\beta-\frac{\alpha}{2})\square^{2}h=0.$
Simplifying terms
$\bigg[\frac{2}{3}\bigg(1+\frac{\kappa^{2}(3\beta-\alpha)}{2}\square\bigg)+\frac{4}{3}\bigg(1+\frac{\kappa^{2}\alpha}{4}\square\bigg)\bigg]\square h-\kappa^{2}(\beta-\frac{\alpha}{2})\square^{2}h=0.$ (2.14)
Therefore only $h^{1}$ may survive, as for $h^{2}$ and $h^{3}$ the last gives an algebraic equation for these modes with solution zero. For instance, for the mode corresponding to $m^{2}_{2}$ one finds
$\frac{2}{3}\bigg(1-\frac{2(3\beta-\alpha)}{\alpha}\bigg)m_{2}^{2}h^{2}+\kappa^{2}(\beta-\frac{\alpha}{2})m_{2}^{2}h^{2}=0.$
The solution is $h^{2}=0$. The same conclusion goes for $h_{3}$, which is $h_{3}=0$. The massive modes are then given by (2.9). For the massless mode, one still has to satisfy the De Donder gauge (2.4). But this situation is known in GR that it constraints the value $h^{1}$ to vanish. To see this, note that the De Donder condition for the polarizations lead to
$2k^{\mu}\epsilon^{1}_{\mu\nu}=k_{\nu}\epsilon^{1}$
This eliminates $4$ polarizations, leading to $6$ independent ones. A further gauge transformation
$h_{\mu\nu}\to h_{\mu\nu}+\partial_{\mu}\xi_{\nu}+\partial_{\nu}\xi_{\mu}.$
This eliminates $4$ adeitional degrees of freedom. The vector $\xi_{\nu}$ can be chosen such that the polarization is (2.8), see for instance formulas (2.37) to (2.43) of reference [158]. Therefore $h_{1}$ can be chosen to be zero. But then $h=0$ and the equation (2.13) reduces to
$\bigg(1-\frac{\kappa^{2}\alpha}{4}\Box\bigg)\square h_{\mu\nu}=0.$
Therefore the equation (2.11) has been found again, implying that $h^{3}_{\mu\nu}=0$.
One may include a more general argument. The equation (2.14) is an equation for the trace of the form
$(1-\frac{\square}{m_{4}^{2}})\square h=0,$
where $m^{2}_{4}$ is a new scale. Then take $\partial^{\mu}$ to the equation (2.13), employ the De Donder gauge in the first term, go to the momentum space and multiply by $k^{\nu}$. Then a relation analogous to (2.7) will be found namely
$F(m_{i}^{2})k_{\mu}\epsilon_{4}=0,$ (2.15)
with $F(m_{i}^{2})$ a function of all the mass scales. Then $\epsilon_{4}=0$ and only the massless mode survives, which the De Donder gauge converts its polarization into a traceless and ”orthogonal” to the propagation direction.
The above conclusion shows that the third mode can not be turned on as a ”plane wave”. However, this does not forbid this term to generate Yukawa like potentials $\frac{e^{-m_{3}r}}{r}$ in spherically symmetric solutions of gravity [32]-[35]. In some sense, the absence of this mode in the free modes imitates confinement. On the other hand, the mode can be seen in classical spherical solutions. It is not clear yet for the author to refer this mode as confined or virtual, or even to employ a new terminology.
One may argue that, at quantum level, the third mode should not be eliminated. Instead, one should employ the Gupta-Bleuler quantization, which imposes weaker conditions at operator level. It will be shown below that nevertheless, even taking into account this quantization, the third mode disappears. This observation raises a very subtle detail. Consider for instance the formula (4.74) for the quartic Pais-Uhlenbeck scalar field $\phi$ given below. These commutator of the quantum field $\phi$ defines the propagator. Here it does not. This is simple to see. If the graviton is quantized and the third mode is absent (as it will be shown below at quantum level), then $\hat{h}_{\alpha\beta}$ does not contain $m_{3}^{2}$. Neither does the commutator
$\hat{h}_{\mu\nu},\hat{h}_{\alpha\beta}]=F(m_{2}^{2},k)$
However, the wave operator and therefore the gauge fixed lagrangian leading to the equations (2.13) characterizing the graviton does depend on the mass scale $m_{3}^{2}$. Therefore
$\hat{h}_{\mu\nu},\hat{h}_{\alpha\beta}]\neq D_{\alpha\beta,\mu\nu},$ (2.16)
where $D_{\alpha\beta,\mu\nu}$ is the propagator of the model. Situations such as (4.74) are lucky, not the general rule. This observation will be fundamental in the analysis below.
The next subsection calculates the propagator on the right hand side of (2.16), thereby confirming that there is no equality between the left and right hand side of this formula.
2.2 The free action in the Stelle gauge
The original reference [31] employs a gauge which is a generalization of the Lorenz gauge employed above, but with the particularity that induces terms which decay as $k^{-4}$. This behavior makes the renormalization analysis presented in that reference more accessible. It is not forbidden to employ other gauges, however, the renormalizability issues raised in that classic reference are much harder to be achieved. For this reason, it is important to discuss this gauge in some extent. Its use may be avoided when dealing with applications related to classical limits for instance, but it is more suited for renormalizability issues.
In some contexts, it is sometimes more convenient to express the action (2.1) and its equations of motion in terms of the Barnes- Rivers operators $P^{a}_{\mu\nu\alpha\beta}$, which are given in terms of the elementary tensors,
$\omega_{\mu\nu}=\frac{k_{\mu}k_{\nu}}{k^{2}},\qquad\theta_{\mu\nu}=\eta_{\mu\nu}-\frac{k_{\mu}k_{\nu}}{k^{2}},$ (2.17)
in the following way
$P^{2}_{\alpha\beta,\mu\nu}=\frac{1}{2}(\theta_{\beta\mu}\theta_{\alpha\nu}+\theta_{\beta\nu}\theta_{\alpha\mu})-\frac{1}{3}\theta_{\beta\alpha}\theta_{\mu\nu}$
$P^{1}_{\alpha\beta,\mu\nu}=\frac{1}{2}(\theta_{\beta\mu}\omega_{\alpha\nu}+\theta_{\beta\nu}\omega_{\alpha\mu}+\theta_{\alpha\mu}\omega_{\beta\nu}+\theta_{\alpha\nu}\omega_{\beta\mu})$
$P^{0-s}_{\alpha\beta,\mu\nu}=\frac{1}{3}\theta_{\beta\alpha}\theta_{\mu\nu},\qquad P^{0-w}_{\alpha\beta,\mu\nu}=\omega_{\beta\alpha}\omega_{\mu\nu},$
$P^{0-sw}_{\alpha\beta,\mu\nu}=\frac{1}{\sqrt{3}}\theta_{\beta\alpha}\omega_{\mu\nu},\qquad P^{0-ws}_{\alpha\beta,\mu\nu}=\frac{1}{\sqrt{3}}\omega_{\beta\alpha}\theta_{\mu\nu}.$ (2.18)
These formulas may be expressed in the coordinate space by making the change
$\omega_{\mu\nu}=\frac{\partial_{\mu}\partial_{\nu}}{\square},\qquad\theta_{\mu\nu}=\eta_{\mu\nu}-\frac{\partial_{\mu}\partial_{\nu}}{\square}.$
resulting in
$P_{\mu\nu\alpha\beta}^{2}=\frac{2}{3\square^{2}}\partial_{\mu}\partial_{\nu}\partial_{\alpha}\partial_{\beta}+\frac{1}{3\square}(\partial_{\mu}\partial_{\nu}\eta_{\alpha\beta}+\partial_{\alpha}\partial_{\beta}\eta_{\mu\nu})$
$\frac{1}{2\square}(\partial_{\mu}\partial_{\alpha}\eta_{\nu\beta}+\partial_{\mu}\partial_{\beta}\eta_{\nu\alpha}+\partial_{\nu}\partial_{\beta}\eta_{\mu\alpha}+\partial_{\alpha}\partial_{\nu}\eta_{\mu\beta})$
$\frac{1}{2}(\eta_{\mu\alpha}\eta_{\nu\beta}+\eta_{\nu\alpha}\eta_{\mu\beta})-\frac{1}{3}\eta_{\mu\nu}\eta_{\alpha\beta},$
$P_{\mu\nu\alpha\beta}^{1}=-\frac{1}{\square^{2}}\partial_{\mu}\partial_{\nu}\partial_{\alpha}\partial_{\beta}+\frac{1}{2\square}(\partial_{\mu}\partial_{\alpha}\eta_{\nu\beta}+\partial_{\mu}\partial_{\beta}\eta_{\nu\alpha}+\partial_{\nu}\partial_{\beta}\eta_{\mu\alpha}+\partial_{\alpha}\partial_{\nu}\eta_{\mu\beta}),$
$P_{\mu\nu\alpha\beta}^{0-s}=-\frac{1}{3\square}(\partial_{\alpha}\partial_{\beta}\eta_{\mu\nu}+\partial_{\mu}\partial_{\nu}\eta_{\alpha\beta})+\frac{1}{3}\eta_{\mu\nu}\eta_{\alpha\beta},\qquad P_{\mu\nu\alpha\beta}^{0-w}=\frac{1}{\square^{2}}\partial_{\mu}\partial_{\nu}\partial_{\alpha}\partial_{\beta},$
$P^{0-sw}_{\mu\nu\alpha\beta}=\frac{1}{\sqrt{3}\square}\eta_{\mu\nu}\partial_{\alpha}\partial_{\beta}-\frac{1}{\sqrt{3}\square^{2}}\partial_{\mu}\partial_{\nu}\partial_{\alpha}\partial_{\beta},$
$P^{0-ws}_{\alpha\beta,\mu\nu}=\frac{1}{\sqrt{3}\square}\eta_{\alpha\beta}\partial_{\mu}\partial_{\nu}-\frac{1}{\sqrt{3}\square^{2}}\partial_{\mu}\partial_{\nu}\partial_{\alpha}\partial_{\beta}.$ (2.19)
The definition of the Barnes-Rivers operators shows that
$(P^{2}+P^{1}+P^{0-2}+P^{0-w})_{\mu\nu}^{\;\;\;\alpha\beta}=\frac{1}{2}(\delta^{\alpha}_{\mu}\delta^{\beta}_{\nu}+\delta^{\alpha}_{\nu}\delta^{\beta}_{\mu}).$
In other words, the last sum is an identity. Therefore
$(P^{2}+P^{1}+P^{0-2}+P^{0-w})_{\mu\nu}^{\;\;\;\alpha\beta}h^{\mu\nu}=h^{\alpha\beta}.$ (2.20)
In addition the set composed by these four operators are such that
$P^{i-a}P^{j-b}=\delta^{ij}\delta^{ab}P^{j-b},$ (2.21)
where $i,j$ takes values between 0 and $2$ and $a,b$ takes values $s$ and $w$. Furthermore, the mixed operators $P^{0-ws}$ and $P^{0-sw}$ are such
$P^{0-ws}P^{0-sw}=P^{0-w},\qquad P^{0-sw}P^{0-ws}=P^{0-s},$
and their non zero actions with the other is
$P^{i-ab}P^{j-c}=\delta^{ij}\delta^{bc}P^{j-ac},\qquad P^{i-a}P^{j-bc}=\delta^{ij}\delta^{ab}P^{j-ac}.$
Note that the first four operators are projection operators, due to (2.21).
The terms of the linearized Stelle action (2.1) can be worked out in terms of the Barnes-Rivers operators as follows. Consider a typical term of the linearized action, mapped to momentum space
$h^{\alpha\beta}\eta_{\alpha\mu}\eta_{\beta\nu}\square^{2}h^{\mu\nu}\to k^{4}h^{\alpha\beta}\eta_{\alpha\mu}\eta_{\beta\nu}h^{\mu\nu}.$
By employing the definition (2.17), it is evident that $\eta_{\mu\nu}=\theta_{\mu\nu}+\omega_{\mu\nu}$, and with this simple identity the last term in momentum space may be worked out as
$k^{4}h^{\alpha\beta}\eta_{\alpha\mu}\eta_{\beta\nu}h^{\mu\nu}=k^{4}h^{\alpha\beta}(\theta_{\alpha\mu}\theta_{\beta\nu}+\theta_{\alpha\mu}\omega_{\beta\nu}+\omega_{\alpha\mu}\theta_{\beta\nu}+\omega_{\alpha\mu}\omega_{\beta\nu})h^{\mu\nu}$
$k^{4}h^{\alpha\beta}(P^{2}_{\alpha\beta,\mu\nu}+P^{0}_{\alpha\beta,\mu\nu}+P^{1}_{\alpha\beta,\mu\nu}+P^{0-w}_{\alpha\beta,\mu\nu})h^{\mu\nu}$
In the last identity, the definition of the Barnes Rivers operators (2.18) was employed. After going back to coordinate space, and doing the same with the other terms of the linearized action, the result is
$S=\frac{2}{\kappa^{2}}\int h^{\mu\nu}\square\bigg[\frac{1}{2}\bigg(1-\frac{\kappa^{2}\alpha}{4}\square\bigg)P^{2}-\bigg(1+\frac{\kappa^{2}(\alpha-3\beta)}{2}\square\bigg)P^{0-s}$
$\bigg(\frac{1}{2}+\frac{\kappa^{2}(4\beta-3\alpha)\square}{8}\bigg)P^{0-w}\bigg]_{\alpha\beta,\mu\nu}h^{\alpha\beta}d^{4}x.$ (2.22)
There are several identities satisfied for these operators, which have been worked out in the literature, for instance in [46]. One of the advantages of expressing the action in terms of the Barnes-Rivers operators is that, given an expression in terms of the Barnes-Rivers operators
$M=a_{2}P^{2}+a_{1}P^{1}+a_{s}P^{0-s}+a_{w}P^{0-w}+a_{sw}\sqrt{3}(P^{0-sw}+P^{0-ws}),$
its inverse is
$M^{-1}=\frac{1}{a_{2}}P^{2}+\frac{1}{a_{1}}P^{1}+\frac{1}{a_{s}a_{w}-3a_{sw}^{2}}\bigg[a_{w}P^{0-s}+a_{s}P^{0-w}-a_{sw}\sqrt{3}(P^{0-sw}+P^{0-ws})\bigg].$ (2.23)
Here the inverse means that
$M_{\mu\nu\gamma\delta}(M^{-1})^{\gamma\delta\alpha\beta}=(M^{-1})^{\gamma\delta\alpha\beta}M_{\mu\nu\gamma\delta}=\frac{1}{2}(\delta^{\alpha}_{\mu}\delta^{\beta}_{\nu}+\delta^{\alpha}_{\nu}\delta^{\beta}_{\mu}).$
The formula (2.23) shows the advantage of employing these operators. If one has an action written as
$S=\int h^{\mu\nu}(x)\hat{O}_{\mu\nu\alpha\beta}h^{\alpha\beta}(x)d^{4}x,$
with the operator $\hat{O}_{\mu\nu\alpha\beta}$ expressed in terms of the Barnes-Rivers operators, then (2.23) gives directly the propagator $\hat{D}_{\mu\nu\alpha\beta}=\hat{O}^{-1}_{\mu\nu\alpha\beta}$. Of course, there is no inverse for the operator in (2.22), since it is not gauge fixed and therefore non invertible. It is easy to see why it is not invertible. The coefficient for $P^{1}$ is zero and the inverse involves the inverse of the coefficient, which is infinite.
Now, a gauge has to be specified in order to obtain an invertible kinetic operator, leading to a non divergent propagtor. The gauge fixing term that the references [31] is the following
$L_{gf}=-\frac{\kappa^{2}}{2\lambda}F_{\tau}\square F^{\tau}.$ (2.24)
Here $F^{\tau}=\partial_{\nu}h^{\tau\nu}$. The use of this gauge may be advantageous in some applications and non practical in others. The advantage of it is that it leads to terms in the gauge fixed propagators which behave like $k^{-4}$ and which facilitate the renormalization program of that reference. In other applications, it may be convenient to employ the de Donder gauge, but the renormalizability of the model, although not spoiled, is considerably obscured.
The gauge fixed action in the Stelle gauge follows by expressing the gauge fixing part in terms of the Barnes-Rivers operators, the result is
$S=\frac{2}{\kappa^{2}}\int h^{\mu\nu}\square\bigg[\frac{1}{2}\bigg(1-\frac{\kappa^{2}\alpha}{4}\square\bigg)P^{2}-\bigg(1+\frac{\kappa^{2}(\alpha-3\beta)}{2}\square\bigg)P^{0-s}$
$\bigg(\frac{1}{2}+\frac{\kappa^{2}(4\beta-3\alpha)\square}{8}\bigg)P^{0-w}+\frac{\kappa^{4}}{8\lambda}\square[P^{1}-P^{0-w}]\bigg]_{\alpha\beta,\mu\nu}h^{\alpha\beta}d^{4}x.$
The inverse of the kinetic operator that can be read from this expression is the gauge fixed propagator, by employing (2.23)
$D_{\alpha\beta,\mu\nu}=\frac{1}{\square}\bigg[\frac{2}{1-\frac{\kappa^{2}\alpha}{4}\square}P^{2}+\frac{16\lambda}{\kappa^{4}\square^{2}}P^{1}-\frac{1}{1-\frac{\kappa^{2}}{2}(3\beta-\alpha)\square}P^{0-s}$
$\frac{2}{1+\frac{\kappa^{4}}{8\lambda}\square+\kappa^{2}(\beta-\frac{3}{4}\alpha)\square}P^{0-w}\bigg]_{\alpha\beta,\mu\nu}.$
It is seen that the gauge independent part namely, the one that corresponds to $\lambda\to 0$, involves three mass scales
$m_{1}^{2}=0,\qquad m_{2}^{2}=\frac{4}{\kappa^{2}\alpha},\qquad m_{3}^{2}=\frac{2}{\kappa^{2}(3\beta-\alpha)}.$ (2.25)
These are the mass scales anticipated in the previous section.
Clearly, the propagator depends on the three mass scales, and therefore the formula (2.16) follows.
2.3 Second quantization of the model
2.3.1 The Gupta-Bleuler scheme
Consider again the Lorenz gauge $\partial_{\mu}h^{\mu\nu}=0$. When making quantization of the model, one should be aware of imposing these classical considerations for the polarization operators. This awareness may be inspired by lessons about the Gupta-Bleuler quantization. Sometimes to impose a gauge at operator level, such as the Lorenz gauge $\partial_{\mu}A^{\mu}=0$ in QED, may be an over request. Instead, imposing a softer condition to the physical states which insures that the Lorenz gauge condition is assured when taking mean values make quantization realizable. An example is $\partial_{\mu}A^{+\mu}|\Psi>=0$, a condition heavily employed in this quantization method.
Consider the non gauged equations of motion without fixing given in (2.22). As before, the Lorenz gauge $\partial_{\mu}h^{\mu\nu}=0$ will be considered. However, both (2.22) and the gauge has to be valid only when mean values between physical states $|\Psi>$ are taken. Therefore the Gupta-Bleuler request is
$\partial^{+}_{\mu}h^{\mu\nu}|\Psi>=0.$
This implies that
$\sum_{\lambda}k^{\mu}\epsilon^{\lambda}_{1\mu\nu}a_{\lambda}(k)|\Psi>=0,\qquad\sum_{\lambda}k^{\mu}\epsilon^{\lambda}_{2\mu\nu}b_{\lambda}(k)|\Psi>=0,$
$\sum_{\lambda}k^{\mu}\epsilon^{\lambda}_{3\mu\nu}c_{\lambda}(k)|\Psi>=0.$ (2.26)
Here $\lambda$ included all the polarizations, not only (2.8) and (2.9). In fact (2.8) and (2.9) do not contribute to this sum, since they are such that $k^{\mu}\epsilon^{\lambda}_{\mu\nu}=0$. The last gives a linear relation between the annihilation operators corresponding to these additional polarizations, whose action destroys the physical states. The averaged equations of motion are now
$\bigg(1+\frac{\kappa^{2}\alpha}{4}\Box\bigg)\square h_{\mu\nu}>-\eta_{\mu\nu}<\bigg[\frac{2}{3}\bigg(1+\frac{\kappa^{2}(3\beta-\alpha)}{2}\square\bigg)+\frac{1}{3}\bigg(1+\frac{\kappa^{2}\alpha}{4}\square\bigg)\bigg]\square h>$
$\bigg[\frac{2}{3}\bigg(1+\frac{\kappa^{2}(3\beta-\alpha)}{2}\square\bigg)+\frac{1}{3}\bigg(1+\frac{\kappa^{2}\alpha}{4}\square\bigg)\bigg]\partial_{\mu}\partial_{\nu}h>=0.$
Contracting the last differential equation with $\eta^{\mu\nu}$ gives
$\bigg(1+\frac{\kappa^{2}(3\beta-\alpha)}{2}\square\bigg)\square h>=0.$
The fact that the mean values do not contain the massive mode with $m_{2}^{2}$ imply that
$\sum_{\lambda}\epsilon_{2}^{\lambda}b_{\lambda}(k)|\Psi>=0.$ (2.27)
The states (2.8) and (2.9) have zero traces. The remaning are not traceless and have non zero $\epsilon^{\lambda}$. The condition (2.27) gives a new linear relation for the annihilation operators. The averaged equations of motion reduce to
$\bigg(1+\frac{\kappa^{2}\alpha}{4}\Box\bigg)\square h_{\mu\nu}>-<\frac{1}{3}\eta_{\mu\nu}\bigg(1+\frac{\kappa^{2}\alpha}{4}\square\bigg)\square h>$
$\bigg[\frac{2}{3}\bigg(1+\frac{\kappa^{2}(3\beta-\alpha)}{2}\square\bigg)+\frac{1}{3}\bigg(1+\frac{\kappa^{2}\alpha}{4}\square\bigg)\bigg]\partial_{\mu}\partial_{\nu}h>=0.$
The absence of the mode with $m_{2}^{2}$ in the average leads to
$\bigg(1-\frac{\kappa^{2}\alpha}{4}m_{3}^{2}\bigg)m_{3}^{2}<h^{3}_{\mu\nu}>+\frac{1}{3}\eta_{\mu\nu}m_{3}^{2}\bigg(1-\frac{\kappa^{2}\alpha}{4}m_{3}^{2}\bigg)<h_{3}>$
$\partial_{\mu}\partial_{\nu}<h_{1}>+\frac{1}{3}\bigg(1-\frac{\kappa^{2}\alpha}{4}m_{3}^{2}\bigg)\partial_{\mu}\partial_{\nu}<h_{3}>=0.$ (2.28)
The independence between the modes $1$ and $3$ leads to
$k_{\mu}k_{\nu}\sum_{\lambda}\epsilon_{1}^{\lambda}a_{\lambda}(k)|\Psi>=0,$
or simply
$\sum_{\lambda}\epsilon_{1}^{\lambda}a_{\lambda}(k)|\Psi>=0,$ (2.29)
and also
$\frac{1}{\omega_{3}(k)}\bigg(1-\frac{\kappa^{2}\alpha}{4}m_{3}^{2}\bigg)m_{3}^{2}\sum_{\lambda}\epsilon^{3}_{\lambda\mu\nu}c^{\lambda}(k)|\Psi>$
$\frac{1}{3\omega_{3}(k)}\eta_{\mu\nu}m_{3}^{2}\bigg(1-\frac{\kappa^{2}\alpha}{4}m_{3}^{2}\bigg)\sum_{\lambda}\epsilon_{\lambda 3}c^{\lambda}(k)|\Psi>$
$\frac{1}{3\omega_{3}(k)}\bigg(1-\frac{\kappa^{2}\alpha}{4}m_{3}^{2}\bigg)k^{3}_{\mu}k^{3}_{\nu}\epsilon_{\lambda 3}c^{\lambda}(k)|\Psi>=0,$
Contracting the first with $k^{3\mu}$ and taking into account the third (2.26) leads to
$\sum_{\lambda}\epsilon_{3}^{\lambda}c_{\lambda}(k)|\Psi>=0,$ (2.30)
The conditions (2.26)-(2.30) are the Gupta-Bleuler constraints on the physical states. The sums here involve all the polarizations, not only the classically allowed (2.8) and (2.9).
2.3.2 Heuristic issues of the model quantization
If one is intended to quantize the model by employing second quantization, then the Ostrogradsky method results in the introduction of a field
$h^{\mu\nu}=h^{1\mu\nu}+h^{2\mu\nu}+h^{3\mu\nu}$
$\sum_{p}\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}\omega_{1}(k)}[a_{p}(k)\epsilon_{1}^{p\mu\nu}(k)e^{i\omega_{1}(k)t-ik\cdot x}+a^{c}_{p}(k)\epsilon_{1}^{\ast p\mu\nu}(k)e^{-i\omega_{1}(k)t+ik\cdot x}]$
$\sum_{p}\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}\omega_{2}(k)}[b_{p}(k)\epsilon_{2}^{p\mu\nu}(k)e^{i\omega_{2}(k)t-ik\cdot x}+b^{c}_{p}(k)\epsilon_{2}^{\ast p\mu\nu}(k)e^{-i\omega_{2}(k)t+ik\cdot x}]$
$\sum_{p}\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}\omega_{3}(k)}[c_{p}(k)\epsilon_{3}^{p\mu\nu}(k)e^{i\omega_{3}(k)t-ik\cdot x}+c^{c}_{p}(k)\epsilon_{3}^{\ast p\mu\nu}(k)e^{-i\omega_{3}(k)t+ik\cdot x}],$ (2.31)
where the component $h^{i\mu\nu}$ corresponds to the mass $m_{i}$. Note that symbol $\dagger$} is replaced by the letter c of ”creation”, anticipating for instance that $a^{c}_{p}(k)$ may not be exactly the adjoint of $a_{p}(k)$. An imaginary unit may be present, or other constants. The time derivatives of the field is
$\dot{h}^{\mu\nu}=\dot{h}^{1\mu\nu}+\dot{h}^{2\mu\nu}+\dot{h}^{3\mu\nu}$
$i\sum_{p}\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}}[a_{p}(k)\epsilon_{1}^{p\mu\nu}(k)e^{i\omega_{1}(k)t-ik\cdot x}-a^{c}_{p}(k)\epsilon_{1}^{\ast p\mu\nu}(k)e^{-i\omega_{1}(k)t+ik\cdot x}]$
$i\sum_{p}\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}}[b_{p}(k)\epsilon_{2}^{p\mu\nu}(k)e^{i\omega_{2}(k)t-ik\cdot x}-b^{c}_{p}(k)\epsilon_{2}^{\ast p\mu\nu}(k)e^{-i\omega_{2}(k)t+ik\cdot x}]$
$i\sum_{p}\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}}[c_{p}(k)\epsilon_{3}^{p\mu\nu}(k)e^{i\omega_{3}(k)t-ik\cdot x}-c^{c}_{p}(k)\epsilon_{3}^{\ast p\mu\nu}(k)e^{-i\omega_{3}(k)t+ik\cdot x}].$ (2.32)
There are two polarizations for the massless mode and five for the massive ones. However, the sum can be generalized to five by postulating that three polarizations for the massless mode are zero. Furthermore, at quantum level, one has to assume that there are further polarization and the states are such that their action leads to zero mean values, which is the indication of the Gupta-Bleuler formalism (2.26).-(2.30).
The coordinates for the four order theory are given by
$Q_{1\mu\nu}=h_{\mu\nu},\qquad Q_{2\mu\nu}=\dot{h}_{\mu\nu},$
while the momentum without imposing the gauge is
$\frac{\kappa^{2}}{2}P_{2\alpha\beta}=\frac{\kappa^{2}\alpha}{4}[\square h_{\alpha\beta}-\partial^{\gamma}\partial_{(\beta}h_{\alpha)\gamma}]-\frac{\kappa^{2}\alpha}{4}\square h_{0(\alpha}\delta_{\beta)0}$
$\frac{\kappa^{2}(4\beta-\alpha)}{4}\delta_{0\alpha}\delta_{0\beta}\square h+\frac{\kappa^{2}(4\beta-\alpha)}{4}\eta_{\alpha\beta}\partial_{\gamma}\partial_{\delta}h^{\gamma\delta}-\frac{\kappa^{2}(4\beta-\alpha)}{4}\eta_{\alpha\beta}\square h$
$2\kappa^{2}(\beta-\frac{\alpha}{2})\delta_{\alpha 0}\delta_{\beta 0}\partial_{\gamma}\partial_{\delta}h^{\gamma\delta},$
$\frac{\kappa^{2}}{2}P_{1\alpha\beta}=-\dot{h}_{\alpha\beta}+\eta_{\alpha\beta}\dot{h}-2\delta_{0(\alpha}\partial^{\sigma}h_{\beta)\sigma}+\eta_{\alpha\beta}\partial_{\epsilon}h^{0\epsilon}-\delta_{0(\alpha}\partial_{\beta)}h$
$\frac{\kappa^{2}\alpha}{4}[\square\dot{h}_{\alpha\beta}+\partial^{\gamma}\partial_{(\beta}\dot{h}_{\alpha)\gamma}]+\frac{\kappa^{2}\alpha}{4}\square\dot{h}_{0(\alpha}\delta_{\beta)0}$
$\frac{\kappa^{2}(4\beta-\alpha)}{4}\delta_{0\alpha}\delta_{0\beta}\square\dot{h}-\frac{\kappa^{2}(4\beta-\alpha)}{4}\eta_{\alpha\beta}\partial_{\gamma}\partial_{\delta}\dot{h}^{\gamma\delta}+\frac{\kappa^{2}(4\beta-\alpha)}{4}\eta_{\alpha\beta}\square\dot{h}$
$2\kappa^{2}(\beta-\frac{\alpha}{2})\delta_{\alpha 0}\delta_{\beta 0}\partial_{\gamma}\partial_{\delta}\dot{h}^{\gamma\delta}.$ (2.33)
Note that
$P_{1\alpha\beta}=\dot{h}_{\alpha\beta}-\eta_{\alpha\beta}\dot{h}+2\delta_{0(\alpha}\partial^{\sigma}h_{\beta)\sigma}-\eta_{\alpha\beta}\partial_{\epsilon}h^{0\epsilon}-\delta_{0(\alpha}\partial_{\beta)}h-\dot{P}_{2\alpha\beta}.$ (2.34)
Before doing large calculations, it is better to focus in simple features first. Forget for a moment that the above is a gauge theory and assume that $Q_{1}^{\mu\nu},Q_{2}^{\alpha\beta}]=0$ was ensured. This imitates the procedure for the Pauli-Uhlenbeck oscillator, defined below in (3.47). This implies that
$h^{\mu\nu},\dot{h}^{\alpha\beta}]=0.$
This, of course, establish a linear relation between $h_{1}^{\mu\nu},\dot{h}_{1}^{\alpha\beta}]$, $h_{2}^{\mu\nu},\dot{h}_{2}^{\alpha\beta}]$ and $h_{3}^{\mu\nu},\dot{h}_{3}^{\alpha\beta}]$.
On the other hand, taking into account the gauge conditions, the momentum in (2.33) reduces to
$\frac{\kappa^{2}}{2}P_{2p\alpha\beta}=\frac{\kappa^{2}\alpha}{4}m_{2}^{2}h^{2}_{\alpha\beta}+\frac{\kappa^{2}\alpha}{4}m_{3}^{2}h^{3}_{\alpha\beta}-\frac{\kappa^{2}\alpha}{4}m_{2}^{2}h_{0(\alpha}\delta_{\beta)0}-\frac{\kappa^{2}\alpha}{4}m_{3}^{2}h^{3}_{0(\alpha}\delta_{\beta)0},$
$\frac{\kappa^{2}}{2}P_{1\alpha\beta}=\dot{h}^{1}_{\alpha\beta}+\dot{h}^{2}_{\alpha\beta}+\dot{h}^{3}_{\alpha\beta}$
$\frac{\kappa^{2}\alpha}{4}m_{2}^{2}\dot{h}^{2}_{\alpha\beta}-\frac{\kappa^{2}\alpha}{4}m_{3}^{2}\dot{h}^{3}_{\alpha\beta}$
$\frac{\kappa^{2}\alpha}{4}m_{2}^{2}\dot{h}^{2}_{0(\alpha}\delta_{\beta)0}+\frac{\kappa^{2}\alpha}{4}m_{3}^{2}\dot{h}^{3}_{0(\alpha}\delta_{\beta)0}.$ (2.35)
Then the requirement $P_{1}^{\mu\nu},P_{2}^{\alpha\beta}]=0$ leads to a new condition relating $h_{i}^{\mu\nu},\dot{h}_{i}^{\alpha\beta}]$, with $i=1,2,3$.
Up to now, there are two homogenous relations for the commutators of each mode $h_{1}^{\mu\nu},\dot{h}_{1}^{\alpha\beta}]$, $h_{2}^{\mu\nu},\dot{h}_{2}^{\alpha\beta}]$ and $h_{3}^{\mu\nu},\dot{h}_{3}^{\alpha\beta}]$.
The last two commutators of interest are $P_{1p\alpha\beta},Q_{1p\alpha\beta}]$ and $P_{2p\alpha\beta},Q_{2p\alpha\beta}]$. Again, a large amount of terms in this expression may be avoided by the Gupta-Bleuler scheme. A very helpful formula is (2.34). It is clear from that formula, the gauge condition, and the definition of $Q_{1p\mu\nu}$, that
$P_{1p\alpha\beta},Q_{1p\alpha\beta}]=[P_{1p\alpha\beta},h_{p\alpha\beta}]=[\dot{h}_{1p\alpha\beta},h_{p\alpha\beta}]-[\dot{P}_{2p\alpha\beta},h_{p\alpha\beta}].$ (2.36)
However, since it has been ensured in (2.39) that $\dot{h}_{p\alpha\beta},h_{p\alpha\beta}]=0$, thus
$P_{1p\alpha\beta},Q_{1p\alpha\beta}]=[h_{p\alpha\beta},\dot{P}_{2p\alpha\beta}].$ (2.37)
On the other hand
$P_{2p\alpha\beta},Q_{2p\alpha\beta}]=[P_{2p\alpha\beta},\dot{h}_{p\alpha\beta}]$ (2.38)
It is not difficult to see that $P_{2p\alpha\beta},\dot{h}_{p\alpha\beta}]=-[h_{p\alpha\beta},\dot{P}_{2p\alpha\beta}]$. This is seen by looking the expressions (2.31), (2.32) and (2.35). Note that the time derivative only changes a sign in one term and multiplies everything by $i\omega_{i}(k)$. These considerations lead to this conclusion. This conclusion together with (2.37) and (2.38) leads to the expectation that
$P_{1\mu\nu}(x),Q_{1\alpha\beta}(y)]=\frac{i}{2}(\eta_{\alpha\mu}\eta_{\beta\nu}+\eta_{\alpha\nu}\eta_{\beta\mu})\delta(x-y),$
$P_{2\mu\nu}(x),Q_{2\alpha\beta}(y)]=-\frac{i}{2}(\eta_{\alpha\mu}\eta_{\beta\nu}+\eta_{\alpha\nu}\eta_{\beta\mu})\delta(x-y).$
As the sign in the second commutator is inverted, thus the variables labelled with $2$ are ghost like. The factor multiplying the identity is only the imaginary unit. This can be warranted by a third condition relating $h_{1}^{\mu\nu},\dot{h}_{1}^{\alpha\beta}]$, $h_{2}^{\mu\nu},\dot{h}_{2}^{\alpha\beta}]$ and $h_{3}^{\mu\nu},\dot{h}_{3}^{\alpha\beta}]$, this time non homogeneous.
Overall, there are three linear conditions for $h_{1}^{\mu\nu},\dot{h}_{1}^{\alpha\beta}]$, $h_{2}^{\mu\nu},\dot{h}_{2}^{\alpha\beta}]$ and $h_{3}^{\mu\nu},\dot{h}_{3}^{\alpha\beta}]$, two homogenous and one not. They may be enough for determining these commutators. However, the above analysis skipped one important issue.
2.3.3 More concrete issues about the quantization of the model
The expectations described above are intentionally naive, however, they capture some helpful features. As anticipated, they do not take into account that the theory is a gauge one. Experience with ordinary QED allows to suspect that the real situation is
$P_{1\mu\nu}(x),Q_{1\alpha\beta}(y)]=\frac{i}{2}[(\eta_{\alpha\mu}\eta_{\beta\nu}+\eta_{\alpha\nu}\eta_{\beta\mu})+\text{gauge terms}]\delta(x-y),$
$P_{2\mu\nu}(x),Q_{2\alpha\beta}(y)]=-\frac{i}{2}[(\eta_{\alpha\mu}\eta_{\beta\nu}+\eta_{\alpha\nu}\eta_{\beta\mu})+\text{other gauge terms}]\delta(x-y).$
In addition,
$Q_{1\mu\nu}(x),Q_{2\alpha\beta}(y)]=\text{gauge terms}.$
$P_{1\mu\nu}(x),P_{2\alpha\beta}(y)]=\text{other gauge terms}.$
Taking into account these expectations, consider again the commutator $Q_{1\mu\nu}(x),Q_{2\alpha\beta}(y)]$ or, what is the same, $h_{\mu\nu},\dot{h}_{\mu\nu}]$. It is given by, up to an integration in $k$, by the following expression
$\sum_{q}\sum_{p}\epsilon_{1}^{p\mu\nu}(k)\epsilon_{1}^{\ast p\mu\nu}(k^{\prime})\frac{[a_{p}(k),a_{q}^{c}(k^{\prime})]}{\omega_{1}(k)}+\epsilon_{2}^{p\mu\nu}(k)\epsilon_{2}^{\ast p\mu\nu}(k^{\prime})\frac{[b_{p}(k),b_{q}^{c}(k^{\prime})]}{\omega_{2}(k)}$
$\epsilon_{3}^{p\mu\nu}(k)\epsilon_{3}^{\ast p\mu\nu}(k^{\prime})\frac{[c_{p}(k),c_{q}^{c}(k^{\prime})]}{\omega_{3}(k)}.$ (2.39)
Suppose that
$a_{p}(k),a_{q}^{c}(k^{\prime})]=c_{1}\delta_{pq}\omega_{1}(k)\delta(k-k^{\prime}),\qquad[b_{p}(k),b_{q}^{c}(k^{\prime})]=c_{2}\delta_{pq}\omega_{2}(k)\delta(k-k^{\prime}),$
$c_{p}(k),c_{q}^{c}(k^{\prime})]=c_{3}\delta_{pq}\omega_{2}(k)\delta(k-k^{\prime}),$ (2.40)
the last constitutes one equation for three structure constants $c_{i}$ with $i=1,2,3$. It is difficult to request that everything cancels out, since the sum of polarizations involve different mass scales and different frequencies. However, one may require that the terms proportional to the identity
$I_{\mu\nu,\alpha\beta}=\frac{1}{2}(\eta_{\alpha\mu}\eta_{\beta\nu}+\eta_{\alpha\nu}\eta_{\beta\mu}).$
cancel each other . This leads to a homogeneous relation between the constants $c_{i}$.
Note that the fact $h_{\mu\nu},\dot{h}_{\mu\nu}]=$”gauge terms”, at least when mean values between physical states are taken, does not imply that $h_{i\mu\nu},\dot{h}_{i\mu\nu}]=$”gauge terms”, for each mass mode. The full sum is the one that vanishes, and also $h_{i\mu\nu},\dot{h}_{j\mu\nu}]=0$ for $i\neq j$.
Consider the momentum part of (2.33). The commutator $P_{1},P_{2}]$ is a cumbersome expression, but it contains a large number of terms that vanish when mean values between the physical states $|\Psi>$ are taken. It is better to avoid them from the beginning, although they are taken care by sandwich between physical states. The relevant part remaining is given in (2.35). It is better to rewrite this formula here, since it is crucial
$\frac{\kappa^{2}}{2}P_{2p\alpha\beta}=\frac{\kappa^{2}\alpha}{4}m_{2}^{2}h^{2}_{\alpha\beta}+\frac{\kappa^{2}\alpha}{4}m_{3}^{2}h^{3}_{\alpha\beta}-\frac{\kappa^{2}\alpha}{4}m_{2}^{2}h_{0(\alpha}\delta_{\beta)0}-\frac{\kappa^{2}\alpha}{4}m_{3}^{2}h^{3}_{0(\alpha}\delta_{\beta)0},$
$\frac{\kappa^{2}}{2}P_{1\alpha\beta}=\dot{h}^{1}_{\alpha\beta}+\dot{h}^{2}_{\alpha\beta}+\dot{h}^{3}_{\alpha\beta}-\frac{\kappa^{2}\alpha}{4}m_{2}^{2}\dot{h}^{2}_{\alpha\beta}-\frac{\kappa^{2}\alpha}{4}m_{3}^{2}\dot{h}^{3}_{\alpha\beta}$
$\frac{\kappa^{2}\alpha}{4}m_{2}^{2}\dot{h}^{2}_{0(\alpha}\delta_{\beta)0}+\frac{\kappa^{2}\alpha}{4}m_{3}^{2}\dot{h}^{3}_{0(\alpha}\delta_{\beta)0}.$
The fact that $m_{2}^{2}=\frac{4}{\kappa^{2}\alpha}$ simplifies several terms. The last expression becomes even simpler
$\frac{\kappa^{2}}{2}P_{2p\alpha\beta}=\frac{\kappa^{2}\alpha}{4}m_{2}^{2}h^{2}_{\alpha\beta}+\frac{\kappa^{2}\alpha}{4}m_{3}^{2}h^{3}_{\alpha\beta}-\frac{\kappa^{2}\alpha}{4}m_{2}^{2}h_{0(\alpha}\delta_{\beta)0}-\frac{\kappa^{2}\alpha}{4}m_{3}^{2}h^{3}_{0(\alpha}\delta_{\beta)0},$
$\frac{\kappa^{2}}{2}P_{1\alpha\beta}=\dot{h}^{1}_{\alpha\beta}+\dot{h}^{3}_{\alpha\beta}-\frac{\kappa^{2}\alpha}{4}m_{3}^{2}\dot{h}^{3}_{\alpha\beta}$
$\frac{\kappa^{2}\alpha}{4}m_{2}^{2}\dot{h}^{2}_{0(\alpha}\delta_{\beta)0}+\frac{\kappa^{2}\alpha}{4}m_{3}^{2}\dot{h}^{3}_{0(\alpha}\delta_{\beta)0}.$
Here only the polarizations (2.8) and (2.9) participate, the others are washed out by taking mean values (see below). The part of the commutator $P_{1p\mu\nu},P_{2p\mu\nu}]$ related to the massless mode $h^{1}_{\mu\nu}$ does not contribute, since it is not present in the expression of $P_{2p\mu\nu}$. The commutator will be an expression related to $b_{p}(k),b_{q}^{c}(k^{\prime})]$ and $c_{p}(k),c_{q}^{c}(k^{\prime})]$ multiplied by expression related by polarization tensors, analogous but not equal to (2.39). The commutator is quite simple. The only expression needed here will be
$\frac{\kappa^{4}}{4}[P_{1p\mu\nu},P_{2p\mu\nu}]=\frac{\kappa^{4}\alpha^{2}}{16}m_{2}^{4}[h^{2}_{\mu\nu},\dot{h}^{2}_{0(\alpha}\delta_{\beta)0}]$
[…]
Choose only spatial indices. Then the terms related to $h^{2}_{\mu\nu}$ are gone and one has
$\frac{\kappa^{4}}{4}[P_{1pij},P_{2pkl}]=\text{terms related to $h^{3}_{ij}$},$
or explicitly
$\frac{\kappa^{4}}{4}[P_{1pij},P_{2pkl}]=\frac{\kappa^{2}\alpha}{4}m_{3}^{2}\bigg(1-\frac{\kappa^{2}\alpha}{4}m_{3}^{2}\bigg)[h^{3}_{ij},\dot{h}^{3}_{kl}]$
The request $P_{1pij},P_{2pkl}]=$”gauge terms”, which is equivalent to the requirement that the terms proportional to the identity cancel out, will give a homegeneous equation for a single constant $c_{3}$ defined in (2.40). It is evident that $c_{3}=0$. Therefore
$c_{p}(k),c^{c}_{p}(k)]=0.$
The coordinates and momentum of this mode are classical. But is hard to put a classical particle in a quantum setting, since it will enter in conflict with the uncertainty principle. Therefore, the following conclusion has been shown.
Conclusion: There are no external lines related to the ghost mode $m_{3}^{2}$. Only internal lines are related to this ghost. The theory is ”free from the ghost mode 33” since there is no ”free ghost mode 33”.
The mode $3$ resembles to the Faddeev-Popov ghosts, except that is not clear that arises as an artifact of gauge fixing.
The above argument relies on throwing the terms that violate the classical gauge condition, by assuming that a sandwich between physical states will take care of these terms. It is convenient perhaps to write the full expression of the commutator, with all the irrelevant terms, and to clarify what does it mean that the terms which vanish with the gauge are annihilated by the physical states. The full commutator is this cumbersome expression
$P_{1\mu\nu},P_{2\mu\nu}]\equiv\sum_{p}\bigg[[\epsilon^{1}_{\alpha\beta}-\eta_{\alpha\beta}\epsilon^{1}+2\delta_{0(\alpha}ik^{\sigma}\epsilon^{1}_{\beta)\sigma}-\eta_{\alpha\beta}ik_{\epsilon}\epsilon^{(1)0\epsilon}-\delta_{0(\alpha}ik_{\beta)}\epsilon^{1}]$
$\frac{\kappa^{2}\alpha}{4}ik^{\gamma}ik_{(\beta}\epsilon^{1}_{\alpha)\gamma}-\bigg[\frac{\kappa^{2}(4\beta-\alpha)}{4}\eta_{\alpha\beta}+2\kappa^{2}(\beta-\frac{\alpha}{2})\delta_{\alpha 0}\delta_{\beta 0}\bigg]ik_{\gamma}ik_{\delta}\epsilon^{1\gamma\delta}\bigg]$
$\bigg[\frac{\kappa^{2}\alpha}{4}ik^{\gamma}ik_{(\beta}\epsilon^{1}_{\alpha)\gamma}-\bigg[\frac{\kappa^{2}(4\beta-\alpha)}{4}\eta_{\alpha\beta}+2\kappa^{2}(\beta-\frac{\alpha}{2})\delta_{\alpha 0}\delta_{\beta 0}\bigg]ik_{\gamma}ik_{\delta}\epsilon^{1\gamma\delta}\bigg]\frac{[a(k),a^{c}(k)]}{\omega_{1}(k)}$
$\bigg[[\epsilon^{2}_{\alpha\beta}-\eta_{\alpha\beta}\epsilon^{2}+2\delta_{0(\alpha}ik^{\sigma}\epsilon^{2}_{\beta)\sigma}-\eta_{\alpha\beta}ik_{\epsilon}\epsilon^{(2)0\epsilon}-\delta_{0(\alpha}ik_{\beta)}\epsilon^{2}]$
$\frac{\kappa^{2}\alpha}{4}ik^{\gamma}ik_{(\beta}\epsilon^{2}_{\alpha)\gamma}-\bigg[\frac{\kappa^{2}(4\beta-\alpha)}{4}\eta_{\alpha\beta}+2\kappa^{2}(\beta-\frac{\alpha}{2})\delta_{\alpha 0}\delta_{\beta 0}\bigg]ik_{\gamma}ik_{\delta}\epsilon^{2\gamma\delta}$
$\frac{\kappa^{2}\alpha m_{2}^{2}}{4}[\epsilon^{2}_{\alpha\beta}-\epsilon^{2}_{0(\alpha}\delta_{\beta)0}]+\frac{\kappa^{2}(4\beta-\alpha)}{4}(\delta_{0\alpha}\delta_{0\beta}+\eta_{\alpha\beta})m_{2}^{2}\eta^{\mu\nu}\epsilon_{\mu\nu}^{2}\bigg]$
$\bigg[\frac{\kappa^{2}\alpha}{4}ik^{\gamma}ik_{(\beta}\epsilon^{2}_{\alpha)\gamma}-\bigg[\frac{\kappa^{2}(4\beta-\alpha)}{4}\eta_{\alpha\beta}+2\kappa^{2}(\beta-\frac{\alpha}{2})\delta_{\alpha 0}\delta_{\beta 0}\bigg]ik_{\gamma}ik_{\delta}\epsilon^{2\gamma\delta}$
$\frac{\kappa^{2}\alpha m_{2}^{2}}{4}[\epsilon^{2}_{\alpha\beta}-\epsilon^{2}_{0(\alpha}\delta_{\beta)0}]+\frac{\kappa^{2}(4\beta-\alpha)}{4}(\delta_{0\alpha}\delta_{0\beta}+\eta_{\alpha\beta})m_{2}^{2}\eta^{\mu\nu}\epsilon_{\mu\nu}^{2}\bigg]\frac{[b(k),b^{c}(k)]}{\omega_{2}(k)}$
$\bigg[[\epsilon^{3}_{\alpha\beta}-\eta_{\alpha\beta}\epsilon^{3}+2\delta_{0(\alpha}ik^{\sigma}\epsilon^{3}_{\beta)\sigma}-\eta_{\alpha\beta}ik_{\epsilon}\epsilon^{(3)0\epsilon}-\delta_{0(\alpha}ik_{\beta)}\epsilon^{3}]$
$\frac{\kappa^{2}\alpha}{4}ik^{\gamma}ik_{(\beta}\epsilon^{3}_{\alpha)\gamma}-\bigg[\frac{\kappa^{2}(4\beta-\alpha)}{4}\eta_{\alpha\beta}+2\kappa^{2}(\beta-\frac{\alpha}{2})\delta_{\alpha 0}\delta_{\beta 0}\bigg]ik_{\gamma}ik_{\delta}\epsilon^{3\gamma\delta}$
$\frac{\kappa^{2}\alpha m_{3}^{2}}{4}[\epsilon^{3}_{\alpha\beta}-\epsilon^{3}_{0(\alpha}\delta_{\beta)0}]+\frac{\kappa^{2}(4\beta-\alpha)}{4}(\delta_{0\alpha}\delta_{0\beta}+\eta_{\alpha\beta})m_{3}^{2}\eta^{\mu\nu}\epsilon_{\mu\nu}^{3}\bigg]$
$\bigg[+\frac{\kappa^{2}\alpha}{4}ik^{\gamma}ik_{(\beta}\epsilon^{3}_{\alpha)\gamma}-\bigg[\frac{\kappa^{2}(4\beta-\alpha)}{4}\eta_{\alpha\beta}+2\kappa^{2}(\beta-\frac{\alpha}{2})\delta_{\alpha 0}\delta_{\beta 0}\bigg]ik_{\gamma}ik_{\delta}\epsilon^{3\gamma\delta}$
$\frac{\kappa^{2}\alpha m_{3}^{2}}{4}[\epsilon^{3}_{\alpha\beta}-\epsilon^{3}_{0(\alpha}\delta_{\beta)0}]+\frac{\kappa^{2}(4\beta-\alpha)}{4}(\delta_{0\alpha}\delta_{0\beta}+\eta_{\alpha\beta})m_{3}^{2}\eta^{\mu\nu}\epsilon_{\mu\nu}^{3}\bigg]\frac{[c(k),c^{c}(k)]}{\omega_{3}(k)}.$
Now act on physical states $|\Psi>$ either on the left or on the right, and employ (2.26)-(2.30) or its adjoint version. A large amount of terms are gone, and one obtains again the simpler expressions employed in this section.
At this point two homogenous conditons have been found. One is $c_{3}=0$ and the other is (2.39). The last condition needed in order to find these coefficients follows by requesting that the identity part of $P_{1\mu\nu},Q_{1\alpha\beta}]$ is multiplied by an imaginary $i$. This leads to a new linear and non homogenous equation for the structure constants $c_{i}$, thus allowing their calculation. The considerations given above show that automatically the other commutator $P_{2\mu\nu},Q_{2\alpha\beta}]$ has an inverted sign in the identity term, and therefore these variables are ghost like.
The calculation of $c_{1}$ and $c_{2}$ follows easily with this conditions from (2.8) and (2.9), by playing with some specific polarizations. The result is so simple as this
$c_{1}=1,\qquad c_{2}=-1.$ (2.41)
These are the same structure constants as in the Pais-Uhlenbeck oscillator, which will be found in the next section.
The third mode, which is usually a ghost, is gone. However, the second mode became a ghost itself. Nevertheless, its presence can be dealt with the methods described in the next sections, based on the quantization program explained in [1].
3. Quantization formulas for ghosts
Before dealing with Quadratic gravity, consider the prototype of a ghost system, namely the Pais-Uhlenbeck lagrangian [44] for a real scalar mode $\phi$
$L=-\frac{3}{2}\partial_{\mu}\phi\partial^{\mu}\phi-\frac{9\gamma}{2}(\square\phi)^{2}-V(\phi).$
The potential $V(\phi)$ will be quite general in the present discussion. As the theory is of higher order, the procedure of doubling the number of variables can be applied, and there will be two coordinates and two momenta involved instead of one. The coordinates will be chosen as
$Q_{1}=\beta\phi,\qquad Q_{2}=\lambda\dot{\phi}.$ (3.42)
with $\beta$ and $\lambda$ constant factors. In the following the values $\beta=\lambda=1$ can be chosen, but these factors are written in order to specify certain liberty on the choice of coordinates and momenta which do not alter the canonical commutation relations. In fact, the quantization to be employed now may allow complex values for these constants and still give real mean values for these operators [29]. This method, maybe ackward at first sight, has roots from Dirac and Pauli times [6]-[7].
The canonical momenta corresponding to (3.42) can be found by the following formula, valid for a higher order formalism
$P_{1}=\frac{\partial L}{\partial\dot{Q_{1}}}-\frac{d}{dt}\bigg(\frac{\partial L}{\partial\ddot{Q_{1}}}\bigg)=\frac{1}{\beta}[3\dot{\phi}-9\gamma\square\dot{\phi}],$
$P_{2}=\frac{\partial L}{\partial\dot{Q_{2}}}-\frac{d}{dt}\bigg(\frac{\partial L}{\partial\ddot{Q_{2}}}\bigg)=\frac{9\gamma}{\lambda}\square\phi,$ (3.43)
while the Hamiltonian density is found by the standard prescription
$H=P_{1}\dot{Q}_{1}+P_{2}\dot{Q}_{2}-L.$
The above procedure ensures that the hamiltonian equations of motion are equivalent to the Euler-Lagrange ones. It is useful to deduce from the second (3.43) that
$\ddot{\phi}=\Delta\phi-\frac{\lambda}{9\gamma}P_{2}.$
The Hamiltonian density can be calculated explicitly in terms of the canonical coordinates (3.42) and momentum (3.43), the result is
$H=\frac{\beta P_{1}Q_{2}}{\lambda}-\frac{\lambda^{2}P_{2}^{2}}{18\gamma}+\frac{\lambda P_{2}\Delta Q_{1}}{\beta}-\frac{3}{2}\frac{Q_{2}^{2}}{\lambda^{2}}+\frac{3}{2\beta^{2}}(\nabla Q_{1})^{2}+V(Q_{1}).$ (3.44)
Here $\Delta$ is the three dimensional Laplacian.
The equation of motion of the model is
$3\square\phi+9\gamma\square^{2}\phi=\frac{\partial L_{i}}{\partial\phi},$ (3.45)
where the interaction lagrangian $L_{1}$ is due to the potential term $V(\phi)$. In the free case, the last equation shows that
$\phi=\phi_{1}+\phi_{2},$
the first mode is massless and the second has mass $m^{2}=\frac{1}{3\gamma}$. This means that the field can be Fourier expanded as
$\phi(x)=\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}\omega(k)}[b(k)e^{i\omega(k)t-ik\cdot x}+b^{\ast}(k)e^{-i\omega(k)t+ik\cdot x}]$
$\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}\omega_{2}(k)}[c(k)e^{i\omega_{2}(k)t-ik\cdot x}+c^{\ast}(k)e^{-i\omega_{2}(k)t+ik\cdot x}],$
where the corresponding dispersion relation is given by $\omega_{2}(k)=\sqrt{k^{2}+m^{2}}$. Also $\omega(k)=|k|$.
The quantization method requires the expressions of the mode coefficients $b(k)$, $b^{\ast}(k)$, $c(k)$ and $c^{\ast}(k)$ in terms of the canonical variables $P^{i}$ and $Q_{i}$ with $i=1,2$. Then, by assuming the standard equal time commutation relations given by
$Q_{i}(x,t),P_{j}(y,t)]=i\hbar\delta_{ij}\delta(x-y),$ (3.46)
the commutation relations for the quantum operators $b(k)$, $b^{\dagger}(k)$, $c(k)$ and $c^{\dagger}(k)$ can be deduced. This is a $4\times 4$ system of equations which, in principle, may be enough for determining those quantities.
This naive expectation however, may have a problem. It will be shown below, it is not possible to enforce this canonical equal time commutators (3.46) if the coordinates commute
$Q_{i}(x,t),Q_{j}(y,t)]=0.$ (3.47)
In fact, it will be seen that the solution of the $4\times 4$ system is such that $Q_{i}(x,t),Q_{j}(y,t)]\neq 0$. In order to avoid this problem one may be tempted to modify one or more of the relations (3.46). If this approach is applied, it leads to apparent negative norm states.
The presence of negative norm states can be justified as follows. The second of the formulas (3.43) shows that this massless mode does not contribute to $P_{2}$ as $\square\phi_{1}=0$. Furthermore, from the first (3.43) and the definition $\dot{\phi}=\lambda Q_{2}$, it is seen that it does not contribute to $\beta P_{1}-\frac{3Q_{2}}{\lambda}$ either. Therefore, only the massive mode $\phi_{2}$ contributes to these two quantities and, since the massive mode involves only the coefficient $c(k)$ and $c^{\ast}(k)$, the system
$\lambda\beta P_{1}-3Q_{2}=-9\lambda\gamma\square\dot{\phi}_{2},\qquad P_{2}=\frac{9\gamma}{\lambda}\square\phi_{2},$ (3.48)
is a $2\times 2$ one allowing the calculation $c(k)$ and $c^{\ast}(k)$ in terms of the modes of $\lambda\beta P_{1}-3Q_{2}$ and $P_{2}$. This makes the problem more tractable. To solve this $2\times 2$ linear system explicitly, write
$\phi_{2}(x,t)=\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}\omega_{2}(k)}[c(k)e^{i\omega_{2}(k)t-ik\cdot x}+c^{\ast}(k)e^{-i\omega_{2}(k)t+ik\cdot x}].$
By taking into account that $\square\phi_{2}=m^{2}\phi_{2}$ with $m^{2}=\frac{1}{3\gamma}$ the equations (3.48) are converted into
$\lambda P_{2}(x,t)=3\phi_{2}(x,t)=3\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}\omega_{2}(k)}[c(k)e^{i\omega_{2}(k)t-ik\cdot x}+c^{\ast}(k)e^{-i\omega_{2}(k)t+ik\cdot x}],$
$3Q_{2}(x,t)-\lambda\beta P_{1}(x,t)=3\dot{\phi}_{2}=3i\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}}[c(k)e^{i\omega_{2}(k)t-ik\cdot x}-c^{\ast}(k)e^{-i\omega_{2}(k)t+ik\cdot x}].$ (3.49)
This can be inverted to give
$c(k)=\frac{1}{3}\int d^{3}xe^{ikx}[\omega_{2}(k)\lambda P_{2}(x,0)-3iQ_{2}(x,0)+i\lambda\beta P_{1}(x,0)],$
$c^{\dagger}(k)=\frac{1}{3}\int d^{3}xe^{-ikx}[\omega_{2}(k)\lambda P_{2}(x,0)+3iQ_{2}(x,0)-i\lambda\beta P_{1}(x,0)].$ (3.50)
If these expressions are prompted to operators $\widetilde{c}_{k}$ and $\widetilde{c}_{k}^{\dagger}$, by assuming the canonical commutation relations (3.46), it is arrived to
$\widetilde{c}_{k^{\prime}},\widetilde{c}^{\dagger}_{k}]=\frac{2\lambda\omega_{2}(k)}{3}\delta(k-k^{\prime}),$ (3.51)
This is the standard commutation relation, up to a factor $3\lambda^{-1}$. Consider now the remaining operators $b(k)$ and $b^{\dagger}(k)$. Their commutation relation can be determined from (3.42), under the assumption that $Q_{1},Q_{2}]=0$. This definition implies that
$Q_{1}=\beta\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}\omega(k)}[b(k)e^{i\omega(k)t-ik\cdot x}+b^{\ast}(k)e^{-i\omega(k)t+ik\cdot x}]$
$\beta\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}\omega_{2}(k)}[c(k)e^{i\omega_{2}(k)t-ik\cdot x}+c^{\ast}(k)e^{-i\omega_{2}(k)t+ik\cdot x}],$
$Q_{2}=i\lambda\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}}[b(k)e^{i\omega(k)t-ik\cdot x}-b^{\ast}(k)e^{-i\omega(k)t+ik\cdot x}]$
$i\lambda\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}}[c(k)e^{i\omega_{2}(k)t-ik\cdot x}-c^{\ast}(k)e^{-i\omega_{2}(k)t+ik\cdot x}].$ (3.52)
As these two quantities commute, if the unknowns $b_{k}$ and $b^{\ast}(k)$ are prompted to operators $\widetilde{b}_{k}$ and $\widetilde{b}_{k}^{\dagger}$ the commutation relation $\widetilde{b}_{k},\widetilde{b}_{k}^{\dagger}]$ has to cancel the terms coming from $c_{k},c_{k}^{\dagger}]$. This leads to
$b_{k},b_{k^{\prime}}^{\dagger}]=-\frac{2\lambda\omega_{2}(k)}{3}\delta(k-k^{\prime}).$ (3.53)
The minus sign in the last expression is the key result. It implies that one of the two pairs of creation/annihilation operators has a non standard 1-1 sign in the commutation relation in order to ensure that $Q_{1},Q_{2}]=0$. This non standard sign is unavoidable. The last commutation relation leads to an expression of $Q_{1},P_{1}]$ at equal times to have a minus sign in their definition, that is
$Q_{1}(x,t),P_{1}(y,t)]=-i\hbar\delta(x-y).$
One may declare a non standard minus sign in $Q_{2},P_{2}]$ instead, and the commutation $Q_{1},P_{1}]$ will be standard. Nevertheless, it is not possible to impose (3.46) for both pairs of canonical variables, at least if $Q_{1},Q_{2}]=0$.
Note that (3.52) and (3.49) imply that
$P_{1}=\frac{3}{\lambda\beta}i\lambda\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}}[b(k)e^{i\omega(k)t-ik\cdot x}-b^{\ast}(k)e^{-i\omega(k)t+ik\cdot x}].$ (3.54)
As $P_{2}$ in (3.49) only involves only $c(k)$ and $c^{\ast}(k)$, the wrong or correct sign with one of the commutators will not enter in conflict with the condition $P_{1},P_{2}]=0$, it only affects the commutator $Q_{1},Q_{2}]$.
The behavior of the energy levels of the system requires the analysis of its Hamiltonian. Its classical version is given explicitly in terms of $\phi$ by
$H=[3\dot{\phi}-9\gamma\square\dot{\phi}]\dot{\phi}+9\gamma\square\phi\ddot{\phi}+\frac{3}{2}\partial_{\mu}\phi\partial^{\mu}\phi+\frac{9\gamma}{2}(\square\phi)^{2}+V(\phi).$ (3.55)
This expression is obtained after some algebra from (3.46) by employing the formulas (3.43) in order to express everything in terms of $\phi$.
By decomposing $\phi=\phi_{1}+\phi_{2}$ and by taking into account their dispersion relations, it can be seen after some calculation, to be presented now, that the mixed terms involving $\phi_{1}$ and $\phi_{2}$ cancel out, and the resulting Hamiltonian density is simply the sum of two free Klein-Gordon ones corresponding to both modes.
The calculation leading to this conclusion goes as follows. After introducing the decomposition $\phi=\phi_{1}+\phi_{2}$ into (3.55) there appear terms that only involve $\phi_{1}$, other that involve $\phi_{2}$ and mixed terms. The mixed ones are
$H_{12}=[3\dot{\phi}_{1}-9\gamma\square\dot{\phi}_{1}]\dot{\phi}_{2}+9\gamma\square\phi_{1}\ddot{\phi}_{2}+3\partial_{\mu}\phi_{1}\partial^{\mu}\phi_{2}+9\gamma\square\phi_{1}\square\phi_{2}$
$3\dot{\phi}_{2}-9\gamma\square\dot{\phi}_{2}]\dot{\phi}_{1}+9\gamma\square\phi_{2}\ddot{\phi}_{1}.$
From the dispersion relation of the modes, it is clear that $\square\phi_{1}=0$ and $3\gamma\square\phi_{2}=\phi_{2}$. This leads to several simplifications, and the last expression becomes
$H_{12}=3\dot{\phi}_{1}\dot{\phi}_{2}+3\partial_{\mu}\phi_{1}\partial^{\mu}\phi_{2}+3\phi_{2}\ddot{\phi}_{1}.$
Collecting the first and the last term as a total derivative, it follows that
$H_{12}=3\partial_{t}(\dot{\phi}_{1}\phi_{2})-3\dot{\phi}_{1}\dot{\phi}_{2}+3\nabla\phi_{1}\cdot\nabla\phi_{2}.$
This can be equivalently written as
$H_{12}=3\phi_{2}\ddot{\phi}_{1}+3\nabla\phi_{1}\cdot\nabla\phi_{2}.$
This is the mixed term of the hamiltonian density. The full hamiltonian is integrated over the spatial volume $\int_{V}d^{3}x$. Inside the spatial volume integral, the last term is equivalent by parts integration to
$H_{12}=3\phi_{2}\ddot{\phi}_{1}-3\phi_{2}\Delta\phi_{1}.$
Since $\ddot{\phi}_{1}\sim-\omega^{2}\phi_{1}$ while $\Delta\phi_{1}\sim-k^{2}\phi_{1}$ inside the mode expansion, the dispersion relation for $\phi_{1}$ $\omega^{2}=k^{2}$ implies that the contribution of these terms cancel. This means that one may set the mixing $H_{12}=0$.
The resulting hamiltonian density is therefore composed only by the decoupled terms involving only $\phi_{1}$ or only $\phi_{2}$. The insertion of the expression $\phi=\phi_{1}+\phi_{2}$ into (3.55) throws the following
$H=\frac{3}{2}\dot{\phi}^{2}_{1}+\frac{3}{2}(\nabla\phi_{1})^{2}+3\phi_{2}\ddot{\phi}_{2}-\frac{3}{2}\dot{\phi}^{2}_{1}+\frac{3}{2}(\nabla\phi_{2})^{2}+\frac{3m^{2}}{2}(\phi_{2})^{2}.$
Inside the Fourier expansion $\ddot{\phi}_{2}=\dot{\phi}_{2}^{2}\sim-\omega_{2}^{2}\phi_{2}$, thus the last expression is equivalent to
$H=\frac{3}{2}\dot{\phi}^{2}_{1}+\frac{3}{2}(\nabla\phi_{1})^{2}+\frac{3}{2}\dot{\phi}^{2}_{2}+\frac{3}{2}(\nabla\phi_{2})^{2}+\frac{3m^{2}}{2}(\phi_{2})^{2}+V(\phi_{1},\phi_{2}).$
This clearly the sum of two free Klein-Gordon hamiltonian densities, for a massless and a massive particle, with an interaction $V(\phi_{1},\phi_{2})$. By omitting the zero point energy, the free part of the hamiltonian written in terms of the modes is given by
$H_{f}=\int d^{3}k[\omega(k)b^{\dagger}(k)b(k)+\omega_{2}(k)c^{\dagger}(k)c(k)].$ (3.56)
This expression would not raise any alarm if the creation and annihilation operators were standard. However, it was shown above that one of them may deviate from the standard commutation relation. Therefore it is needed to understand if this Hamiltonian is bounded or instead unbounded from below, if these non standard commutators are employed.
The problems described above may be understood in terms of path integrals. The Hamiltonian (3.46) depends only linearly on $P_{1}$. This raises a problem if one tries to quantize the model by the method of path integrals. If the coefficients $\beta$ and $\lambda$ are chosen equal to one, then the euclidean path version of this theory wlll be divergent, since the integration on the variable $P_{1}$ yields an infinite result. This divergence is clear, since the integral in $P_{1}$ involves a real exponential and the integrand will be unbounded. The fact that the Hamiltonian is linear in $P_{1}$ seems then problematic. This is not the case for other choices of these constants, for instance imaginary ones. These choices do not affect the canonical commutation relations $[P,Q]$ as they multiply $Q$ in inverse way as for $P$. The point is to understand the consequences of declaring that $\lambda$ or $\beta$ are imaginary  [^1] Related work is the reference [36], where apparently a self-consistent way to quantize these systems has been presented..
3.1 A toy quantum mechanical model
At this point it is instructive to study the non standard harmonic oscillator
$H=\hbar\omega u_{c}u_{a},\qquad[u_{a},u_{c}]=-1.$
The non standard nature of this oscillator is due to the factor 1-1 in the commutator, while in standard QM it is $1$. This situation is the QM analogous of the second term of (3.56). One may interpret these operators in terms of the standard creation and annihilation operators as
$u_{a}=ia,\qquad u_{c}=ia^{\dagger}$
These operators are not adjoint each other. In fact $u_{a}^{\dagger}=-u_{c}$. The standard states of an harmonic oscillator $|n>$ can be employed in order to find
$n|H|n>=-\hbar\omega n.$
The energy seems to be not bounded from below. However, the well known relation
$a=\frac{\hat{x}+i\hat{p}}{\sqrt{2}}.$
suggests that the definition $u_{a}=ia$ is equivalent to the operator redefinition $\hat{p}\to\hat{P}=i\hat{p}$ and $\hat{x}\to\hat{X}=i\hat{x}$. The Hamiltonian $H$ is quadratic in these quantities and implies the redefinition $H\to-H$.
Therefore, it is reasonable. to postulate that the mean values of these quantities are
$\overline{p}=-i<\hat{P}>,\qquad\overline{x}=-i<\hat{X}>,\qquad\overline{E}=-<H>.$
Already this is in concordance with the continuation procedure in [1]. Also for example
$\overline{p}^{n}=(-i)^{n}<\hat{P}^{n}>,\qquad\overline{x}^{n}=(-i)^{n}<\hat{X}^{n}>$
$\overline{p^{n}x^{m}}=(-i)^{n+m}<\hat{P}^{n}\hat{X}^{m}>$
For a generic observable $O(P,X)$ constructed in powers of $\hat{X}$ and $\hat{P}$ these rules give the corresponding mean values. Calculate the mean value <O(P,X)><O(P,X)> and analitically continue to values $P=\to-iP$ and $X\to-iX$. The energy mean values are now positive, and the impulse and coordinates mean values are real..
Another subtle detail is related to probabilities. The basis $|n>$ employed above is the standard one. Then <n|n>=1<n|n>=1, which is perfectly fine. However, if this state is expressed in terms of the creation operators, then
$|n>\neq\frac{1}{\sqrt{n!}}u_{c}^{n}|0>,$
In other words, the standard expression in QM does not hold here. Instead
$|n>=\frac{(-i)^{n}}{\sqrt{n!}}u_{c}^{n}|0>.$ (3.57)
The reason for this formula is simple. The standard oscillator states $|n>$ are created in terms of $a^{\dagger}=-iu_{c}$, and this leads to the conclusion. This distinction will be important in the derivation of the LSZ formulas. In fact, consider the scalar field of the previous section $\phi=\phi_{1}+\phi_{2}$. The interaction potential may generate terms of the form $\phi_{1}^{2}\phi_{2}^{3}$. If one calculates $V(\phi)>$, then terms will appear such as
$\phi_{1}^{3}\phi_{2}^{3}>$
As the ghost part $\phi_{2}$ is constructed with creation and annihilation operators which has an imaginary factor $i$, this will lead to an imaginary value of the last expression. On the other hand, it is reasonable to expect the mean value of a real scalar field to be real. Therefore, the above prescription is consistent if the mean values are calculated, for example, as follows
$\overline{\phi_{1}^{3}\phi_{2}^{3}}=(-i)^{3}<\phi_{1}^{3}\phi_{2}^{3}>.$ (3.58)
In general, in order to calculate mean values of a given observable, one must make the replacement $\phi=\phi_{1}+\phi_{2}\to\phi_{1}-i\phi_{2}$, with the 2 index denoting the ghost variable.
It seems that the above prescription can be expressed as follows. The second quantization of the ghost corresponds to skew hermitian operators of the coordinates and impulses. The fundamental postulates of QM states that the operator corresponding to an observable $Q$ should be hermitian. In the present model, the hermiticity is spoiled by an i-i unit in the creation and annihilation operators. For the Pais-Uhlenbeck oscillator, it is $P_{2}$ and $Q_{2}$ which have this unwanted factor. The above procedure corresponds to throwing away the $i$ factor that spoils hermiticity in all calculations. That is, to go from $P_{2}$ and $Q_{2}$ to $iP_{2}$ and $iQ_{2}$. This is in agreement with the postulates of Quantum Mechanics presented in the Solvay Congress in 1927.
The prescription of changing $P_{2}$ and $Q_{2}$ to $iP_{2}$ and $iQ_{2}$ in the Hamiltonian (3.56) changes the unwanted negative sign in the ghost contribution, and the energy becomes bounded from below. This is the desired feature. Also, the fact that the states employed are the standard one do not enter in conflict with negative probabilities.
3.2 The path integral formulation
The path integral formulation of Quantum Mechanics in the presence of both standard $x$ and ghost like $X$ variables is obtained by considering the transition amplitude [13], [4]-[29], [1]-[2]
$x_{f},X_{f},t_{f}|x_{i},X_{i},t_{i}>=\prod_{l=1}^{N}\int[dx_{l}][dX_{l}]<x_{f},X_{f},t_{f}|x_{N},X_{N},t_{N}>$...
$x_{1},X_{1},t_{1}|x_{i},X_{i},t_{i}>,$
this results follows by successively inserting identities
$1=\int|x,X,t><x,X,t|dxdX.$
In addition
$x_{l+1},X_{l+1},t_{l+1}|x_{l},X_{l},t_{l}>=\int<x_{l+1},X_{l+1},t_{l+1}|p_{l+1},P_{l+1},t_{l+1}>$
$p_{l+1},P_{l+1},t_{l+1}|x_{l},X_{l},t_{l}>dp_{l}dP_{l}.$
From the expression of the evolution at small $\delta t_{l}=t_{l+1}-t_{l}$ given in terms of the system‘s Hamiltonian by
$p_{l+1},P_{l+1},t_{l+1}|x_{l},X_{l},t_{l}>=<p_{l+1},P_{l+1},t_{l}|e^{-iH\delta t}|x_{l},X_{l},t_{l}>$
$\frac{1}{(2\pi)^{m}}\exp\{-iH(p_{l},x_{l},P_{l},X_{l})\delta t_{l}-i(p_{l}x_{l}+P_{l}X_{l})\},$
where $m$ is the number of degrees of freedom and $H(p_{l},x_{l},P_{l},X_{l})$ is the mean value
$H_{a}(p_{l},x_{l},P_{l},X_{l})=\frac{<p_{l},P_{l},t_{l}|H(p_{l},x_{l},P_{l},X_{l})|x_{l},X_{l},t_{l}>}{<p_{l},P_{l},t_{l}|x_{l},X_{l},t_{l}>},$
An important point is that the above mean value has to be constructed taking into account that the eigenvalues of $P_{i}$ and $X_{i}$ are imaginary. Therefore, at the classical level it corresponds to put imaginary values $iP_{l}$ and $iX_{l}$. The minus in the momentum is due to the fact that <P|<P| is on the left when calculating the mean value, and that $P$ is skew hermitian. Collecting all these results, it follows that
$x_{f},X_{f},t_{f}|x_{i},X_{i},t_{i}>=\int[Dx][DX][Dp][DP]\exp\bigg\{i\int_{t_{i}}^{t_{f}}dt[p\dot{x}+P\dot{X}-H_{a}(p,x,P,X)]\bigg\}$
The last expression imitates usual path integral in QM, however with the subtle detail that the Hamiltonian is continued to imaginary values of the ghost variables. This may generate real divergent exponentials, and the integral may not exist. This existence issue has to be analyzed model by model. Fortunately, in the present case, this unpleasant feature does not take place [1]-[2], [4]-[29].
In order to see that the path integral does not have the apparent divergencies problem, note that the Hamiltonian (3.46), which it written again here by convenience after choosing $\beta=\lambda=1$
$H=P_{1}Q_{2}-\frac{P_{2}^{2}}{18\gamma}+P_{2}\Delta Q_{1}-\frac{3}{2}Q_{2}^{2}+\frac{3}{2}(\nabla Q_{1})^{2}+V(Q_{1})$ (3.59)
only depends on $P_{1}$ by a linear term, which in the path integral may lead to a divergent exponential. However, under the prescription of continue the values of $Q_{2}\to iQ_{2}$ and $P_{2}\to-iP_{2}$ it becomes
$H=iP_{1}Q_{2}+\frac{P_{2}^{2}}{18\gamma}-iP_{2}\Delta Q_{1}+\frac{3}{2}Q_{2}^{2}+\frac{3}{2}(\nabla Q_{1})^{2}+V(Q_{1}).$ (3.60)
The dangerous term now is converted in a harmless imaginary term $iP_{1}Q_{2}$ whose exponentiation and integration does not lead to divergent results. To see this note that the euclidean path integral is
$Z_{e}(0)=\int DP^{1}D\phi_{1}DP^{2}D\phi_{2}\exp\bigg\{-\int[iP_{1}\dot{Q}_{1}+iP_{2}\dot{Q}_{2}-H]dVd\tau\bigg\}$
$\int DP^{1}D\phi_{1}DP^{2}D\phi_{2}$
$\exp\bigg\{-\int[iP_{1}\dot{\phi}_{1}+iP_{2}\dot{\phi}_{2}-iP_{1}Q_{2}-\frac{P_{2}^{2}}{18\gamma}+iP_{2}\Delta Q_{1}-\frac{3}{2}Q_{2}^{2}-\frac{3}{2}(\nabla Q_{1})^{2}]dVd\tau\bigg\}$
The integration over $P_{1}$ leads to a Dirac delta that enforces $Q_{2}=-Q^{\prime}_{1}$, where the prime indicates derivative with respect to the euclidean time. The path integral now is related to $D\dot{Q}_{1}$ due to this identification, but a change to the variable $DQ_{1}$ may be done by multiplying the integration measure by $\prod_{x_{E^{\prime}}}\Delta\tau^{-1}$. The integral over $P_{2}$ is a Gaussian integral after setting $Q_{2}=-Q^{\prime}_{1}$. The general formula
$\int d^{N}xe^{x^{T}Ax+bx}=\frac{\pi^{\frac{N}{2}}}{\sqrt{\det A}}e^{-\frac{1}{4}b^{T}Ab},$
can be employed to make this integral. After some calculation, it is found that the result with currents $J$ turned on given by
$Z_{e}(J)=\int D\phi\exp\bigg\{-\int L_{e}dVd\tau+\int J\phi dVd\tau\bigg\}.$ (3.61)
Here $L_{e}$ is the euclidean version of the model corresponding to the action
$S=\int d\tau\int d^{3}x\frac{3}{2}\partial_{\mu}\phi\partial_{\mu}\phi+\frac{9\gamma}{2}(\partial_{\mu}\partial_{\mu}\phi)^{2}+V(\phi).$
Therefore, the apparent divergence which can invalidate the model is avoided with this approach. The correlation functions can be calculated in the euclidean formalism and analitically continued to the Lorenzian setting, by going from imaginary $\tau$ to real $t=i\tau$ times.
A further comment about the literature about the subject is mandatory. The prescription given above, which requires to continue the values of the coordinates and momentum to imaginary values, is inspired in some deep works, for instance [7]-[30]. I believe that at the end it is equivalent to the prescription of [1]. Although the procedure may look naive, it has it roots in works of Dirac and Pauli, and was extensively studied also in [13]. The main difference with these approaches is to put the imaginary unit $i$ not in the creation and annihilation operators, but in the states themselves. In other words, these references work with modified states $|n>$ whose norm is not insured to be positive due to the inclusion of $i$ unit. The norm of these states $\eta_{nm}=<n|m>$ is has negative entries. However, the mixed index metric $\eta_{n}^{m}=\delta_{n}^{m}$ is positive definite. These references distinguish between covariant, contravariant and mixed Hilbert space metrics. The mixed one is adapted to describe probabilities [13], [4]-[29], [1]. The mean values can be given by taking into account the covariant or contravariant metric. This is the formalism employed in [42]. Our conjecture is that the prescription presented here will lead, at least in these models, the same or analogous results than those pioneer references.
4. LSZ reduction for ghost variables
Given the previous discussion, consider again the $S$ matrix for a given process involving the scalar $\phi$. The action on the creation operator for a state $b_{in}^{\dagger}(k)$ of momentum $k$ is defined by
$S^{\dagger}b^{\dagger}_{in}(k)S=(\sum_{\text{all states}}|r,\text{out}><r,\text{in}|)(\sum_{\text{all states}}c_{k}\sqrt{n(k)+1}|h+k,\text{in}><h,\text{out}|)$
$\sum_{\text{all states}}c_{k}\sqrt{n(k)+1}|h+k,\text{out}><h,\text{out}|.$ (4.62)
Here $n(k)$ are the number of quanta of momentum $k$ that were already present in the state $h$, and $h+k$ denotes the state with one of such quanta added. A constant $c_{k}$ is included, in order to anticipate that these operators may not be standard.
The following analysis repeats the standard one adapted to a theory of second order to one of fourth order, with ghost included. The idea is to understand in which part of the deduction the fact that the creation or annihilation operators are non standard may affect the calculations. Therefore, although the analysis may look straightforward, it is necessary to work out all the details in order to be sure what are the consequences, if any, of the non standard creation and annihilation nature of the operators.
Clearly, the last expression represents $b^{\dagger}_{out}(k)$, which leads to the conclusion that
$S^{\dagger}b^{\dagger}_{in}(k)S=b^{\dagger}_{out}(k),\qquad S^{\dagger}b_{in}(k)S=b_{out}(k),$
the second relation is completely analogous to the first one. The value of the constant $c_{k}$ parameterizing the non standard nature of the operator is irrelevant in order to obtain this conclusion. The same relation follows for $c(k)$ and $c^{\dagger}(k)$. Therefore, it is deduced that
$S^{\dagger}\phi_{in}(k)S=\phi_{out}(k),$ (4.63)
regardless the non standard commutation of the creation and annihilation operators.
The use of the standard Fock space ensures that, at this point, the matrix $S$ in (4.62) is unitary.
Before deducing the LSZ formula, it is mandatory to deduce some classical relations for the fields in the model. The first is related to the equations of motion of $\phi(x)$ is given by
$\hat{O}\phi_{1}=c\frac{\partial L_{int}}{\partial\phi},$ (4.64)
where $L_{int}$ contains the infinite vertices involving the interaction of $\phi_{1}$ with the rest of the fields. The operator $\hat{O}$, corresponding to the equations of motion of the scalar field of the previous section, is a quartic operator
$\hat{O}=\square^{2}-\frac{1}{3\gamma}\square,$
up to a constant. That is the reason for adding a constant $c$ in the interaction term. The following discussion follows closely the methods of [154]. The kernel of the last operator, that is, a Green function for $\hat{O}$ is given by
$\hat{O}_{y}G(y-x)=\delta(y-x).$
It is clear that
$G(x-y)=\frac{1}{\hat{O}_{y}}=\frac{1}{\square^{2}-\frac{1}{3\gamma}\square}=\bigg(\frac{1}{\square}-\frac{1}{\square-\frac{1}{3\gamma}}\bigg)\frac{1}{3\gamma}.$
Therefore the Green function G(xy)G(x-y) is given by
$G(x-y)=\frac{1}{3\gamma}[G^{m=0}(x-y)-G^{m}(x-y)],$
being $G^{m}(x-y)$ any of the standard Green functions for a massive scalar field. There are plenty of such Green functions since the addition of a homogeneous solution of the wave operator leads to a new Green function. In terms of any of these Geen kernel, the following integral relation for $\phi_{1}$
$\int[\phi_{1}(y)\hat{O}_{y}G(y-x)-G(y-x)\hat{O}_{y}\phi_{1}(y)]d^{4}y=\phi_{1}(x)-c\int G(y-x)\frac{\partial L_{int}}{\partial\phi(y)}d^{4}y,$ (4.65)
is found. By assuming that $\phi_{1}$ vanish fast enough at the spatial infinite, it is concluded that the spatial part of the left hand side is related to two types of integrals. The first is related to the laplacian $\Delta_{y}$ in three dimensions, appearing in $\square_{y}$,. This contribution is
$\int[\phi_{1}(y)\Delta_{y}G(y-x)-G(y-x)\Delta_{y}\phi_{1}(y)]d^{3}ydy_{0}$
$\int[\phi_{1}(y)\nabla_{y}G(y-x)-G(y-x)\nabla_{y}\phi_{1}(y)]dS_{y}dy_{0}.$ (4.66)
The right hand side is obtained from the left by noticing that the left integrand is the gradient of $\phi_{1}(y)\nabla_{y}G(y-x)-G(y-x)\nabla_{y}\phi_{1}(y)$. If everything vanish in the infinite surface $S_{y}$ this gives no contributions. The other spatial contribution, due to the terms with $\square_{y}^{2}$, are given by
$\int[\phi_{1}(y)\Delta_{y}\Delta_{y}G(y-x)-G(y-x)\Delta_{y}\Delta_{y}\phi_{1}(y)]d^{3}ydy_{0}=0,$
after integration by parts. There are also terms in $\square^{2}_{y}$ mixing spatial and time derivatives. These are of the form
$\int[\phi_{1}(y)\Delta_{y}\partial^{2}_{y_{0}}G(y-x)-G(y-x)\Delta_{y}\partial^{2}_{y_{0}}\phi_{1}(y)]d^{3}ydy_{0}$
$\int\partial_{y_{0}}[\phi_{1}(y)\Delta_{y}\partial_{y_{0}}G(y-x)-G(y-x)\Delta_{y}\partial_{y_{0}}\phi_{1}(y)]d^{3}ydy_{0}$
$\int[\partial_{y_{0}}\phi_{1}(y)\Delta_{y}\partial_{y_{0}}G(y-x)-\partial_{y_{0}}G(y-x)\Delta_{y}\partial_{y_{0}}\phi_{1}(y)]d^{3}ydy_{0}.$
The last two terms cancel by a same argument as (4.66). Therefore the last equality can be worked out as follows
$\int[\phi_{1}(y)\Delta_{y}\partial^{2}_{y_{0}}G(y-x)-G(y-x)\Delta_{y}\partial^{2}_{y_{0}}\phi_{1}(y)]d^{3}ydy_{0}$
$\int\partial_{y_{0}}[\phi_{1}(y)\Delta_{y}\partial_{y_{0}}G(y-x)-G(y-x)\Delta_{y}\partial_{y_{0}}\phi_{1}(y)]d^{3}ydy_{0}$
$\int\partial_{y_{0}}[\Delta_{y}\phi_{1}(y)\partial_{y_{0}}G(y-x)-G(y-x)\Delta_{y}\partial_{y_{0}}\phi_{1}(y)]d^{3}ydy_{0}.$ (4.67)
The unique spatial contribution is then (4.67). The pure time components that come from $\square_{y}$ are of the form
$\int[\phi_{1}(y)\partial^{2}_{y_{0}}G(y-x)-G(y-x)\partial^{2}_{y_{0}}\phi_{1}(y)]d^{4}y$
$\int\partial_{y_{0}}[\phi_{1}(y)\partial_{y_{0}}G(y-x)-G(y-x)\partial_{y_{0}}\phi_{1}(y)]d^{4}y$
$\bigg(\int_{y^{-}_{0}}-\int_{y^{+}_{0}}\bigg)G(y-x)\overleftrightarrow{\partial}_{y_{0}}\phi_{1}(y)d^{3}y.$ (4.68)
The first two integrals are three dimensional, the first is done at an spatial surface corresponding to the time $y_{0}^{-}$at the past of $x_{0}$ and the second at another time $y_{0}^{+}$ at the future of $x_{0}$. The last integral is four dimensional, and is performed on the $4$-volume delimited by these two three dimensional surfaces. The last pure time contribution is due to the time part of $\square_{y}^{2}$. It is, and can be worked out as
$\int[\phi_{1}(y)\partial^{4}_{y_{0}}G(y-x)-G(y-x)\partial^{4}_{y_{0}}\phi_{1}(y)]d^{4}y$
$\int\partial_{y_{0}}[G(y-x)\partial^{3}_{y_{0}}\phi_{1}(y)-\phi_{1}(y)\partial^{3}_{y_{0}}G(y-x)]d^{4}y$
$\int\partial_{y_{0}}[\partial^{2}_{y_{0}}G(y-x)\partial_{y_{0}}\phi_{1}(y)-\partial^{2}_{y_{0}}\phi_{1}(y)\partial_{y_{0}}G(y-x)]d^{4}y$
$\bigg(\int_{y^{-}_{0}}-\int_{y^{+}_{0}}\bigg)[G(y-x)\overleftrightarrow{\partial}_{y_{0}}\partial_{y_{0}}^{2}\phi_{1}(y)+\partial_{y_{0}}^{2}G(y-x)\overleftrightarrow{\partial}_{y_{0}}\phi_{1}(y)]d^{3}y.$ (4.69)
By collecting all the above formulas (4.66)-(4.69), the expression (4.65)t becomes
$\phi_{1}(x)=-\bigg(\int_{y^{-}_{0}}-\int_{y^{+}_{0}}\bigg)G(y-x)\overleftrightarrow{\partial}_{y_{0}}\phi_{1}(y)d^{3}y$
$\frac{1}{3\gamma}\bigg(\int_{y^{-}_{0}}-\int_{y^{+}_{0}}\bigg)[G(y-x)\overleftrightarrow{\partial}_{y_{0}}\partial_{y_{0}}^{2}\phi_{1}(y)+\partial_{y_{0}}^{2}G(y-x)\overleftrightarrow{\partial}_{y_{0}}\phi_{1}(y)]d^{3}y$
$c\int_{y_{0}^{-}}^{y_{0}^{+}}G(y-x)\frac{\partial L_{int}}{\partial\phi(y)}d^{4}y.$
At the present point G(yx)G(y-x) can be any Green function, defined by any possible boundary condition. Particularly useful examples are the advanced $\Delta_{a}(x)$ or retarded $\Delta_{r}(x)$ Green functions, which vanish for $x_{0}>0$ and $x_{0}<0$, respectively, and vanish outside the light cone. If these functions are employed, then the two spatial integrals in the last formulas are reduced into one. For instance
$\phi_{1}(x)=\int_{y^{-}_{0}}\Delta_{r}(y-x)\overleftrightarrow{\partial}_{y_{0}}\phi_{1}(y)d^{3}y$
$c\int_{y^{-}_{0}}[\Delta_{r}(y-x)\overleftrightarrow{\partial}_{y_{0}}\partial_{y_{0}}^{2}\phi_{1}(y)+\partial_{y_{0}}^{2}\Delta_{r}(y-x)\overleftrightarrow{\partial}_{y_{0}}\phi_{1}(y)]d^{3}y+c\int\Delta_{r}(y-x)\frac{\partial L_{int}}{\partial\phi(y)}d^{4}y.$
The first term in the last formula has a simple interpretation. This term is such that
$\hat{O}_{x}\int_{y^{-}_{0}}\Delta_{r}(y-x)\overleftrightarrow{\partial}_{y_{0}}\phi_{1}(y)d^{3}y-\hat{O}_{x}\int_{y^{-}_{0}}[\Delta_{r}(y-x)\overleftrightarrow{\partial}_{y_{0}}\partial_{y_{0}}^{2}\phi_{1}(y)+\partial_{y_{0}}^{2}\Delta_{r}(y-x)\overleftrightarrow{\partial}_{y_{0}}\phi_{1}(y)]d^{3}y$
$\int_{y^{-}_{0}}\delta(y-x)\overleftrightarrow{\partial}_{y_{0}}\phi_{1}(y)d^{3}y-\int_{y^{-}_{0}}[\delta(y-x)\overleftrightarrow{\partial}_{y_{0}}\partial_{y_{0}}^{2}\phi_{1}(y)+\partial_{y_{0}}^{2}\delta(y-x)\overleftrightarrow{\partial}_{y_{0}}\phi_{1}(y)]d^{3}y$
In the limit $y_{0}^{-}\to$-$\infty$ the last integral vanishes if $x_{0}$ is fixed, since this point will be always at the future of $y_{0}$. Therefore this particular quantity is killed by the action of $\hat{O}_{x}$ and may be identified as a free field, the incoming field. The previous formula, by taking into account the equations of motion (4.64), may be expressed after taking this limit as
$\phi_{1}(x)=\phi_{1in}(x)+c\int\Delta_{r}(y-x)\frac{\partial L_{int}}{\partial\phi(y)}d^{4}y,$
or by use of the equations of motion
$\phi_{1}(x)=\phi_{1in}(x)+c\int\Delta_{r}(x-y)\hat{O}_{y}\phi_{1}(y)d^{4}y.$ (4.70)
Here, the second integral can be done in the full space time volume, as the past time is taken to -$\infty$ and the retarded Green function vanishes at the future.
The last two formulas make sense, as this field is determined by time derivatives up to order three, consistent with a four order equation of motion.
Analogously
$\phi_{1}(x)=\phi_{1iut}(x)+c\int\Delta_{a}(y-x)\frac{\partial L_{int}}{\partial\phi(y)}d^{4}y,$
or
$\phi_{1}(x)=\phi_{1out}(x)+c\int\Delta_{a}(x-y)\hat{O}_{y}\phi_{1}(y)d^{4}y.$ (4.71)
All the above discussion is purely classical, at the moment this solution has not been prompted to operators.
Consider now the operator version of these classical identities. From the mode expansion
$\phi_{in}=\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}\omega(k)}[b(k)e^{i\omega(k)t-ik\cdot x}+b^{\ast}(k)e^{-i\omega(k)t+ik\cdot x}]$
$\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}\omega_{2}(k)}[c(k)e^{i\omega_{2}(k)t-ik\cdot x}+c^{\ast}(k)e^{-i\omega_{2}(k)t+ik\cdot x}],$
it is seen that, due to the non standard oscillator algebra of $b(k)$ and $b^{\dagger}(k)$
$\phi_{1in}(x),\phi_{1in}(y)]=\int\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}\omega_{1}(k)}\frac{d^{3}l}{(2\pi)^{\frac{3}{2}}\omega_{1}(l)}\bigg[[b(k),b^{\dagger}(l)]e^{-i\omega(k)x_{0}+ik\cdot x}e^{i\omega(l)y_{0}-il\cdot y}$
$b(l),b^{\dagger}(k)]e^{i\omega(k)x_{0}-ik\cdot x}e^{-i\omega(l)y_{0}+il\cdot y}\bigg].$
$\int\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}\omega_{2}(k)}\frac{d^{3}l}{(2\pi)^{\frac{3}{2}}\omega_{2}(l)}\bigg[[c(k),c^{\dagger}(l)]e^{-i\omega(k)x_{0}+ik\cdot x}e^{i\omega(l)y_{0}-il\cdot y}$
$c(l),c^{\dagger}(k)]e^{i\omega(k)x_{0}-ik\cdot x}e^{-i\omega(l)y_{0}+il\cdot y}\bigg].$
$i\int\frac{d^{3}k}{(\pi)^{\frac{3}{2}}\omega_{1}(k)}e^{ik\cdot(x-y)}\sin(\omega_{1}(k)(x_{0}-y_{0}))-i\int\frac{d^{3}k}{(\pi)^{\frac{3}{2}}\omega_{2}(k)}e^{ik\cdot(x-y)}\sin(\omega_{2}(k)(x_{0}-y_{0})).$
This is a linear combination of two propagators for a massless and massive particle, but one of them has a non standard minus sign in front. Each of these propagators are known to be written as a difference of a retarded and advanced Green functions
$\Delta^{m}(x-y)=\Delta^{m}_{a}(x-y)-\Delta^{m}_{r}(x-y),$ (4.72)
and the analogous formula holds for the massless case, with dispersion $\omega_{1}(k)$. In other words, by denoting the last commutator as $i\Delta(x-y)$ it is clear that
$\phi_{1in}(x),\phi_{1in}(y)]=<0|[\phi_{1in}(x),\phi_{1in}(y)]|0>=i\Delta(x-y)=i\Delta_{a}(x-y)-i\Delta_{r}(x-y)$
$i\Delta^{m=0}_{a}(x-y)-i\Delta^{m=0}_{r}(x-y)-i\Delta^{m}_{a}(x-y)+i\Delta^{m}_{r}(x-y),$
two propagators are due to the massless mode and two due to the massive mode. The first identity is due to the fact that the commutator is a c-number. The last expression allows to infer that
$0|[\phi_{1in}(x),\phi_{1in}(y)]|0>=<0|[\phi_{1}(x),\phi_{1}(y)]|0>.$ (4.73)
This is seen as follows. The last identity, deduced from (4.70) is true up to a term proportional to
$0|[\phi_{in}(x),\hat{O}_{y}\phi(y)]|0>.$
Therefore, the identity will be valid if this term is zero. It is, since
$0|\hat{O}_{y}\phi_{1}(y)|p>=\hat{O}_{y}e^{ipy}<0|\phi(0)|p>=p^{2}(1-(3\gamma)^{-1}p^{2})e^{ipy}<0|\phi(0)|p>=0.$
The last identity is due to the fact that the modes satisfy the dispersion relation $p^{2}(1-(3\gamma)^{-1}p^{2})=0$. Due to this property the insertion of the identity $\int dp|p><p|$ in this extra piece gives vanishing result and the desired identity is proved. In addition, the value of $\phi_{1}(x),\phi_{1}(y)]$ is known to be a c-number [154] and therefore (4.73) may be replaced by
$\phi_{1in}(x),\phi_{1in}(y)]=[\phi_{1}(x),\phi_{1}(y)]=i\Delta(x-y),$ (4.74)
in other words, the vacuum expectation value may be deleted.
By collecting all the above information, consider the fundamental operator for scattering in QFT namely
$Z(J)=<0|I(J)|0>,\qquad I(J)=Te^{i\int\phi(x)J(x)dx},$
the last quantity satisfies
$i\frac{\delta I(J)}{\delta J(x)}=T(\phi(x)I(J)).$
The last formula combined with formulas (4.70) and (4.71) leads to the following formulas
$i\frac{\delta I(J)}{\delta J(x)}=I(x)\phi_{1in}(x)+i\int\Delta_{r}(x-y)\hat{O}_{y}\frac{\delta I(J)}{\delta J(y)}d^{4}y,$ (4.75)
$i\frac{\delta I(J)}{\delta J(x)}=\phi_{1out}(x)I(x)+i\int\Delta_{a}(x-y)\hat{O}_{y}\frac{\delta I(J)}{\delta J(y)}d^{4}y.$ (4.76)
From the two equations (4.75) and (4.76) it is arrived to
$\phi_{1out}(x)I(x)-\phi_{1in}(x)I(x)=-i\int\Delta(x-y)\hat{O}_{y}\frac{\delta I(J)}{\delta J(y)}d^{4}y,$
and this, together with the action of the $S$ matrix defined in (4.63) leads to the following commutator
$\phi_{1in}(x),SI(J)]=i\int\Delta(x-y)\hat{O}_{y}\frac{\delta SI(J)}{\delta J(y)}d^{4}y.$ (4.77)
In addition, it has been shown in (4.74) that the commutator of the fields is a c-number. By applying the following consequence of the Bakker-Campbell formula, which is true only for commutators which are c-numbers,
$A,e^{B}]=[A,B]e^{B},$
it can be seen that the general solution of the last equation (4.77) is
$SI(J)=e^{\int\phi_{1in}(z)\hat{O}_{z}\frac{\delta}{\delta J(z)}d^{4}z}F(J),$
with $F(J)$ arbitrary. The proof is found from the version of the Baker Campbell formula given above. To visualize this statement, note that formula and the last trial function imply
$\phi_{1in}(x),SI(J)]=\bigg[\int[\phi_{1in}(x),\phi_{1in}(z)]\hat{O}_{z}\frac{\delta}{\delta J(z)}d^{4}z\bigg]e^{-\int\phi_{1in}(y)\hat{O}_{y}\frac{\delta}{\delta J(y)}d^{4}y}F(J)$
$ie^{-\int\phi_{1in}(y)\hat{O}_{y}\frac{\delta}{\delta J(y)}d^{4}y}\bigg[\int\Delta(x-z)\hat{O}_{z}\frac{\delta F(J)}{\delta J(z)}d^{4}z\bigg].$
This characterizes the left hand side of the equation (4.77) The right hand side is
$i\int\Delta(x-y)\hat{O}_{y}\frac{\delta SI(J)}{\delta J(y)}d^{4}y=-i\int\Delta(x-y)\hat{O}_{y}e^{-\int\phi_{1in}(z)\hat{O}_{z}\frac{\delta}{\delta J(z)}d^{4}z}\frac{\delta F(J)}{\delta J(y)}d^{4}y$
The last two formulas are equal, and this shows that (4.77) is satisfied for every $F(J)$. This unknown function is determined by taking into account that, for flat Minkowski space, $S|0>=|0>$, and that
$0|e^{A}|0>=<0|:e^{A}:|0>=1,$
for any operator $A$. In these terms
<0|IS(J)|0>=<0|F(J)|0>=<0|I(J)|0>=Z(J).<0|IS(J)|0>=<0|F(J)|0>=<0|I(J)|0>=Z(J).
In the last step it was assumed that there are no external fields, and that this implies $S|0>=|0>$. In presence of such fields, this hypothesis has to be revised.
Therefore, it is concluded that the scattering operator is given by
$S=:e^{\int\phi_{1in}(z)\hat{O}\frac{\delta}{\delta J(z)}d^{4}z}:Z(J)\bigg|_{J=0},$ (4.78)
The LSZ reduction formula is then as in standard QFT, but the wave operator is replaced by the quartic wave one $\hat{O}$. The propagator $\Delta(x-y)$ has disappeared in the left hands side of the final formula, although it is inside the path integral $Z(J)$. As in (3.61), there is not a modification of $\phi$ in all this approach. All the calculation for scattering amplitudes is consistent with the field $\phi$ unmodified. However, if this reduction formula is expanded, it will lead to non standard results due to the non standard commutation relation.
It is needed to clarify how to operate with the last formula. However, this will be discussed in detail in the next sections, in a model which is more interesting namely, Quadratic Gravity.
5. The full Stelle action in synchronous
The previous ideas will be applied first to Stelle gravity [31]-[32] in the synchronous gauge. One of the main results of reference [1] is that the quantization presented above also cure the behavior of $Z(J)$ for this theory. Here the main points are reviewed, further details can be found in the original reference.
The Stelle lagrangian is
$S=\int[-\frac{2}{\kappa^{2}}R+bR^{2}+\frac{a}{2}W_{\mu\nu}W^{\mu\nu}]\sqrt{-g}d^{4}x,$
with the first term is the Einstein-Hilbert lagrangian and
$W_{\mu\nu}W^{\mu\nu}=\frac{1}{2}R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}-R_{\mu\nu}R^{\mu\nu}+\frac{R^{2}}{6}.$
The metric in the synchronous gauge is
$ds^{2}=-dt^{2}+g_{ij}(x,t)dx^{i}dx^{j}.$ (5.79)
In these coordinates the Christofell symbols read as follows
$\Gamma_{\mu\nu}^{\rho}=\frac{1}{2}g^{\rho\sigma}(\partial_{\mu}g_{\sigma\nu}+\partial_{\nu}g_{\sigma\mu}-\partial_{\sigma}g_{\mu\nu})$ (5.80)
$\Gamma^{0}_{ij}=\Gamma^{0}_{ji}=\frac{1}{2}\dot{g}_{ij}\equiv-K_{ij},\quad\Gamma^{l}_{ij},\quad\Gamma^{l}_{0j}=\Gamma^{l}_{j0}=\frac{1}{2}g^{lm}\dot{g}_{mj}\equiv-K^{l}_{~j},$ (5.81)
where a dot represents a derivative with respect to $t$. The curvature terms in the synchronous gauge is decomposed as [1]
$R=R^{3}-2g^{ij}\dot{K}_{ij}-3K_{ij}K^{ij}+K^{2},$
$R_{\mu\nu}R^{\mu\nu}=\left[g^{ij}(\dot{K}_{ij}+K_{il}K^{l}_{~j})\right]^{2}-2(D_{i}K-D^{j}K_{ji})(D^{i}K-D^{l}K_{l}^{~i})$
$(~R^{3}_{ij}-\dot{K}_{ij}-2K_{il}K^{l}_{~j}+KK_{ij})g^{il}g^{jm}(~R^{3}_{lm}-\dot{K}_{lm}-2K_{lp}K^{p}_{~m}+KK_{lm}),$
$R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}=4(\dot{K}_{ij}+K_{il}K^{l}_{~j})g^{il}g^{jm}(\dot{K}_{lm}+K_{lp}K^{p}_{~m})$
$4(D^{j}K_{il}-D_{l}K^{j}_{~i})(D_{j}K^{il}-D^{l}K_{j}^{~i})$
$(~R^{3}_{ijlm}+K_{il}K_{jm}-K_{jl}K_{im})(~R^{\hskip-11.38092pt3\hskip 8.5359ptijlm}+K^{il}K^{jm}-K^{jl}K^{im}).$ (5.82)
Here  $R^{3}_{ijlm}$,  $R^{3}_{ij}$,  $R^{3}$ and $D_{i}$ are, respectively, the three-dimensional Riemann tensor, Ricci tensor, Ricci scalar and covariant derivative built with the three-dimensional metric $g_{ij}$.the inverse three-dimensional metric is represented as usual by $g^{ij}$, and the following quantity $K\equiv K_{i}^{~i}$ was also introduced in those formulas.
The quantity $\dot{K}_{ij}$ is time-reversal invariant and appears linearly, while $K_{ij}\equiv-\dot{g}_{ij}/2$ is odd but appears quadratically. The full lagrangian is time reversal.
As done in previous section by following the Ostrogradsky method, the variables $g_{ij}$, $\dot{g}_{ij}$ or, equivalently, of $g_{ij}$, $K_{ij}$ can be chosen. In this formalism $g_{ij}$ and $K_{ij}$ are independent canonical coordinates. The second choice will employed leading to the following conjugate momentum densities
$\pi^{ij}\equiv\frac{\partial L}{\partial\dot{g}_{ij}}-\frac{d}{dt}\frac{\partial L}{\partial\ddot{g}_{ij}},\quad P^{ij}\equiv\frac{\partial L}{\partial\dot{K}_{ij}}-\frac{d}{dt}\frac{\partial L}{\partial\ddot{K}_{ij}}$ (5.83)
Since $L_{s}$ is independent of $\ddot{K}_{ij}$,
$P^{ij}=\frac{\partial L}{\partial\dot{K}_{ij}}.$
This leads to
$P^{ij}=-2\sqrt{-g}[G^{ijlm}\dot{K}_{lm}+aK^{ij}K-ag^{ij}K_{lm}K^{lm}+aR^{ij}$
$g^{ij}[(\frac{a}{3}+2b)(~R^{3}-3K_{lm}K^{lm}+K^{2})],$ (5.84)
where the quantity
$G^{ijlm}=a\,\frac{g^{il}g^{jm}+g^{im}g^{jl}}{2}+\left(4b-\frac{a}{3}\right)g^{ij}g^{lm},$
has been introduced. Note that, for non vanishing $\alpha$ and $\beta$ the quantity
$\Gamma_{pqij}\equiv\frac{1}{a}\frac{g_{pi}g_{qj}+g_{pj}g_{qi}}{2}-\frac{4b-a/3}{12ba}g_{pq}g_{ij}$ (5.85)
is the inverse of $G^{ijlm}$ , with the definition of inverse given by
$\Gamma_{pqij}G^{ijlm}=\frac{1}{2}(\delta_{p}^{l}\delta_{q}^{m}+\delta_{p}^{m}\delta_{q}^{l}).$
The crucial point is the following. The last expressions show that $\dot{K}_{ij}$ can be expressed in terms of $P^{ij}$, $g_{ij}$ and $K_{ij}$. Inserting this expression in $L_{s}$ the resulting functional $g_{ij}$, $K_{ij}$ and $P^{ij}$ with no dependence on $\pi^{ij}$. Therefore the Ostrogradsky Hamiltonian
$H=\pi^{ij}\dot{g}_{ij}+P^{ij}\dot{K}_{ij}-L_{s}.$ (5.86)
As the Stelle lagrangian $L_{s}$ does not depend on $\pi_{ij}$ it is clear that the only dependence is given by
$\pi^{ij}\dot{g}_{ij}=-2\pi^{ij}K_{ij}.$ (5.87)
This is a linear term in $\pi_{ij}$ and, if the path integral formalism is applied, the resulting path integral $Z(J)$ will involve a linear exponential in $\pi_{ij}$ whose integration will diverge. Therefore the path integral would not exist.
The above non existence argument does not hold if the quantization presented in previous sections is applied. This remarkable observation was done in [1]. The point is that $K_{ij}$ and its conjugate momentum $\pi_{ij}$ are ghost like variables, since they change the sign under time reversal, and in this quantization the Hamiltonian must be continued to imaginary values of these variables. This will convert the divergent exponential into an imaginary one, which is harmless.
To see this statement more explicitly, consider the quantization of Stelle gravity in the synchronous gauge. From (5.84) it is obtained that
$\dot{K}_{ab}=\Gamma_{abij}\bigg\{\frac{P^{ij}}{2\sqrt{-g}}+aK^{ij}K-ag^{ij}K_{lm}K^{lm}+a_{ij}$
$g^{ij}\left[\left(\frac{a}{3}+2b\right)(~R^{3}-3K_{lm}K^{lm}+K^{2}\right]\bigg\},$
By use of this formula replace $\dot{K}_{ij}$ in all the curvatures given in (5.82). The result will depend on $P^{ij}$, $g_{ij}$ and $K_{ij}$ but not on $\pi_{ij}$, as the last expression (5.84) does not depend on this quantity. The Hamiltoian is then an expression of the form
$H=\pi^{ij}\dot{g}_{ij}+P^{ij}\dot{K}_{ij}-L_{s}(g_{ij},P^{ij},K_{ij}).$
It is needed to analitically continue the classical Hamiltonian to imaginary values of the variables: $K_{ij}\to iK_{ij}$, $P^{ij}\to-iP^{ij}$ and remembering that $\dot{g}_{ij}=-2K_{ij}$
$H=-2i\pi^{ij}K_{ij}-iP^{ij}\dot{K}_{ij}-L_{s}(g_{ij},-iP^{ij},iK_{ij}).$
In the last expression it is understood that $K_{ij}$ should be written as (5.84). The path integral, in its euclidean version, where the metric is
$ds^{2}_{E}=d\tau^{2}+g_{ij}(x)dx^{i}dx^{j}.$
is given by
$Z_{e}(0)=\int D\pi^{ij}DK_{ij}DP^{ij}Dg_{ij}\exp\bigg\{-\int[i\pi^{ij}\dot{g}_{ij}+P^{ij}\dot{K}_{ij}-H(g_{ij},\pi_{ij},-iP^{ij},iK_{ij})]d\tau\bigg\}$
$\int D\pi^{ij}DK_{ij}DP^{ij}Dg_{ij}$
$\exp\bigg\{-\int[i\pi^{ij}\dot{g}_{ij}+iP^{ij}\dot{K}_{ij}-2i\pi^{ij}K_{ij}-iP^{ij}\dot{K}_{ij}-L_{s}(g_{ij},-iP^{ij},iK_{ij})]d\tau\bigg\}$
[…]
The integral over $P^{ij}$ is a Gaussian integral after setting $K_{ij}=-\frac{1}{2}g_{ij}^{\prime}$ , this follows after inspection of (5.82). The general formula
$\int d^{N}xe^{x^{T}Ax+bx}=\frac{\pi^{\frac{N}{2}}}{\sqrt{\det A}}e^{-\frac{1}{4}b^{T}Ab},$
can be employed to make this integral. Alternatively, they can be performed with a standard stationary point method. The stationary point satisfies now the equation
$iK_{ij}^{\prime}=\frac{\partial{\cal H}}{\partial P^{ij}},$
results in
$\frac{1}{2}g_{ij}^{\prime\prime}=\frac{\partial{\cal H}}{\partial(-iP^{ij})}.$
The last equation allows to express $P^{ij}$ in terms of $g_{ij}$, $\dot{g}_{ij}$ and $\ddot{g}_{ij}$ in classical physics but with $P^{ij}$ substituted by $iP_{ij}$. Regardless the method employed, the calculation is not difficult (it is just gaussian integration) the intermediate steps are a bit cumbersome. The final result is remarkable [1]-[2]. For finite initial and final euclidean times, after turning on a current $J^{ij}$, it is
$Z_{e}(J,\tau_{i},\tau_{f})=\int^{q(\tau_{f})=q_{f}}_{q(\tau_{i})=q_{i}}\,CDg_{ij}\,\exp\left(-\frac{S_{E}}{\hbar}+\int_{\tau_{i}}^{\tau_{f}}d\tau\int d^{3}x\,J^{ij}g_{ij}\right),$ (5.88)
where $S_{E}$ is the Euclidean version of the Stelle action
$S_{E}=\int_{\tau_{i}}^{\tau_{f}}d\tau\int d^{3}x\sqrt{g}\left(\frac{a}{2}W_{E}^{2}+bR_{E}^{2}+\frac{2}{\kappa^{2}}R_{E}\right).$ (5.89)
This result is important for the following reason. The theory that is constructed with this quantization is going to be unitary. However, it was not warranted that it will coincide with the euclidean Stelle gravity version. The Stelle theory is known to be renormalizable, and this property would be broken if another lagrangian was to be found. The fact that it coincides means that all the good properties of the Stelle model will be preserved, in particular renormalizability, while keeping unitarity [1].
The extra factor in the last path integral is
$C\sim\prod_{x_{E}^{\prime}}\frac{1}{\sqrt{g(x_{E})}}.$
This factor follows from the Gaussian integration formula directly. It is well known that it may introduce $\delta^{4}(0)$ singularities. However, if perturbation around flat spaces $\eta_{\mu\nu}$ are considered and dimensional regularization is employed, the well known Veltmann identities throw these terms to zero. The quantity $J^{ij}$ is the external “current” corresponding to $g_{ij}$ in the generating functional.
The above path integral is done in the synchronous gauge. Other gauges may be of interest, such as the De Donder one. The change to a general gauge leads to
$Z(J)=N\int\,{\cal D}g\,\left(\det\frac{\delta f}{\delta\xi}\right)\,\delta(f)\exp\left(-\frac{S_{E}}{\hbar}+\int d^{4}x_{E}\,J^{\mu\nu}g_{\mu\nu}\right),$ (5.90)
where the normalization factor is
$N=\frac{1}{\int\,{\cal D}g\,\left(\det\frac{\delta f}{\delta\xi}\right)\,\delta(f)\exp\left(-\frac{S_{E}}{\hbar}\right)}$
The determinant $\det\frac{\delta f}{\delta\xi}$ is the Faddeev-Popov determinant that usually appears when fixing a gauge in such theories.
After having the euclidean version of the model, the Lorentzian Green’s functions may be found in the following way. Replace the euclidean currents $J^{\mu\nu}(\tau)$ with a Lorentzian currents
$J^{kl}(\tau)\to J^{kl}(it)/\hbar,\qquad J^{k4}(\tau)\to iJ^{k0}(it)/\hbar,\qquad J^{44}(\tau)\to-J^{00}(it)/\hbar,$
and also make the replacement
$g_{kl}(\tau)\to g_{kl}(it),\qquad g_{k4}(\tau)\to-ig_{k0}(it),\qquad g_{44}(\tau)\to-g_{00}(it).$
These substitutions lead to
$Z(J)=N\int\,Dg\,\left(\det\frac{\delta f}{\delta\xi}\right)\,\delta(f)\exp\left(\frac{iS}{\hbar}+i\int d^{4}x\,J^{\mu\nu}g_{\mu\nu}/\hbar\right),$ (5.91)
where ${\cal Z}(J)$ is the generating functional of the Lorentzian Green’s functions. It is defined only as an analytic continuation of the Euclidean one.
The next task is to characterize the second quantization and the LSZ rules of the model.
6. The Stelle model reconsidered
Consider again the formulas (6.93). These formulas show that the oscillator structure constants corresponding to the free Stelle model are the same as the Pais-Uhlenbeck oscillator found above sections. The task is to understand which new constants does the continuation $P_{2}\to iP_{2}$ and $Q_{2}\to iQ_{2}$ to imaginary values induces.
6.1 About the mapping $P_{2}\to iP_{2}$ and $Q_{2}\to iQ_{2}$
The reasoning given above leads to the coefficients $c_{i}$ defining the commutation relations for the creation and annihilation operators in (2.40). However, the quantization prescription requires to make the continuation of the ghost variables to imaginary values, that is, $P_{2}\to iP_{2}$ and $Q_{2}\to iQ_{2}$. It is natural to ask how (2.40) is changed under this continuation. Assume that it leads to new commutation relations
$A_{p}(k),A_{q}^{c}(k^{\prime})]=d_{1}\delta_{pq}\omega_{1}(k)\delta(k-k^{\prime}),\qquad[B_{p}(k),B_{q}^{c}(k^{\prime})]=d_{2}\delta_{pq}\omega_{2}(k)\delta(k-k^{\prime}),$
$C_{p}(k),C_{q}^{c}(k^{\prime})]=d_{3}\delta_{pq}\omega_{2}(k)\delta(k-k^{\prime}),$ (6.92)
with new constants $d_{i}$ with $i=1,2,3$. The capital letter is to indicate that these creation and annihilation operators are not the same as in (2.40). The task is to understand how to calculate these new constants $d_{i}$. Note that even though $c_{3}=0$, it was not assumed here that $d_{3}=0$, the idea is to be as general as possible in the following discussion.
In the above presentation, one condition was to ensure that
$Q_{1\mu\nu},Q_{2\mu\nu}]=[h_{1\mu\nu},\dot{h}_{2\mu\nu}]=\text{pure gauge terms}.$
This is not the case now, and can be seen as follows. As shown above in (2.36)
$Q_{2\mu\nu},P_{2\mu\nu}]=[h_{1\mu\nu},\dot{h}_{2\mu\nu}]-[P_{1\mu\nu},Q_{1\mu\nu}].$
Since the continuation $P_{2}\to iP_{2}$ and $Q_{2}\to iQ_{2}$ change the sign of the commutator of the left on the last formula, now the identity term of $h_{1\mu\nu},\dot{h}_{2\mu\nu}]-[P_{1\mu\nu},Q_{1\mu\nu}]$ has to have the right sign. This is possible only if now $h_{1\mu\nu},\dot{h}_{2\mu\nu}]$ has a non zero identity term, engineered in such a way that it changes the sign of $P_{1\mu\nu},Q_{1\mu\nu}]$. Therefore now it has to required that both
$P_{1\mu\nu}(x),Q_{1\alpha\beta}(y)]=\frac{i}{2}[(\eta_{\alpha\mu}\eta_{\beta\nu}+\eta_{\alpha\nu}\eta_{\beta\mu})+\text{gauge terms}]\delta(x-y),$
$P_{2\mu\nu}(x),Q_{2\alpha\beta}(y)]=\frac{i}{2}[(\eta_{\alpha\mu}\eta_{\beta\nu}+\eta_{\alpha\nu}\eta_{\beta\mu})+\text{other gauge terms}]\delta(x-y).$
The minus sign of the second commutator is now avoided. This leads to two conditions for the coefficients $d_{i}$.
It is interesting that one condition is still missing in the analysis. There are three coefficients and, up to now, two equations. It may be enforced that $P_{1\mu\nu}(x),P_{2\alpha\beta}(y)]$ is a pure gauge, as before. This was the case in the Pais-Uhlenbeck oscillator, as show in (3.54). Based on this, the same condition may be assumed here. This leads to $d_{3}=0$, again the third mode is not present as an external line. The resulting mapped constants are
$d_{1}=d_{2}=1.$ (6.93)
Again, these are the same as the transformed constants for the Pais-Uhlenbeck oscillator,, described in section 2.
6.2 The LSZ formula for Stelle gravity in detail
By collecting all the information given above, the LSZ rules for the model may be written. They are given by analogy with (4.78)
$\hat{S}=:\exp\bigg\{\frac{2}{\kappa^{2}}\int h^{\mu\nu}_{in}(x)\square\bigg[\frac{1}{2}\bigg(1-\frac{\kappa^{2}\alpha}{4}\square\bigg)P^{2}-\bigg(1+\frac{\kappa^{2}(\alpha-3\beta)}{2}\square\bigg)P^{0-s}$
$\bigg(\frac{1}{2}+\frac{\kappa^{2}(4\beta-3\alpha)\square}{8}\bigg)P^{0-w}+\frac{\kappa^{4}}{8\lambda}\square[P^{1}-P^{0-w}]\bigg]_{\alpha\beta,\mu\nu}\frac{\delta}{\delta J_{\alpha\beta}(x)}d^{4}x.\bigg\}:Z(J)\bigg|_{J=0}.$ (6.94)
Here $h^{\mu\nu}_{in}$ is a free field graviton operator of the form described in the previous sections. Despite the similarity with the standard textbook formula, there are differences. The prescription needed to sandwich this operator between the physical states is the crucial point for these differences.
As the quantization prescription requires $P_{2}\to iP_{2}$ and $Q_{2}\to iQ_{2}$, it follows that the creation and annihilation operators (6.92) in the field $h_{\mu\nu}^{in}$ in (6.94) have to be employed. In other words, the operators $A_{p}(k)$, $B_{p}(k)$ are the ones inside $h^{in}_{\mu\nu}$, these correspond to the structure constants $d_{i}$ with $i=1,2$. The operator $C_{p}(k)$ is gone, since $d_{3}=0$. The third mode does not participate as an external line. Explicitly the free graviton operator is
$h_{in}^{\mu\nu}=\sum_{p}\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}\omega_{1}(k)}[A_{p}(k)\epsilon_{1}^{p\mu\nu}(k)e^{i\omega_{1}(k)t-ik\cdot x}+A^{c}_{p}(k)\epsilon_{1}^{\ast p\mu\nu}(k)e^{-i\omega_{1}(k)t+ik\cdot x}]$
$\sum_{p}\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}\omega_{2}(k)}[B_{p}(k)\epsilon_{2}^{p\mu\nu}(k)e^{i\omega_{2}(k)t-ik\cdot x}+B^{c}_{p}(k)\epsilon_{2}^{\ast p\mu\nu}(k)e^{-i\omega_{2}(k)t+ik\cdot x}].$
There is no need to do the continuation $P_{2}\to iP_{2}$ and $Q_{2}\to iQ_{2}$ with these operators, they are constructed in such a way that this is automatic. The operators satisfy the algebra
$A_{p}(k),A_{q}^{c}(k^{\prime})]=\delta_{pq}\omega_{1}(k)\delta(k-k^{\prime}),\qquad[B_{p}(k),B_{q}^{c}(k^{\prime})]=\delta_{pq}\omega_{2}(k)\delta(k-k^{\prime}),$
that is, the standard one. The expansion of the exponential in (6.94) leads to terms related to correlation functions of the following form
$\prod_{i=1}^{n}\hat{h}^{in}_{\gamma\delta}(x^{i})\hat{O}^{\gamma\delta\mu\nu}_{x^{i}}<\hat{T}h_{\mu\nu}(x^{1})...h_{\alpha\beta}(x^{n})>,$ (6.95)
where $\hat{O}$ is the operator defined in the exponential in (6.94). This operator removes external legs. The full element (6.95) is to be sandwiched between states which a definite momentum. The prescription is that these states have to be constructed in standard fashion, for instance, a state with definite momentum for the free $h_{3in}^{\mu\nu}$ component
$|k_{1},p>_{in}=A^{c}_{p(in)}(k_{3})|0>,\qquad|k_{2},p>_{in}=B^{c}_{p(in)}(k_{3})|0>.$
The tensor product of these states gives multi-particle states.
The use of these states ensures that the $S$ matrix (4.62) is unitary, as they are standard. The part of the Fock space $|k_{3}>$ corresponding the third mode and their tensor produces is decoupled. The unitarity of the $S$ matrix (4.62) is not spoiled. Two modes are interacting and one is decoupled.
6.3 Universality properties of the effective action
The above discussion was related to free graviton fields $h_{in\mu\nu}$, which are relevant for studying scattering as they enter in the LSZ formula. It is very difficult to make a discussion about this reduction formula without a proper second quantization of the free or asymptotic fields.
The other important quantity is the partition function $Z(J)$, which is fundamental for calculating the correlation functions in (6.95). The expression consistent with the gauge considered in the original reference [31], described here in (2.24) is
$Z(\xi,\overline{\xi},K,L,J)=e^{iG(\xi,\overline{\xi},K,L,J)}=\int Dh^{\mu\nu}D\eta^{a}D\overline{\eta}^{b}\exp\{i\int d^{4}x[\overline{\xi}_{\alpha}\eta^{\alpha}+\xi_{\alpha}\overline{\eta}^{\alpha}+\kappa J_{\mu\nu}h^{\mu\nu}]\}$
$\exp\{i\int d^{4}x[L_{s}-\frac{\kappa^{2}}{2\lambda}F_{\tau}\square F^{\tau}+\overline{\eta}_{\tau}F_{\mu\nu}^{\tau}D^{\mu\nu}_{\alpha}\eta^{\alpha}+\kappa K_{\mu\nu}D^{\mu\nu}_{\alpha}\eta^{\alpha}+\kappa^{2}L_{\alpha}\partial_{\beta}\eta^{\alpha}\eta^{\beta}]\}.$ (6.96)
Here the current coupling is $J_{\mu\nu}h^{\mu\nu}$ and $J_{\mu\nu}$ may be identified as the energy momentum tensor $J_{\mu\nu}=T_{\mu\nu}$ for an external source. The quantity $\lambda$ is simple a constant parameter in the gauge fixing term, as discussed in the previous section. As in those theories, the sources $K_{\mu\nu}$ and $L_{\alpha}$ are coupled to the nilpotent part of the BRST invariance. The $\eta_{a}$ and $\overline{\eta}_{a}$ fields are the Faddeev-Popov ghosts and anti-ghost respectively. The standard transition amplitude is
$Z(\xi,\overline{\xi},J)=Z(\xi,\overline{\xi},J,K,L)|_{K=L=0}$
From the expression of the above functional, it follows that the gauge ghost action for the gravitational perturbation can be expressed as follows
$S_{gf}=\int d^{4}x[L_{s}-\frac{\kappa^{2}}{2\lambda}F_{\tau}\square F^{\tau}+\overline{\eta}_{\tau}F_{\mu\nu}^{\tau}D^{\mu\nu}_{\alpha}\eta^{\alpha}+\kappa K_{\mu\nu}D^{\mu\nu}_{\alpha}\eta^{\alpha}+\kappa^{2}L_{\alpha}\partial_{\beta}\eta^{\alpha}\eta^{\beta}].$
The operator related to the Stelle gauge choice is $F^{\tau}=\partial_{\nu}h^{\tau\nu}$. Furthermore $F^{\tau}_{\mu\nu}=\delta^{\tau}_{\mu}\partial_{\nu}$. The mass dimensions of the fields are directly seen from this action
$h^{\mu\nu}]=0,\qquad[\eta^{a}]=[\overline{\eta}^{a}]=1,\qquad[K_{\mu\nu}]=[L_{\alpha}]=3.$ (6.97)
Here, as before, the values for the ghost and anti-ghost are non uniquely defined. The choice above is however valid, and is the one that is usually employed. The operator $D_{\alpha}^{\beta\gamma}$ is defined by
$D_{\alpha}^{\mu\nu}\xi^{\alpha}=\partial^{\mu}\xi^{\nu}+\partial^{\nu}\xi^{\mu}-\eta^{\mu\nu}\partial_{\alpha}\xi^{\alpha}+h^{\alpha\mu}\partial_{\alpha}\xi^{\nu}+h^{\alpha\nu}\partial_{\alpha}\xi^{\mu}-\xi^{\alpha}\partial_{\alpha}h^{\mu\nu}-h^{\nu\nu}\partial_{\alpha}\xi^{\alpha}.$ (6.98)
As is well known, there is a deep relation between the functional $G$ given above and the effective action $\Gamma$. The relation is
$e^{iG(J,\xi,\overline{\xi},K,L)}=\int_{\text{C.T.D}}Dh^{\mu\nu}D\eta^{a}D\overline{\eta}^{a}\exp\{i\int\;d^{4}x\;[\Gamma(h,\eta,\overline{\eta},K,L)+$
$\overline{\xi}_{\alpha}\eta^{\alpha}+\xi_{\alpha}\overline{\eta}^{\alpha}+\kappa J_{\mu\nu}h^{\mu\nu}]\},$ (6.99)
where the notation C.T.D. stands for connected three diagrams. The physical meaning of the last formula is the following. If functional form of the effective action $\Gamma(h,\eta,\overline{\eta},K,L)$ has been obtained exactly, the functional $G(J,\xi,\overline{\xi},K,L)$ can be found by using the effective action in place of the original action and by calculating Feynman diagrams only at tree level, without loops.
An important point is that, with the gauge described in the above sections, the Stelle effective action has the universal form
$\Gamma_{n}=\int d^{4}x\bigg[A(h_{\mu\nu})+\frac{\delta S_{0}}{\delta h^{\mu\nu}}P^{\mu\nu}+(\kappa K_{\rho\sigma}-\overline{\eta}_{\tau}\overleftarrow{F}^{\tau}_{\rho\sigma})\bigg(\frac{\delta D^{\rho\sigma}_{\alpha}}{\delta h^{\mu\nu}}\eta^{\alpha}P^{\mu\nu}$
$\frac{\partial P^{\rho\sigma}}{\delta h^{\mu\nu}}D^{\mu\nu}_{\alpha}\eta^{\alpha}-D^{\rho\sigma}_{\alpha}(Q^{\alpha}_{\tau}\eta^{\tau})\bigg)$
$\kappa L_{\sigma}\bigg(\kappa Q^{\sigma}_{\tau}\partial_{\beta}\eta^{\tau}-\kappa\partial_{\beta}(Q^{\sigma}_{\tau}\eta^{\tau})-\kappa(\partial_{\tau}\eta^{\sigma})Q^{\tau}_{\beta}-\frac{\delta Q^{\sigma}_{\tau}}{\delta h^{\mu\nu}}\eta^{\tau}D_{\beta}^{\mu\nu}\bigg)\eta^{\beta}\bigg].$ (6.100)
This is exactly the formula (6.17) of the original reference [32]. Here the following quantity
$A(h_{\mu\nu})=\sqrt{-g}\bigg[-\frac{2}{\kappa^{2}+\delta\kappa^{2}}R+(\beta+\delta\beta)R^{2}-(\alpha+\delta\alpha)R_{\mu\nu}R^{\mu\nu}\bigg],$
has been introduced and $S_{0}$ is the classical lagrangian of Stelle gravity, in other words, the limit of $A(h_{\mu\nu})$ with the $\delta$ quantities taken to zero. The quantities $P^{\mu\nu}(h^{\alpha\beta})$ and $Q_{\nu}^{\mu}(h^{\alpha\beta})$ are functions of $h^{\alpha\beta}$ solely, and are not determined without a proper loop analysis for the effective action. Other gauges can be employed, this is allowed, however this universal form is not ensured. Its new expression should have to be worked out independently.
The third massive mode with mass $m_{3}^{2}$ is fundamental for the finding (6.100). The fact that is absent as an external line does not mean that it not contribute to the full correlation functions. In fact, these can be expressed in terms of the effective action $\Gamma_{n}$, whose form does take into account the effect of the third mode.
7. Final comments and possible future research programs
Before discussing future research programs, some comments about the prescription described between (6.94) and (6.95) are in order. Although the corresponding line of reasoning was described throughout the text, a further justification comes from Chapter 9 of the well known book [155]. In this chapter, these authors develop the path integral formalism starting with creation $a$ and annihilation operators $a^{\dagger}$ and take into account coherent states. On one hand, of the main points of the present work is that the third Stelle mode is gone as an external line. On the other hand, the full path integral $Z(J)$, or the effective action $\Gamma$ depend on the mode, even after the prescription that reference [1]. This may raise a doubt about the consistency of the throwing away procedure of the third mode. More precisely, in Chapter 9, the book [155] develop formulas such as (9-87) or (9.990), which state that the free path integral is
$Z_{0}(J)=\exp[\frac{i}{2}\int d^{4}xd^{4}yJ_{0}(x)\Delta(x-y)J_{0}(y)],$
while the interacting one is
$Z(J)=\exp[-i\int d^{4}xV(-\frac{i\delta}{\delta J})]Z_{0}(J).$
Here $V$ is the interaction term. These authors also derive the scattering LSZ in formula (9-88). These formulas are of course taught in every QFT course. But the particularity of this book presentation is that these formulas are deduced starting with the second quantized theory. In the present context, the interaction $V$ may be interpreted as all the interaction vertices that appear when expanding the Stelle action. The quantity $\Delta(x-y)$ would correspond to the Stelle propagator. This quantity depends on the three mass scales $m_{i}^{2}$ (2.25) with $i=1,2,3$ for each graviton mode. The formalism of [1], which includes the analytic continuation $P_{2}\to iP_{2}$ and $Q_{2}\to iQ_{2}$ rediscovers, in the euclidean setting, the euclidean version of Stelle gravity. The second quantization presented here is claimed to be in agreement with this continuation, and eliminates the mode with mass scale $m^{2}_{3}$. It may sound odd to work with a theory with a propagator depending on an scale $m_{3}^{2}$ that is not present in the mode expansion. However, this is not an inconsistency, and the comment made in (2.16) is done precisely for clarifying this. The point is that
$\hat{h}_{\mu\nu},\hat{h}_{\alpha\beta}]\neq D_{\alpha\beta,\mu\nu},$
so the left hand side may be independent on $m_{3}^{2}$ while being perfectly possible that the right hand is dependent. It is legal that the propagator depends on a mode which is absent in the free field expansion, although this situation may not be usual in standard Quantum Field Theory.
Note that the LSZ formulas derived in section 4 contains in some intermediary step the commutator of the scalar fields. These steps, adapted to Stelle gravity, will contain the commutator $\hat{h}_{\mu\nu},\hat{h}_{\alpha\beta}]$. However, at the end, these commutators do not appear in the final LSZ formula.
Another issue is that the $S$ matrix presented in this work is unitary. The explanation is simple. In the present work, the Fock states are taken as standard with positive norm, and the usually norm problems are attributed to bad creation and annihilation operators instead to bad norm states. In author´s opinion, the formalism presented here is equivalent to the pioneer ones, in particular [6], [7],[1], [2], [4]-[29]. The use of these states ensures that the $S$ matrix (4.62) is unitary. In fact, consider the Fock spaces $|k_{3}>$ corresponding the third mode and their tensor produces. There appears tensor products with the other components of the Fock space, $|k_{1}>$ and $|k_{2}>$. The Fock space corresponding to the third mode can be considered to decoupled. The unitarity of the $S$ matrix
$S=\sum_{\text{all states}}|r,\text{out}><r,\text{in}|,$
is not spoiled from this decoupling. There is one state that is decoupled and the other two interact each other.
The only point may be if continuation $P_{2}\to iP_{2}$ and $Q_{2}\to iQ_{2}$ spoils this property. But this prescription leads to Stelle gravity in an euclidean setting, and one mode is gone as external line. On the other hand, the Fock space is a tensor product, and this procedure, after continuing to Lorenz quantities, leads to a unitary theory with two graviton modes, one massless and one massive.
The theory is renormalizable. The path integral can be given in terms of an effective action that can be renormalized with a finite number of counter terms [31] which, in some specific term, leads to the general form (6.100). One then may employ tree diagrams approximation for calculating correlations. The Slavnov-Taylor identities [37]-[38] ensure order by order gauge invariance. A review is [39], further details can be found in the textbooks [156]-[161].
Recently, a reference [36] appeared which apparently presents a consistent quantization of the Pais-Uhlenbeck oscillator. It is of interest to understand how the methods developed there apply for Quadratic Gravity. A work that seems to have started this program is [43], which is of clear interest.
The formulation of the present model in different gauges is of course relevant. In this context, the quantization of [47]-[48] can be of special interest, see also [49]-[50]. I believe that the disappearance of the third mode is gage independent, but the study of different gauges is interesting to check this.
There are several potential applications related to the present formalism. A first one is related to Bremsstrahlung in gravity theories, and classical limits of gravity. A reference that considered this issue in detail is [82]. The calculation of this effect requires to take the mean value of the wave form $h_{\mu\nu}>$. This calculation, for Quadratic Gravity, involves the prescription presented here, in particular the mapping of the oscillator structure constants $c_{1}=1$, $c_{2}=-1$ to $d_{1}=d_{2}=1$ or, equivalently, $P_{2}\to iP_{2}$ and $Q_{2}\to iQ_{2}$. There are several works related to this topic, restricted to GR. The work [81] for instance, considers a worldline formalism which may be useful to be generalized to the present context. Other references related to this topic are [51]-[56], [136]-[137]. The works [57]-[60] present interesting results about wave forms, bound states and Hawking radiation which are worthy to be reconsidered in the present context.
The use of coherent states in order to study the classical limit of gravity and scattering process, which was employed in [61]-[62] may be of usefl in this context. In special, regarding the discussion about the consistency of the oscillator prescription described in the above paragraphs.
Not less important, it may be of clear interest to reformulate the issues regarded to the eikonal limit of gravity, in view of the quantization presented here. An extensive but perhaps incomplete list of works dealing with such topics are [63]-[148]. There are computational codes [149]-[152] and new formalisms [153] that may facilitate the calculations required when studying these subjects.
The renormalization of the model presented here is related to perturbations along flat spaces. It may seem that applications related to curved backgrounds such as [120]-[121] are outside of the present formalism. However, already in 1973 it was found in reference [75] a connection between quantum Feynman diagrams and the perturbation series of the classical Schwarzschild background. This topic was explored further in [76]-[85], related work is [92]-[98]. These connections between classical and quantum aspects of gravity are worthy to be generalized to the present context. Further applications may be found in [167].
Also, it may be interesting to study the Lyra version of the Stelle model [168]-[174] and consider processes which involve the interactions between graviton, matter and the scale fields. Finally, there are apparent super renormalizable generalizations of the Stelle model [162]-[166]. A treatment of their ghost structure is a desirable analysis for the future.
Acknowledgements
The authors are supported by CONICET, Argentina and O. S is supported by the Grant PICT 2020-02181.