### v4

# Scattering amplitudes in Quadratic Gravity in a general formalism

Abstract
In [1], inspired by the works [7]-[30], a prescription for calculating the correlation functions in Quadratic Gravity [31]-[32] was presented and further exploited in [2]-[5]. By construction, it is likely that this procedure does not enter in conflict with unitarity. The corresponding Hamiltonian quantization is based on a covariant and contra-variant distinction in the non positive definite metric in the space of states [13]. The Gauss-Ostrogradsky method for higher order theories defines two momentum densities $P_{1}$ and $P_{2}$ and two coordinate densities $Q_{1}$ and $Q_{2}$, one pair is standard, the other ghost like. The approach in [1] involves the continuation $P_{2}\to iP_{2}$ and $Q_{2}\to iQ_{2}$ of the ghost variables acting on kets $|>$ after taking mean values. In the present work, following [42], the LSZ rules are derived, with a formalism adapted to full quartic or higher order theories. The derivation of these rules is possible only after the second quantization of the free theory. This is considered here, in a more general setting than [42]. It includes a Gupta-Bleuler characterization of the physical states. The effective action $\Gamma$ can be calculated following [31], and can be constrained by the Slavnov-Taylor identities [37]-[38].
1. Introduction
Stelle’s gravity [31]-[32] is a renormalizable gravity model, which contains a massless mode and two massive ones. In the original reference, the renormalization program for this model was achieved in terms the Slavnov-Taylor quantization method [37]-[38]. However, one of the massive modes contributes to the kinetic energy with a wrong sign, and was interpreted as a ghost which spoils unitarity. The theory is of higher order, its equations of motion are of fourth order and its Cauchy problem is well posed [40]-[41]. The unitarity problem was already pointed out as the main issue which obstructs it as a quantum gravity model, at least when the quantization is done around a flat background.
On the other hand, starting from the seminal works of Dirac and Pauli [6]-[7], several authors [8]-[30] and more recently [36] started different programs for quantization of higher order theories, with the intention of avoiding ghosts. These works are not only related to Quantum Gravity, instead they present different motivations, some of them [8]-[9] are even related to Statistical Physics. All of them bring interesting features which are not evident at first sight. An interesting approach was due to Boulware and Gross [13], who made a formal description of the non positive defined metric in the Hilbert space. These authors consider states such that their inner product $n|m>=\eta_{mn}$ is not positive definite. This reference describes a covariant $\eta_{nm}$, contra-variant $\eta^{nm}$ and mixed metric $\delta_{n}^{m}$. The mixed one is positive definite and is the one to be employed for calculating probabilities. The Hilbert space itself has covariant and contra variant components. The mean values have to be adapted to the covariant or contra variant Hilbert space, but usually negative and non bounded from below quantities become bounded by this covariant and contra-variant distinction. These authors [13] also present several useful formulas for the path integral for these theories, together with some interesting insights. In particular, that depending on the use of the covariant, contra-variant or mixed version, a self adjoint observable may be presented in terms of non hermitian matrices. Despite these interesting findings, those authors open the possibility that the path integral they define may not exist [13]. The main point is the possible presence of real and unbounded exponentials in the path integral, which make the full result divergent.
On the other hand, the authors [4]-[29] made several efforts to physically interpret the results of these references. In particular, in [1] the author made the remarkable observation that the quantization scheme of Boulware-Gross applied to the Stelle model leads, in the euclidean setting, to a perfectly defined path integral. This was further exploited in [2]-[5]. The potentially dangerous exponentials become imaginary and the full result is well defined (in a path integral sense). The main observation is that, in a Hamiltonian setting, there are variables which behave normally under time reversal, and there are ones that not. It is not possible to state in brief the results of these detailed works, but it is worthy to mention that the quantization scheme with these variables, denoted as $P_{2}$ and $Q_{2}$ is consistent after the continuation to imaginary values $P_{2}\to iP_{2}$ and $Q_{2}\to iQ_{2}$. Heuristically, if the commutator $a,a^{\dagger}]$ of an state of the mode has a wrong sign, this continuation removes the unwanted minus sign.
Furthermore, both the results of [13] and [1], strongly point out that this quantization does not enter in any conflict with unitarity. The final result of the whole continuation is the euclidean version of Stelle gravity, one may calculate the physical quantities in the euclidean setting, and then go to the Lorenzian one by continuation to real times. The results will be renormalizable, and is possible that unitarity will be preserved.
The results of all the references mentioned above are of significant interest. However, a full presentation requires an analysis of the LSZ formulas which, in this context, have to be taken with care. The point is that, in this quantization method, the creation and annihilation algebra may have some particularities not present in an standard quantum theory. A program about this topic was initiated in [42]. The LSZ rules require a second quantization of the free model consistent with the principles employed in all these references. In the present work, the idea is to present an LSZ deduction fully adapted to a fourth order theory, regardless how the modes are decomposed. A proper hamiltonian treatement, missing in [42] is included here.
It is important to remark that the bulk of this text presents a quantization scheme which the author believes to be equivalent to the one in [1]-[2]. A full proof will be desirable. The advantage is its compact form, which avoids the above mentioned covariant and contra-variant distinction. However, the main result presented here is related to the LSZ rules, the quantization program is a result that has to be attributed to all the references.
Furthermore, there is nothing wrong with the covariant and contra-variant formalism of [13]. The presentation here is not a critic against this formalism, it just a way for presenting the results in more concise manner. At the end, the references [1]-[5], [13] and [28]-[29] contain the full formalism, and can be consulted if some clarifications are needed since they include a large amount of details.
The present work is organized as follows. Section 2 shows the classical modes of the Stelle model. Also, the Gupta-Bleuler quantization is outlined. Section 3 reviews the Gauss-Ostrogradsky quantization for the Pais-Uhlenbeck oscillator, and how the ghost emerges. A simple prescription for avoiding the effect of these ghosts is presented. I believe that this prescription is equivalent to the one presented in [1], and is heavily inspired in that reference and also by [13]. The presentation is intended to be more compact, and this has a price of containing less details. In Section 4 the LSZ rules for the Pais-Uhlenbeck oscillator are derived, in a formalism fully adapted for higher order theories. Section 5 reviews the notable result of reference [1] about the finiteness of the path integral when the above described quantization is presented, which presents a quantization of the model which does not allow divergences due to exponential integrations in the path integral and, at the end, results in the euclidean version of Stelle gravity. In Section 6 the LSZ formulas adapted to this fourth order model is derived, the universal form of the effective action is specified and the effect of the non standard creation annihilation operators is clarified in the last part of this section. Section 7 contains possible research programs related to these results.
2. Classical and quantum aspects of Quadratic Gravity
2.1 The free Stelle action in the De Donder gauge
Consider the Stelle action [31]-[32], which will be taken as
$S_{s}=\int\sqrt{-g}[-\frac{\gamma}{\kappa^{2}}R+\beta R^{2}-\alpha R_{\mu\nu}R^{\mu\nu}]d^{4}x,$
Here $\gamma=2$ and $\kappa^{2}=32\pi G_{N}$. The last action may be linearized in $h^{\mu\nu}=g^{\mu\nu}-\eta^{\mu\nu}$ by taking into account that $\sqrt{g}=1+\frac{h}{2}+O(h^{2})$ and the curvature expansions
$R_{\mu\nu}=\frac{1}{2}[\partial^{\rho}\partial_{\mu}h_{\nu\rho}+\partial^{\rho}\partial_{\nu}h_{\mu\rho}-\square h_{\mu\nu}-\partial_{\mu}\partial_{\nu}h],\qquad R=\partial_{\mu}\partial_{\nu}h^{\mu\nu}-\square h+O(h^{2}),$
$\beta R^{2}=\beta h\square^{2}h+\beta h^{\mu\nu}\partial_{\mu}\partial_{\nu}\partial_{\alpha}\partial_{\beta}h^{\alpha\beta}-2\beta h\square\partial_{\mu}\partial_{\nu}h^{\mu\nu}+\text{total derivatives},$
$\alpha R_{\mu\nu}R^{\mu\nu}=-\frac{\alpha}{4}h\square^{2}h-\frac{\alpha}{4}h_{\mu\nu}\square^{2}h^{\mu\nu}+\frac{\alpha}{2}h_{\mu}^{\sigma}\square\partial_{\sigma}\partial_{\nu}h^{\nu\mu}$
$\frac{\alpha}{2}h\square\partial_{\mu}\partial_{\nu}h^{\nu\mu}-\frac{\alpha}{2}h^{\mu\nu}\partial_{\mu}\partial_{\nu}\partial_{\alpha}\partial_{\beta}h^{\alpha\beta}.$
$\text{total derivatives}$
$\beta R^{2}-\alpha R_{\mu\nu}R^{\mu\nu}=-\frac{\alpha}{4}h^{\mu\nu}\square[\square h_{\mu\nu}-2\partial_{\sigma}\partial^{\nu}h_{\nu\mu}]+(\beta-\frac{\alpha}{4})h\square[\square h-2\partial_{\mu}\partial_{\nu}h^{\nu\mu}]$
$(\beta-\frac{\alpha}{2})h^{\mu\nu}\partial_{\mu}\partial_{\nu}\partial_{\alpha}\partial_{\beta}h^{\alpha\beta}+\text{total derivatives}.$
At this order, the Fierz-Pauli action corresponding to GR is
$\frac{1}{16\pi G}\int\sqrt{g}Rd^{4}x\simeq-\frac{1}{\kappa^{2}}\int h^{\mu\nu}[\square h_{\mu\nu}-2\partial^{\rho}\partial_{(\mu}h_{\nu)\rho}]+\frac{h}{\kappa^{2}}[\square h-2\partial_{\mu}\partial_{\nu}h^{\mu\nu}]d^{3}x,$
this result follows only after expanding the curvature $R$ up to second order in $h^{\mu\nu}$.
The free Stelle action corresponds to the terms quadratic in $h^{\mu\nu}$. With these formulas, it is straightforward to find it, which will be the generalization of the Fierz-Pauli lagrangian adapted to the present context. The result is
$S_{s}=-\frac{2}{\kappa^{2}}\int\bigg[\frac{1}{2}h^{\mu\nu}\bigg(1-\frac{\kappa^{2}\alpha}{4}\Box\bigg)[\square h_{\mu\nu}-2\partial^{\gamma}\partial_{(\nu}h_{\mu)\gamma}]$
$\frac{1}{2}h\bigg(1-\frac{\kappa^{2}(4\beta-\alpha)}{4}\square\bigg)[2\partial_{\alpha}\partial_{\beta}h^{\alpha\beta}-\square h]-\frac{\kappa^{2}}{2}(\beta-\frac{\alpha}{2})h^{\mu\nu}\partial_{\mu}\partial_{\nu}\partial_{\alpha}\partial_{\beta}h^{\alpha\beta}\bigg]d^{4}x.$ (2.1)
If the parameters $\alpha$ and $\beta$ vanish, this reduces to the well known Fierz-Pauli action. The above deduction requires to take into account the $R$ terms up to second order and the remaining ones to lower order, since they will be quadratic in $h_{\mu\nu}$ automatically. The non gauged Stelle action may be expressed in an alternative form
$S_{s}=-\frac{2}{\kappa^{2}}\int\bigg[\frac{1}{2}h^{\mu\nu}\bigg(1-\frac{\kappa^{2}\alpha}{4}\Box\bigg)[\square h_{\mu\nu}-2\partial^{\gamma}\partial_{(\nu}h_{\mu)\gamma}]$
$\frac{1}{2}h\bigg[\frac{2}{3}\bigg(1-\frac{\kappa^{2}(3\beta-\alpha)}{2}\square\bigg)+\frac{1}{3}\bigg(1-\frac{\kappa^{2}\alpha}{4}\square\bigg)\bigg][2\partial_{\alpha}\partial_{\beta}h^{\alpha\beta}-\square h]$
$h^{\mu\nu}\bigg[\bigg(1-\frac{\kappa^{2}\alpha}{4}\partial_{\mu}\partial_{\nu}\bigg)-\bigg(1-\frac{\kappa^{2}(3\beta-\alpha)}{2}\partial_{\mu}\partial_{\nu}\bigg)\bigg]\partial_{\alpha}\partial_{\beta}h^{\alpha\beta}\bigg]d^{4}x.$ (2.2)
The reason for writing all in this form is that it is expressed in terms of the mass scales that will be found below in (2.22) namely
$m_{1}^{2}=0,\qquad m_{2}^{2}=\frac{4}{\kappa^{2}\alpha},\qquad m_{3}^{2}=\frac{2}{\kappa^{2}(3\beta-\alpha)}.$
However, one simply may consider this as a valid way to write the action. In any case, the equations of motion without fixing the gauge are
$\bigg(1-\frac{\kappa^{2}\alpha}{4}\Box\bigg)\square h_{\mu\nu}-\eta_{\mu\nu}\bigg[\frac{2}{3}\bigg(1-\frac{\kappa^{2}(3\beta-\alpha)}{2}\square\bigg)+\frac{1}{3}\bigg(1-\frac{\kappa^{2}\alpha}{4}\square\bigg)\bigg]\square h$
$2\bigg(1-\frac{\kappa^{2}\alpha}{4}\Box\bigg)\partial^{\gamma}\partial_{(\nu}h_{\mu)\gamma}+\eta_{\mu\nu}\bigg[\frac{2}{3}\bigg(1-\frac{\kappa^{2}(3\beta-\alpha)}{2}\square\bigg)+\frac{1}{3}\bigg(1-\frac{\kappa^{2}\alpha}{4}\square\bigg)\bigg]\partial_{\alpha}\partial_{\beta}h^{\alpha\beta}$
$\bigg[\frac{2}{3}\bigg(1-\frac{\kappa^{2}(3\beta-\alpha)}{2}\square\bigg)+\frac{1}{3}\bigg(1-\frac{\kappa^{2}\alpha}{4}\square\bigg)\bigg]\partial_{\mu}\partial_{\nu}h$
$\kappa^{2}(\beta-\frac{\alpha}{2})\partial_{\mu}\partial_{\nu}\partial_{\alpha}\partial_{\beta}h^{\alpha\beta}=0.$ (2.3)
The classical graviton field can be Fourier expanded. Now comes the problem of gauge fixing the model. For an easy visualization of results, consider the Lorenz gauge $\partial_{\mu}h^{\mu\nu}=0$ applied to the equations of motions. The De Donder gauge will be applied later on. The equations in the Lorenz gauge are
$\bigg(1-\frac{\kappa^{2}\alpha}{4}\Box\bigg)\square h_{\mu\nu}-\eta_{\mu\nu}\bigg[\frac{2}{3}\bigg(1-\frac{\kappa^{2}(3\beta-\alpha)}{2}\square\bigg)+\frac{1}{3}\bigg(1-\frac{\kappa^{2}\alpha}{4}\square\bigg)\bigg]\square h$
$\bigg[\frac{2}{3}\bigg(1-\frac{\kappa^{2}(3\beta-\alpha)}{2}\square\bigg)+\frac{1}{3}\bigg(1-\frac{\kappa^{2}\alpha}{4}\square\bigg)\bigg]\partial_{\mu}\partial_{\nu}h=0.$
Contracting the last differential equation with $\eta^{\mu\nu}$ gives
$\bigg(1-\frac{\kappa^{2}(3\beta-\alpha)}{2}\square\bigg)\square h=0.$ (2.4)
Therefore the trace of the metric $h$ is composed by masses $m^{2}_{1}=0$ and $m_{3}^{2}$. The mode with mass $m_{2}^{2}$ does not give contribution $h_{2}$ to the trace. The equations of motion reduce to
$\bigg(1-\frac{\kappa^{2}\alpha}{4}\Box\bigg)\square h_{\mu\nu}-\frac{1}{3}\eta_{\mu\nu}\bigg(1-\frac{\kappa^{2}\alpha}{4}\square\bigg)\square h$
$\bigg[\frac{2}{3}\bigg(1-\frac{\kappa^{2}(3\beta-\alpha)}{2}\square\bigg)+\frac{1}{3}\bigg(1-\frac{\kappa^{2}\alpha}{4}\square\bigg)\bigg]\partial_{\mu}\partial_{\nu}h=0.$ (2.5)
Decompose the trace $h$ into its components $h_{1}$ and $h_{3}$. The last equation is then
$\bigg(1-\frac{\kappa^{2}\alpha}{4}m_{3}^{2}\bigg)m_{3}^{2}h^{3}_{\mu\nu}+\frac{1}{3}\eta_{\mu\nu}m_{3}^{2}\bigg(1-\frac{\kappa^{2}\alpha}{4}m_{3}^{2}\bigg)h_{3}$
$\partial_{\mu}\partial_{\nu}h_{1}+\frac{1}{3}\bigg(1-\frac{\kappa^{2}\alpha}{4}m_{3}^{2}\bigg)\partial_{\mu}\partial_{\nu}h_{3}=0.$ (2.6)
The task is now to constraints the polarizations. For this, note that the Lorenz gauge condition implies that
$k^{i}_{\mu}\epsilon^{i\mu\nu}=0,$
where the index $i$ is indicating that the four momenta for every mode are different, since they satisfy a different mass shell relation. All three polarizations $\epsilon^{i}_{\mu\nu}$ with $i=1,2,3$ satisfy the last condition. Also, the fact that $h_{2}$ is zero is equivalent to
$\epsilon_{2}^{p\mu\nu}\eta_{\mu\nu}=0.$
The remaining condition follows from (2.6) and is
$\frac{1}{\omega_{3}(k)}\bigg(1-\frac{\kappa^{2}\alpha}{4}m_{3}^{2}\bigg)m_{3}^{2}\epsilon^{3}_{\mu\nu}+\frac{1}{3\omega_{3}(k)}\eta_{\mu\nu}m_{3}^{2}\bigg(1-\frac{\kappa^{2}\alpha}{4}m_{3}^{2}\bigg)\epsilon_{3}$
$\frac{1}{3\omega_{3}(k)}\bigg(1-\frac{\kappa^{2}\alpha}{4}m_{3}^{2}\bigg)k^{3}_{\mu}k^{3}_{\nu}\epsilon_{3}=0,$
$k_{\mu}k_{\nu}\epsilon_{1}=0.$
The second one follows since $h_{1}$ and $h_{3}$ have different Fourier expansions and have to be considered as independent. But this implies that $\epsilon_{1}=0$. Contracting the first with $k^{3\mu}$ and taking into account that leads to nothing. Therefore $h_{2}=h_{1}=0$ and the polarization $\epsilon^{i}_{\mu\nu}$ is orthogonal to $k_{i}^{\mu}$ (this is a bit loosely speaking, as the definition of orthogonality in four dimensions is subtle). If the direction of propagation is $\hat{z}$ it is known that the polarizations for the massless mode are
$\epsilon^{+}_{1\mu\nu}=\left(\begin{array}[]{cccc}0&0&0&0\\ 0&1&0&0\\ 0&0&-1&0\\ 0&0&0&0\end{array}\right),\qquad\epsilon^{x}_{1\mu\nu}=\left(\begin{array}[]{cccc}0&0&0&0\\ 0&0&1&0\\ 0&1&0&0\\ 0&0&0&0\end{array}\right)$ (2.7)
For the massive mode there more possible polarizations. A massive graviton with mass $m$ is not relativistic and in the rest frame its 4-momentum is $k^{\mu}=(m,0,0,0)$. A linearly independent set satisfying the transversality and trace zero conditions are
$\epsilon^{+}_{\mu\nu}=\left(\begin{array}[]{cccc}0&0&0&0\\ 0&1&0&0\\ 0&0&-1&0\\ 0&0&0&0\end{array}\right),\qquad\epsilon^{1}_{\mu\nu}=\left(\begin{array}[]{cccc}0&0&0&0\\ 0&1&0&0\\ 0&0&0&0\\ 0&0&0&-1\end{array}\right),$
$\epsilon^{x}_{\mu\nu}=\left(\begin{array}[]{cccc}0&0&0&0\\ 0&0&1&0\\ 0&1&0&0\\ 0&0&0&0\end{array}\right),\qquad\epsilon^{2}_{\mu\nu}=\left(\begin{array}[]{cccc}0&0&0&0\\ 0&0&0&1\\ 0&0&0&0\\ 0&1&0&0\end{array}\right),\qquad\epsilon^{3}_{\mu\nu}=\left(\begin{array}[]{cccc}0&0&0&0\\ 0&0&0&0\\ 0&0&0&1\\ 0&0&1&0\end{array}\right).$ (2.8)
Defining the sum over polarizations as
$P_{\mu\nu,\alpha\beta}=\sum_{\lambda}\epsilon^{\lambda}_{1\mu\nu}\epsilon^{\lambda}_{1\alpha\beta},$
it follows that the non zero components are
$P_{00,00}=0,\qquad P_{11,11}=2,\qquad P_{22,22}=P_{33,33}=1,$
$P_{12,12}=P_{21,21}=P_{13,13}=P_{31,31}=P_{23,23}=P_{32,32}=1,$
$P_{12,21}=P_{21,12}=P_{13,31}=P_{31,13}=P_{23,32}=P_{32,23}=1.$
By making a Lorenz transformation to a moving frame $k^{\mu}=(\omega(k),0,0,k)$
$\Lambda^{\mu}_{\nu}=\frac{1}{m}\left(\begin{array}[]{cccc}\omega&0&0&k\\ 0&m&0&0\\ 0&0&m&0\\ k&0&0&\omega\end{array}\right),$ (2.9)
the sum of polarizations can be calculated as
$P^{\prime}_{\mu\nu,\alpha\beta}=\Lambda^{\gamma}_{\mu}\Lambda^{\delta}_{\nu}\Lambda^{\epsilon}_{\alpha}\Lambda^{\kappa}_{\beta}P_{\gamma\delta,\epsilon\kappa}$
The sum depends on the masses. The polarization tensors in the new frame can also be calculated by making Lorentz transformations to the above ones (2.8).
The last thing to analyze is the third mode. After setting $h_{1}=h_{2}=0$ one has that
$\bigg(1-\frac{\kappa^{2}\alpha}{4}m_{3}^{2}\bigg)m_{3}^{2}h_{\mu\nu}-\frac{1}{3}\eta_{\mu\nu}\bigg(1-\frac{\kappa^{2}\alpha}{4}m_{3}^{2}\bigg)m_{3}^{2}h$
$\frac{1}{3}\bigg(1-\frac{\kappa^{2}\alpha}{4}m_{3}^{2}\bigg)\partial_{\mu}\partial_{\nu}h=0.$
The polarization is then
$\epsilon^{3}_{\mu\nu}=\frac{1}{3}\bigg(\eta_{\mu\nu}+\frac{k_{\mu}k_{\nu}}{m_{3}^{2}}\bigg)\epsilon.$ (2.10)
Therefore if the direction of propagation is $k_{\mu}=(\omega(k),0,0,k_{3})$ the polarization matrices reduce to a single one given by
$\epsilon^{+}_{\mu\nu}=\frac{1}{3}\left(\begin{array}[]{cccc}-1+\frac{\omega_{3}^{2}}{m_{3}^{2}}&0&0&\frac{k\omega_{3}}{m_{3}^{2}}\\ 0&1&0&0\\ 0&0&1&0\\ \frac{k\omega_{3}}{m_{3}^{2}}&0&0&1+\frac{k^{2}}{m_{3}^{2}}\end{array}\right).$
The field is decomposed as
$h_{in}^{\mu\nu}=\sum_{p}\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}\omega(k)}[\epsilon_{1\mu\nu}e^{i\omega(k)t-ik\cdot x}+\epsilon_{1\mu\nu}^{\ast}e^{-i\omega(k)t+ik\cdot x}]$
$\sum_{p}\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}\omega_{2}(k)}[\epsilon_{2\mu\nu}e^{i\omega_{2}(k)t-ik\cdot x}+\epsilon^{\ast}_{2\mu\nu}e^{-i\omega_{2}(k)t+ik\cdot x}].$
$\sum_{p}\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}\omega_{3}(k)}[\epsilon_{3\mu\nu}e^{i\omega_{3}(k)t-ik\cdot x}+\epsilon^{\ast}_{3\mu\nu}e^{-i\omega_{3}(k)t+ik\cdot x}].$
The sum is over all allowed polarizations, described above. The next topic is how this gauge conditions are imposed after quantization on the physical states.
2.1.1 The Gupta-Bleuler quantization scheme
Consider again the Lorenz gauge $\partial_{\mu}h^{\mu\nu}=0$. When making quantization of the model, one should be aware of imposing these classical considerations for the polarization operators. This awareness may be inspired by lessons about the Gupta-Bleuler quantization. Sometimes to impose a gauge at operator level, such as the Lorenz gauge $\partial_{\mu}A^{\mu}=0$ in QED, may be an over request. Instead, imposing a softer condition to the physical states which insures that the Lorenz gauge condition is assured when taking mean values make quantization realizable. An example is $\partial_{\mu}A^{+\mu}|\Psi>=0$, a condition heavily employed in this quantization method.
Consider the non gauged equations of motion without fixing given in (2.19). As before, the Lorenz gauge $\partial_{\mu}h^{\mu\nu}=0$ will be considered. However, both (2.19) and the gauge has to be valid only when mean values between physical states $|\Psi>$ are taken. The field is expanded as
$h^{\mu\nu}=\sum_{p}\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}\omega(k)}[a_{\mu\nu}a_{p}(k)e^{i\omega(k)t-ik\cdot x}+a_{\mu\nu}^{\ast}a_{p}^{c}(k)e^{-i\omega(k)t+ik\cdot x}]$
$\sum_{p}\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}\omega_{2}(k)}[b_{\mu\nu}b_{p}(k)e^{i\omega_{2}(k)t-ik\cdot x}+b^{\ast}_{\mu\nu}b_{p}^{c}(k)e^{-i\omega_{2}(k)t+ik\cdot x}].$
$\sum_{p}\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}\omega_{3}(k)}[c_{\mu\nu}c_{p}(k)e^{i\omega_{3}(k)t-ik\cdot x}+c^{\ast}_{\mu\nu}c_{p}^{c}(k)e^{-i\omega_{3}(k)t+ik\cdot x}].$
Here $a_{p}(k)$, $b_{p}(k)$ and $c_{p}(k)$ are annihilation operators, and the ones with the upperindex ”c” are the creation ones. The notation ”c” replaces $\dagger$} in order to leave the possibility for a non standard creation-annihiliation algebra.
The sum above is over all the polarizations, not only the allowed by the gauge conditions. The physical states are the ones which will restrict the sums. Therefore the Gupta-Bleuler request is
$\partial^{+}_{\mu}h^{\mu\nu}|\Psi>=0.$
This implies that
$\sum_{\lambda}k^{\mu}\epsilon^{\lambda}_{1\mu\nu}a_{\lambda}(k)|\Psi>=0,\qquad\sum_{\lambda}k^{\mu}\epsilon^{\lambda}_{2\mu\nu}b_{\lambda}(k)|\Psi>=0,$
$\sum_{\lambda}k^{\mu}\epsilon^{\lambda}_{3\mu\nu}c_{\lambda}(k)|\Psi>=0.$ (2.11)
Here $\lambda$ included all the polarizations, not only (2.7) and (2.8). The last gives a linear relation between the annihilation operators corresponding to these additional polarizations, whose action destroys the physical states. The averaged equations of motion are now
$\bigg(1+\frac{\kappa^{2}\alpha}{4}\Box\bigg)\square h_{\mu\nu}>-\eta_{\mu\nu}<\bigg[\frac{2}{3}\bigg(1+\frac{\kappa^{2}(3\beta-\alpha)}{2}\square\bigg)+\frac{1}{3}\bigg(1+\frac{\kappa^{2}\alpha}{4}\square\bigg)\bigg]\square h>$
$\bigg[\frac{2}{3}\bigg(1+\frac{\kappa^{2}(3\beta-\alpha)}{2}\square\bigg)+\frac{1}{3}\bigg(1+\frac{\kappa^{2}\alpha}{4}\square\bigg)\bigg]\partial_{\mu}\partial_{\nu}h>=0.$
Contracting the last differential equation with $\eta^{\mu\nu}$ gives
$\bigg(1+\frac{\kappa^{2}(3\beta-\alpha)}{2}\square\bigg)\square h>=0.$
The fact that the mean values do not contain the massive mode with $m_{2}^{2}$ imply that
$\sum_{\lambda}\epsilon_{2}^{\lambda}b_{\lambda}(k)|\Psi>=0.$ (2.12)
The states (2.7) and (2.8) have zero traces. The remaning are not traceless and have non zero $\epsilon^{\lambda}$. The condition (2.12) gives a new linear relation for the annihilation operators. The averaged equations of motion reduce to
$\bigg(1+\frac{\kappa^{2}\alpha}{4}\Box\bigg)\square h_{\mu\nu}>-<\frac{1}{3}\eta_{\mu\nu}\bigg(1+\frac{\kappa^{2}\alpha}{4}\square\bigg)\square h>$
$\bigg[\frac{2}{3}\bigg(1+\frac{\kappa^{2}(3\beta-\alpha)}{2}\square\bigg)+\frac{1}{3}\bigg(1+\frac{\kappa^{2}\alpha}{4}\square\bigg)\bigg]\partial_{\mu}\partial_{\nu}h>=0.$
The absence of the mode with $m_{2}^{2}$ in the average leads to
$\bigg(1-\frac{\kappa^{2}\alpha}{4}m_{3}^{2}\bigg)m_{3}^{2}<h^{3}_{\mu\nu}>+\frac{1}{3}\eta_{\mu\nu}m_{3}^{2}\bigg(1-\frac{\kappa^{2}\alpha}{4}m_{3}^{2}\bigg)<h_{3}>$
$\partial_{\mu}\partial_{\nu}<h_{1}>+\frac{1}{3}\bigg(1-\frac{\kappa^{2}\alpha}{4}m_{3}^{2}\bigg)\partial_{\mu}\partial_{\nu}<h_{3}>=0.$ (2.13)
The independence between the modes $1$ and $3$ leads to
$k_{\mu}k_{\nu}\sum_{\lambda}\epsilon_{1}^{\lambda}a_{\lambda}(k)|\Psi>=0,$
For the third mode one has from (2.10) that
$\sum_{\lambda}\epsilon^{3\lambda}_{\mu\nu}c_{\lambda}(k)||\Psi>=\sum_{\lambda}\frac{1}{3}\bigg(\eta_{\mu\nu}+\frac{k_{\mu}k_{\nu}}{m_{3}^{2}}\bigg)\epsilon^{\lambda}c_{\lambda}|\Psi>$
Quantization should be performed by considering general polarizaion and assuming that the physical states are constrained by the above conditions.
2.2 The free action in the Stelle gauge
The original reference [31] employs a gauge which is a generalization of the Lorenz gauge employed above, but with the particularity that induces terms which decay as $k^{-4}$. This behavior makes the renormalization analysis presented in that reference more accessible. It is not forbidden to employ other gauges, however, the renormalizability issues raised in that classic reference are much harder to be achieved. For this reason, it is important to discuss this gauge in some extent. Its use may be avoided when dealing with applications related to classical limits for instance, but it is more suited for renormalizability issues.
In some contexts, it is sometimes more convenient to express the action (2.1) and its equations of motion in terms of the Barnes- Rivers operators $P^{a}_{\mu\nu\alpha\beta}$, which are given in terms of the elementary tensors,
$\omega_{\mu\nu}=\frac{k_{\mu}k_{\nu}}{k^{2}},\qquad\theta_{\mu\nu}=\eta_{\mu\nu}-\frac{k_{\mu}k_{\nu}}{k^{2}},$ (2.14)
in the following way
$P^{2}_{\alpha\beta,\mu\nu}=\frac{1}{2}(\theta_{\beta\mu}\theta_{\alpha\nu}+\theta_{\beta\nu}\theta_{\alpha\mu})-\frac{1}{3}\theta_{\beta\alpha}\theta_{\mu\nu},$
$P^{1}_{\alpha\beta,\mu\nu}=\frac{1}{2}(\theta_{\beta\mu}\omega_{\alpha\nu}+\theta_{\beta\nu}\omega_{\alpha\mu}+\theta_{\alpha\mu}\omega_{\beta\nu}+\theta_{\alpha\nu}\omega_{\beta\mu}),$
$P^{0-s}_{\alpha\beta,\mu\nu}=\frac{1}{3}\theta_{\beta\alpha}\theta_{\mu\nu},\qquad P^{0-w}_{\alpha\beta,\mu\nu}=\omega_{\beta\alpha}\omega_{\mu\nu},$
$P^{0-sw}_{\alpha\beta,\mu\nu}=\frac{1}{\sqrt{3}}\theta_{\beta\alpha}\omega_{\mu\nu},\qquad P^{0-ws}_{\alpha\beta,\mu\nu}=\frac{1}{\sqrt{3}}\omega_{\beta\alpha}\theta_{\mu\nu}.$ (2.15)
These formulas may be expressed in the coordinate space by making the change
$\omega_{\mu\nu}=\frac{\partial_{\mu}\partial_{\nu}}{\square},\qquad\theta_{\mu\nu}=\eta_{\mu\nu}-\frac{\partial_{\mu}\partial_{\nu}}{\square},$
resulting in
$P_{\mu\nu\alpha\beta}^{2}=\frac{2}{3\square^{2}}\partial_{\mu}\partial_{\nu}\partial_{\alpha}\partial_{\beta}+\frac{1}{3\square}(\partial_{\mu}\partial_{\nu}\eta_{\alpha\beta}+\partial_{\alpha}\partial_{\beta}\eta_{\mu\nu})$
$\frac{1}{2\square}(\partial_{\mu}\partial_{\alpha}\eta_{\nu\beta}+\partial_{\mu}\partial_{\beta}\eta_{\nu\alpha}+\partial_{\nu}\partial_{\beta}\eta_{\mu\alpha}+\partial_{\alpha}\partial_{\nu}\eta_{\mu\beta})$
$\frac{1}{2}(\eta_{\mu\alpha}\eta_{\nu\beta}+\eta_{\nu\alpha}\eta_{\mu\beta})-\frac{1}{3}\eta_{\mu\nu}\eta_{\alpha\beta},$
$P_{\mu\nu\alpha\beta}^{1}=-\frac{1}{\square^{2}}\partial_{\mu}\partial_{\nu}\partial_{\alpha}\partial_{\beta}+\frac{1}{2\square}(\partial_{\mu}\partial_{\alpha}\eta_{\nu\beta}+\partial_{\mu}\partial_{\beta}\eta_{\nu\alpha}+\partial_{\nu}\partial_{\beta}\eta_{\mu\alpha}+\partial_{\alpha}\partial_{\nu}\eta_{\mu\beta}),$
$P_{\mu\nu\alpha\beta}^{0-s}=-\frac{1}{3\square}(\partial_{\alpha}\partial_{\beta}\eta_{\mu\nu}+\partial_{\mu}\partial_{\nu}\eta_{\alpha\beta})+\frac{1}{3}\eta_{\mu\nu}\eta_{\alpha\beta},\qquad P_{\mu\nu\alpha\beta}^{0-w}=\frac{1}{\square^{2}}\partial_{\mu}\partial_{\nu}\partial_{\alpha}\partial_{\beta},$
$P^{0-sw}_{\mu\nu\alpha\beta}=\frac{1}{\sqrt{3}\square}\eta_{\mu\nu}\partial_{\alpha}\partial_{\beta}-\frac{1}{\sqrt{3}\square^{2}}\partial_{\mu}\partial_{\nu}\partial_{\alpha}\partial_{\beta},$
$P^{0-ws}_{\alpha\beta,\mu\nu}=\frac{1}{\sqrt{3}\square}\eta_{\alpha\beta}\partial_{\mu}\partial_{\nu}-\frac{1}{\sqrt{3}\square^{2}}\partial_{\mu}\partial_{\nu}\partial_{\alpha}\partial_{\beta}.$ (2.16)
The definition of the Barnes-Rivers operators shows that
$(P^{2}+P^{1}+P^{0-2}+P^{0-w})_{\mu\nu}^{\;\;\;\alpha\beta}=\frac{1}{2}(\delta^{\alpha}_{\mu}\delta^{\beta}_{\nu}+\delta^{\alpha}_{\nu}\delta^{\beta}_{\mu}).$
In other words, the last sum is an identity. Therefore
$(P^{2}+P^{1}+P^{0-2}+P^{0-w})_{\mu\nu}^{\;\;\;\alpha\beta}h^{\mu\nu}=h^{\alpha\beta}.$ (2.17)
In addition the set composed by these four operators are such that
$P^{i-a}P^{j-b}=\delta^{ij}\delta^{ab}P^{j-b},$ (2.18)
where $i,j$ takes values between 0 and $2$ and $a,b$ takes values $s$ and $w$. Furthermore, the mixed operators $P^{0-ws}$ and $P^{0-sw}$ are such
$P^{0-ws}P^{0-sw}=P^{0-w},\qquad P^{0-sw}P^{0-ws}=P^{0-s},$
and their non zero actions with the other is
$P^{i-ab}P^{j-c}=\delta^{ij}\delta^{bc}P^{j-ac},\qquad P^{i-a}P^{j-bc}=\delta^{ij}\delta^{ab}P^{j-ac}.$
Note that the first four operators are projection operators, due to (2.18).
The terms of the linearized Stelle action (2.1) can be worked out in terms of the Barnes-Rivers operators as follows. Consider a typical term of the linearized action, mapped to momentum space
$h^{\alpha\beta}\eta_{\alpha\mu}\eta_{\beta\nu}\square^{2}h^{\mu\nu}\to k^{4}h^{\alpha\beta}\eta_{\alpha\mu}\eta_{\beta\nu}h^{\mu\nu}.$
By employing the definition (2.14), it is evident that $\eta_{\mu\nu}=\theta_{\mu\nu}+\omega_{\mu\nu}$, and with this simple identity the last term in momentum space may be worked out as
$k^{4}h^{\alpha\beta}\eta_{\alpha\mu}\eta_{\beta\nu}h^{\mu\nu}=k^{4}h^{\alpha\beta}(\theta_{\alpha\mu}\theta_{\beta\nu}+\theta_{\alpha\mu}\omega_{\beta\nu}+\omega_{\alpha\mu}\theta_{\beta\nu}+\omega_{\alpha\mu}\omega_{\beta\nu})h^{\mu\nu}$
$k^{4}h^{\alpha\beta}(P^{2}_{\alpha\beta,\mu\nu}+P^{0}_{\alpha\beta,\mu\nu}+P^{1}_{\alpha\beta,\mu\nu}+P^{0-w}_{\alpha\beta,\mu\nu})h^{\mu\nu}$
In the last identity, the definition of the Barnes Rivers operators (2.15) was employed. After going back to coordinate space, and doing the same with the other terms of the linearized action, the result is
$S=\frac{2}{\kappa^{2}}\int h^{\mu\nu}\square\bigg[\frac{1}{2}\bigg(1-\frac{\kappa^{2}\alpha}{4}\square\bigg)P^{2}-\bigg(1-\frac{\kappa^{2}(\alpha-3\beta)}{2}\square\bigg)P^{0-s}$
$\bigg(\frac{1}{2}-\frac{\kappa^{2}(4\beta-3\alpha)\square}{8}\bigg)P^{0-w}\bigg]_{\alpha\beta,\mu\nu}h^{\alpha\beta}d^{4}x.$ (2.19)
There are several identities satisfied for these operators, which have been worked out in the literature, for instance in [46]. One of the advantages of expressing the action in terms of the Barnes-Rivers operators is that, given an expression in terms of the Barnes-Rivers operators
$M=a_{2}P^{2}+a_{1}P^{1}+a_{s}P^{0-s}+a_{w}P^{0-w}+a_{sw}\sqrt{3}(P^{0-sw}+P^{0-ws}),$
its inverse is
$M^{-1}=\frac{1}{a_{2}}P^{2}+\frac{1}{a_{1}}P^{1}+\frac{1}{a_{s}a_{w}-3a_{sw}^{2}}\bigg[a_{w}P^{0-s}+a_{s}P^{0-w}-a_{sw}\sqrt{3}(P^{0-sw}+P^{0-ws})\bigg].$ (2.20)
Here the inverse means that
$M_{\mu\nu\gamma\delta}(M^{-1})^{\gamma\delta\alpha\beta}=(M^{-1})^{\gamma\delta\alpha\beta}M_{\mu\nu\gamma\delta}=\frac{1}{2}(\delta^{\alpha}_{\mu}\delta^{\beta}_{\nu}+\delta^{\alpha}_{\nu}\delta^{\beta}_{\mu}).$
The formula (2.20) shows the advantage of employing these operators. If one has an action written as
$S=\int h^{\mu\nu}(x)\hat{O}_{\mu\nu\alpha\beta}h^{\alpha\beta}(x)d^{4}x,$
with the operator $\hat{O}_{\mu\nu\alpha\beta}$ expressed in terms of the Barnes-Rivers operators, then (2.20) gives directly the propagator $\hat{D}_{\mu\nu\alpha\beta}=\hat{O}^{-1}_{\mu\nu\alpha\beta}$. Of course, there is no inverse for the operator in (2.19), since it is not gauge fixed and therefore non invertible. It is easy to see why it is not invertible. The coefficient for $P^{1}$ is zero and the inverse involves the inverse of the coefficient, which is infinite.
Now, a gauge has to be specified in order to obtain an invertible kinetic operator, leading to a non divergent propagtor. The gauge fixing term that the references [31] is the following
$L_{gf}=-\frac{\kappa^{2}}{2\lambda}F_{\tau}\square F^{\tau}.$ (2.21)
Here $F^{\tau}=\partial_{\nu}h^{\tau\nu}$. The use of this gauge may be advantageous in some applications and non practical in others. The advantage of it is that it leads to terms in the gauge fixed propagators which behave like $k^{-4}$ and which facilitate the renormalization program of that reference. In other applications, it may be convenient to employ the de Donder gauge, but the renormalizability of the model, although not spoiled, is considerably obscured.
The gauge fixed action in the Stelle gauge follows by expressing the gauge fixing part in terms of the Barnes-Rivers operators, the result is
$S=\frac{2}{\kappa^{2}}\int h^{\mu\nu}\square\bigg[\frac{1}{2}\bigg(1-\frac{\kappa^{2}\alpha}{4}\square\bigg)P^{2}-\bigg(1-\frac{\kappa^{2}(\alpha-3\beta)}{2}\square\bigg)P^{0-s}$
$\bigg(\frac{1}{2}-\frac{\kappa^{2}(4\beta-3\alpha)\square}{8}\bigg)P^{0-w}+\frac{\kappa^{4}}{8\lambda}\square[P^{1}-P^{0-w}]\bigg]_{\alpha\beta,\mu\nu}h^{\alpha\beta}d^{4}x.$
The inverse of the kinetic operator that can be read from this expression is the gauge fixed propagator, by employing (2.20)
$D_{\alpha\beta,\mu\nu}=\frac{1}{\square}\bigg[\frac{2}{1-\frac{\kappa^{2}\alpha}{4}\square}P^{2}+\frac{16\lambda}{\kappa^{4}\square^{2}}P^{1}-\frac{1}{1-\frac{\kappa^{2}}{2}(3\beta-\alpha)\square}P^{0-s}$
$\frac{2}{1+\frac{\kappa^{4}}{8\lambda}\square+\kappa^{2}(\beta-\frac{3}{4}\alpha)\square}P^{0-w}\bigg]_{\alpha\beta,\mu\nu}.$
It is seen that the gauge independent part namely, the one that corresponds to $\lambda\to 0$, involves three mass scales
$m_{1}^{2}=0,\qquad m_{2}^{2}=\frac{4}{\kappa^{2}\alpha},\qquad m_{3}^{2}=\frac{2}{\kappa^{2}(3\beta-\alpha)}.$ (2.22)
These are the mass scales anticipated in the previous section.
3. Quantization formulas for ghosts
3.1 The Pais-Uhlenbeck quartic oscillator
Before dealing with Quadratic gravity, consider the prototype of a ghost system, namely the Pais-Uhlenbeck lagrangian [44] for a real scalar mode $\phi$
$L=-\frac{3}{2}\partial_{\mu}\phi\partial^{\mu}\phi-\frac{9\gamma}{2}(\square\phi)^{2}-V(\phi).$
The potential $V(\phi)$ will be quite general in the present discussion. As the theory is of higher order, the procedure of doubling the number of variables can be applied, and there will be two coordinates and two momenta involved instead of one. The coordinates will be chosen as
$Q_{1}=\beta\phi,\qquad Q_{2}=\lambda\dot{\phi}.$ (3.23)
with $\beta$ and $\lambda$ constant factors. In the following the values $\beta=\lambda=1$ can be chosen, but these factors are written in order to specify certain liberty on the choice of coordinates and momenta which do not alter the canonical commutation relations. In fact, the quantization to be employed now may allow complex values for these constants and still give real mean values for these operators [29]. This method, maybe ackward at first sight, has roots from Dirac and Pauli times [6]-[7].
The canonical momenta corresponding to (3.23) can be found by the following formula, valid for a higher order formalism
$P_{1}=\frac{\partial L}{\partial\dot{Q_{1}}}-\frac{d}{dt}\bigg(\frac{\partial L}{\partial\ddot{Q_{1}}}\bigg)=\frac{1}{\beta}[3\dot{\phi}-9\gamma\square\dot{\phi}],$
$P_{2}=\frac{\partial L}{\partial\dot{Q_{2}}}-\frac{d}{dt}\bigg(\frac{\partial L}{\partial\ddot{Q_{2}}}\bigg)=\frac{9\gamma}{\lambda}\square\phi,$ (3.24)
while the Hamiltonian density is found by the standard prescription
$H=P_{1}\dot{Q}_{1}+P_{2}\dot{Q}_{2}-L.$
The above procedure ensures that the hamiltonian equations of motion are equivalent to the Euler-Lagrange ones. It is useful to deduce from the second (3.24) that
$\ddot{\phi}=\Delta\phi-\frac{\lambda}{9\gamma}P_{2}.$
The Hamiltonian density can be calculated explicitly in terms of the canonical coordinates (3.23) and momentum (3.24), the result is
$H=\frac{\beta P_{1}Q_{2}}{\lambda}-\frac{\lambda^{2}P_{2}^{2}}{18\gamma}+\frac{\lambda P_{2}\Delta Q_{1}}{\beta}-\frac{3}{2}\frac{Q_{2}^{2}}{\lambda^{2}}+\frac{3}{2\beta^{2}}(\nabla Q_{1})^{2}+V(Q_{1}).$ (3.25)
Here $\Delta$ is the three dimensional Laplacian.
The equation of motion of the model is
$3\square\phi+9\gamma\square^{2}\phi=\frac{\partial L_{i}}{\partial\phi},$ (3.26)
where the interaction lagrangian $L_{1}$ is due to the potential term $V(\phi)$. In the free case, the last equation shows that
$\phi=\phi_{1}+\phi_{2},$
the first mode is massless and the second has mass $m^{2}=\frac{1}{3\gamma}$. This means that the field can be Fourier expanded as
$\phi(x)=\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}\omega(k)}[b(k)e^{i\omega(k)t-ik\cdot x}+b^{\ast}(k)e^{-i\omega(k)t+ik\cdot x}]$
$\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}\omega_{2}(k)}[c(k)e^{i\omega_{2}(k)t-ik\cdot x}+c^{\ast}(k)e^{-i\omega_{2}(k)t+ik\cdot x}],$
where the corresponding dispersion relation is given by $\omega_{2}(k)=\sqrt{k^{2}+m^{2}}$. Also $\omega(k)=|k|$.
The quantization method requires the expressions of the mode coefficients $b(k)$, $b^{\ast}(k)$, $c(k)$ and $c^{\ast}(k)$ in terms of the canonical variables $P^{i}$ and $Q_{i}$ with $i=1,2$. Then, by assuming the standard equal time commutation relations given by
$Q_{i}(x,t),P_{j}(y,t)]=i\hbar\delta_{ij}\delta(x-y),$ (3.27)
the commutation relations for the quantum operators $b(k)$, $b^{\dagger}(k)$, $c(k)$ and $c^{\dagger}(k)$ can be deduced. This is a $4\times 4$ system of equations which, in principle, may be enough for determining those quantities.
This naive expectation however, may have a problem. It will be shown below, it is not possible to enforce this canonical equal time commutators (3.27) if the coordinates commute [42]
$Q_{i}(x,t),Q_{j}(y,t)]=0.$ (3.28)
In fact, it will be seen that the solution of the $4\times 4$ system is such that $Q_{i}(x,t),Q_{j}(y,t)]\neq 0$. In order to avoid this problem one may be tempted to modify one or more of the relations (3.27). If this approach is applied, it leads to apparent negative norm states.
The presence of negative norm states can be justified as follows. The second of the formulas (3.24) shows that this massless mode does not contribute to $P_{2}$ as $\square\phi_{1}=0$. Furthermore, from the first (3.24) and the definition $\dot{\phi}=\lambda Q_{2}$, it is seen that it does not contribute to $\beta P_{1}-\frac{3Q_{2}}{\lambda}$ either. Therefore, only the massive mode $\phi_{2}$ contributes to these two quantities and, since the massive mode involves only the coefficient $c(k)$ and $c^{\ast}(k)$, the system
$\lambda\beta P_{1}-3Q_{2}=-9\lambda\gamma\square\dot{\phi}_{2},\qquad P_{2}=\frac{9\gamma}{\lambda}\square\phi_{2},$ (3.29)
is a $2\times 2$ one allowing the calculation $c(k)$ and $c^{\ast}(k)$ in terms of the modes of $\lambda\beta P_{1}-3Q_{2}$ and $P_{2}$. This makes the problem more tractable. To solve this $2\times 2$ linear system explicitly, write
$\phi_{2}(x,t)=\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}\omega_{2}(k)}[c(k)e^{i\omega_{2}(k)t-ik\cdot x}+c^{\ast}(k)e^{-i\omega_{2}(k)t+ik\cdot x}].$
By taking into account that $\square\phi_{2}=m^{2}\phi_{2}$ with $m^{2}=\frac{1}{3\gamma}$ the equations (3.29) are converted into
$\lambda P_{2}(x,t)=3\phi_{2}(x,t)=3\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}\omega_{2}(k)}[c(k)e^{i\omega_{2}(k)t-ik\cdot x}+c^{\ast}(k)e^{-i\omega_{2}(k)t+ik\cdot x}],$
$3Q_{2}(x,t)-\lambda\beta P_{1}(x,t)=3\dot{\phi}_{2}=3i\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}}[c(k)e^{i\omega_{2}(k)t-ik\cdot x}-c^{\ast}(k)e^{-i\omega_{2}(k)t+ik\cdot x}].$ (3.30)
This can be inverted to give
$c(k)=\frac{1}{3}\int d^{3}xe^{ikx}[\omega_{2}(k)\lambda P_{2}(x,0)-3iQ_{2}(x,0)+i\lambda\beta P_{1}(x,0)],$
$c^{\dagger}(k)=\frac{1}{3}\int d^{3}xe^{-ikx}[\omega_{2}(k)\lambda P_{2}(x,0)+3iQ_{2}(x,0)-i\lambda\beta P_{1}(x,0)].$ (3.31)
If these expressions are prompted to operators $\widetilde{c}_{k}$ and $\widetilde{c}_{k}^{\dagger}$, by assuming the canonical commutation relations (3.27), it is arrived to
$\widetilde{c}_{k^{\prime}},\widetilde{c}^{\dagger}_{k}]=\frac{2\lambda\omega_{2}(k)}{3}\delta(k-k^{\prime}),$ (3.32)
This is the standard commutation relation, up to a factor $3\lambda^{-1}$. Consider now the remaining operators $b(k)$ and $b^{\dagger}(k)$. Their commutation relation can be determined from (3.23), under the assumption that $Q_{1},Q_{2}]=0$. This definition implies that
$Q_{1}=\beta\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}\omega(k)}[b(k)e^{i\omega(k)t-ik\cdot x}+b^{\ast}(k)e^{-i\omega(k)t+ik\cdot x}]$
$\beta\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}\omega_{2}(k)}[c(k)e^{i\omega_{2}(k)t-ik\cdot x}+c^{\ast}(k)e^{-i\omega_{2}(k)t+ik\cdot x}],$
$Q_{2}=i\lambda\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}}[b(k)e^{i\omega(k)t-ik\cdot x}-b^{\ast}(k)e^{-i\omega(k)t+ik\cdot x}]$
$i\lambda\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}}[c(k)e^{i\omega_{2}(k)t-ik\cdot x}-c^{\ast}(k)e^{-i\omega_{2}(k)t+ik\cdot x}].$ (3.33)
As these two quantities commute, if the unknowns $b_{k}$ and $b^{\ast}(k)$ are prompted to operators $\widetilde{b}_{k}$ and $\widetilde{b}_{k}^{\dagger}$ the commutation relation $\widetilde{b}_{k},\widetilde{b}_{k}^{\dagger}]$ has to cancel the terms coming from $c_{k},c_{k}^{\dagger}]$. This leads to
$b_{k},b_{k^{\prime}}^{\dagger}]=-\frac{2\lambda\omega_{2}(k)}{3}\delta(k-k^{\prime}).$ (3.34)
The minus sign in the last expression is the key result. It implies that one of the two pairs of creation/annihilation operators has a non standard 1-1 sign in the commutation relation in order to ensure that $Q_{1},Q_{2}]=0$. This non standard sign is unavoidable. The last commutation relation leads to an expression of $Q_{1},P_{1}]$ at equal times to have a minus sign in their definition, that is
$Q_{1}(x,t),P_{1}(y,t)]=-i\hbar\delta(x-y).$
One may declare a non standard minus sign in $Q_{2},P_{2}]$ instead, and the commutation $Q_{1},P_{1}]$ will be standard. Nevertheless, it is not possible to impose (3.27) for both pairs of canonical variables, at least if $Q_{1},Q_{2}]=0$.
Note that (3.33) and (3.30) imply that
$P_{1}=\frac{3}{\lambda\beta}i\lambda\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}}[b(k)e^{i\omega(k)t-ik\cdot x}-b^{\ast}(k)e^{-i\omega(k)t+ik\cdot x}].$ (3.35)
As $P_{2}$ in (3.30) only involves only $c(k)$ and $c^{\ast}(k)$, the wrong or correct sign with one of the commutators will not enter in conflict with the condition $P_{1},P_{2}]=0$, it only affects the commutator $Q_{1},Q_{2}]$.
The behavior of the energy levels of the system requires the analysis of its Hamiltonian. Its classical version is given explicitly in terms of $\phi$ by
$H=[3\dot{\phi}-9\gamma\square\dot{\phi}]\dot{\phi}+9\gamma\square\phi\ddot{\phi}+\frac{3}{2}\partial_{\mu}\phi\partial^{\mu}\phi+\frac{9\gamma}{2}(\square\phi)^{2}+V(\phi).$ (3.36)
This expression is obtained after some algebra from (3.27) by employing the formulas (3.24) in order to express everything in terms of $\phi$.
By decomposing $\phi=\phi_{1}+\phi_{2}$ and by taking into account their dispersion relations, it can be seen after some calculation, to be presented now, that the mixed terms involving $\phi_{1}$ and $\phi_{2}$ cancel out, and the resulting Hamiltonian density is simply the sum of two free Klein-Gordon ones corresponding to both modes.
The calculation leading to this conclusion goes as follows. After introducing the decomposition $\phi=\phi_{1}+\phi_{2}$ into (3.36) there appear terms that only involve $\phi_{1}$, other that involve $\phi_{2}$ and mixed terms. The mixed ones are
$H_{12}=[3\dot{\phi}_{1}-9\gamma\square\dot{\phi}_{1}]\dot{\phi}_{2}+9\gamma\square\phi_{1}\ddot{\phi}_{2}+3\partial_{\mu}\phi_{1}\partial^{\mu}\phi_{2}+9\gamma\square\phi_{1}\square\phi_{2}$
$3\dot{\phi}_{2}-9\gamma\square\dot{\phi}_{2}]\dot{\phi}_{1}+9\gamma\square\phi_{2}\ddot{\phi}_{1}.$
From the dispersion relation of the modes, it is clear that $\square\phi_{1}=0$ and $3\gamma\square\phi_{2}=\phi_{2}$. This leads to several simplifications, and the last expression becomes
$H_{12}=3\dot{\phi}_{1}\dot{\phi}_{2}+3\partial_{\mu}\phi_{1}\partial^{\mu}\phi_{2}+3\phi_{2}\ddot{\phi}_{1}.$
Collecting the first and the last term as a total derivative, it follows that
$H_{12}=3\partial_{t}(\dot{\phi}_{1}\phi_{2})-3\dot{\phi}_{1}\dot{\phi}_{2}+3\nabla\phi_{1}\cdot\nabla\phi_{2}.$
This can be equivalently written as
$H_{12}=3\phi_{2}\ddot{\phi}_{1}+3\nabla\phi_{1}\cdot\nabla\phi_{2}.$
This is the mixed term of the hamiltonian density. The full hamiltonian is integrated over the spatial volume $\int_{V}d^{3}x$. Inside the spatial volume integral, the last term is equivalent by parts integration to
$H_{12}=3\phi_{2}\ddot{\phi}_{1}-3\phi_{2}\Delta\phi_{1}.$
Since $\ddot{\phi}_{1}\sim-\omega^{2}\phi_{1}$ while $\Delta\phi_{1}\sim-k^{2}\phi_{1}$ inside the mode expansion, the dispersion relation for $\phi_{1}$ $\omega^{2}=k^{2}$ implies that the contribution of these terms cancel. This means that one may set the mixing $H_{12}=0$.
The resulting hamiltonian density is therefore composed only by the decoupled terms involving only $\phi_{1}$ or only $\phi_{2}$. The insertion of the expression $\phi=\phi_{1}+\phi_{2}$ into (3.36) throws the following
$H=\frac{3}{2}\dot{\phi}^{2}_{1}+\frac{3}{2}(\nabla\phi_{1})^{2}+3\phi_{2}\ddot{\phi}_{2}-\frac{3}{2}\dot{\phi}^{2}_{1}+\frac{3}{2}(\nabla\phi_{2})^{2}+\frac{3m^{2}}{2}(\phi_{2})^{2}.$
Inside the Fourier expansion $\ddot{\phi}_{2}=\dot{\phi}_{2}^{2}\sim-\omega_{2}^{2}\phi_{2}$, thus the last expression is equivalent to
$H=\frac{3}{2}\dot{\phi}^{2}_{1}+\frac{3}{2}(\nabla\phi_{1})^{2}+\frac{3}{2}\dot{\phi}^{2}_{2}+\frac{3}{2}(\nabla\phi_{2})^{2}+\frac{3m^{2}}{2}(\phi_{2})^{2}+V(\phi_{1},\phi_{2}).$
This clearly the sum of two free Klein-Gordon hamiltonian densities, for a massless and a massive particle, with an interaction $V(\phi_{1},\phi_{2})$. By omitting the zero point energy, the free part of the hamiltonian written in terms of the modes is given by
$H_{f}=\int d^{3}k[\omega(k)b^{\dagger}(k)b(k)+\omega_{2}(k)c^{\dagger}(k)c(k)].$ (3.37)
This expression would not raise any alarm if the creation and annihilation operators were standard. However, it was shown above that one of them may deviate from the standard commutation relation. Therefore it is needed to understand if this Hamiltonian is bounded or instead unbounded from below, if these non standard commutators are employed.
The problems described above may be understood in terms of path integrals. The Hamiltonian (3.27) depends only linearly on $P_{1}$. This raises a problem if one tries to quantize the model by the method of path integrals. If the coefficients $\beta$ and $\lambda$ are chosen equal to one, then the euclidean path version of this theory wlll be divergent, since the integration on the variable $P_{1}$ yields an infinite result. This divergence is clear, since the integral in $P_{1}$ involves a real exponential and the integrand will be unbounded. The fact that the Hamiltonian is linear in $P_{1}$ seems then problematic. This is not the case for other choices of these constants, for instance imaginary ones. These choices do not affect the canonical commutation relations $[P,Q]$ as they multiply $Q$ in inverse way as for $P$. The point is to understand the consequences of declaring that $\lambda$ or $\beta$ are imaginary  [^1] Related work is the reference [36], where apparently a self-consistent way to quantize these systems has been presented..
3.2 A toy quantum mechanical model
At this point it is instructive to study the non standard harmonic oscillator
$H=\hbar\omega u_{c}u_{a},\qquad[u_{a},u_{c}]=-1.$
The non standard nature of this oscillator is due to the factor 1-1 in the commutator, while in standard QM it is $1$. This situation is the QM analogous of the second term of (3.37). One may interpret these operators in terms of the standard creation and annihilation operators as
$u_{a}=ia,\qquad u_{c}=ia^{\dagger}$
These operators are not adjoint each other. In fact $u_{a}^{\dagger}=-u_{c}$. The standard states of an harmonic oscillator $|n>$ can be employed in order to find
$n|H|n>=-\hbar\omega n.$
The energy seems to be not bounded from below. However, the well known relation
$a=\frac{\hat{x}+i\hat{p}}{\sqrt{2}}.$
suggests that the definition $u_{a}=ia$ is equivalent to the operator redefinition $\hat{p}\to\hat{P}=i\hat{p}$ and $\hat{x}\to\hat{X}=i\hat{x}$. The Hamiltonian $H$ is quadratic in these quantities and implies the redefinition $H\to-H$.
Therefore, it is reasonable. to postulate that the mean values of these quantities are
$\overline{p}=-i<\hat{P}>,\qquad\overline{x}=-i<\hat{X}>,\qquad\overline{E}=-<H>.$
Already this is in concordance with the continuation procedure in [1]. Also for example
$\overline{p}^{n}=(-i)^{n}<\hat{P}^{n}>,\qquad\overline{x}^{n}=(-i)^{n}<\hat{X}^{n}>$
$\overline{p^{n}x^{m}}=(-i)^{n+m}<\hat{P}^{n}\hat{X}^{m}>$
For a generic observable $O(P,X)$ constructed in powers of $\hat{X}$ and $\hat{P}$ these rules give the corresponding mean values. Calculate the mean value <O(P,X)><O(P,X)> and analitically continue to values $P=\to-iP$ and $X\to-iX$. The energy mean values are now positive, and the impulse and coordinates mean values are real..
Another subtle detail is related to probabilities. The basis $|n>$ employed above is the standard one. Then <n|n>=1<n|n>=1, which is perfectly fine. However, if this state is expressed in terms of the creation operators, then
$|n>\neq\frac{1}{\sqrt{n!}}u_{c}^{n}|0>,$
In other words, the standard expression in QM does not hold here. Instead
$|n>=\frac{(-i)^{n}}{\sqrt{n!}}u_{c}^{n}|0>.$ (3.38)
The reason for this formula is simple. The standard oscillator states $|n>$ are created in terms of $a^{\dagger}=-iu_{c}$, and this leads to the conclusion. This distinction will be important in the derivation of the LSZ formulas. In fact, consider the scalar field of the previous section $\phi=\phi_{1}+\phi_{2}$. The interaction potential may generate terms of the form $\phi_{1}^{2}\phi_{2}^{3}$. If one calculates $V(\phi)>$, then terms will appear such as
$\phi_{1}^{3}\phi_{2}^{3}>$
As the ghost part $\phi_{2}$ is constructed with creation and annihilation operators which has an imaginary factor $i$, this will lead to an imaginary value of the last expression. On the other hand, it is reasonable to expect the mean value of a real scalar field to be real. Therefore, the above prescription is consistent if the mean values are calculated, for example, as follows
$\overline{\phi_{1}^{3}\phi_{2}^{3}}=(-i)^{3}<\phi_{1}^{3}\phi_{2}^{3}>.$ (3.39)
In general, in order to calculate mean values of a given observable, one must make the replacement $\phi=\phi_{1}+\phi_{2}\to\phi_{1}-i\phi_{2}$, with the 2 index denoting the ghost variable.
The prescription of changing $P_{2}$ and $Q_{2}$ to $iP_{2}$ and $iQ_{2}$ in the Hamiltonian (3.37) changes the unwanted negative sign in the ghost contribution, and the energy becomes bounded from below. This is the desired feature. Also, the fact that the states employed are the standard one do not enter in conflict with negative probabilities.
3.3 The path integral formulation
The path integral formulation of Quantum Mechanics in the presence of both standard $x$ and ghost like $X$ variables is obtained by considering the transition amplitude [13], [4]-[29], [1]-[2]
$x_{f},X_{f},t_{f}|x_{i},X_{i},t_{i}>=\prod_{l=1}^{N}\int[dx_{l}][dX_{l}]<x_{f},X_{f},t_{f}|x_{N},X_{N},t_{N}>$...
$x_{1},X_{1},t_{1}|x_{i},X_{i},t_{i}>,$
this results follows by successively inserting identities
$1=\int|x,X,t><x,X,t|dxdX.$
In addition
$x_{l+1},X_{l+1},t_{l+1}|x_{l},X_{l},t_{l}>=\int<x_{l+1},X_{l+1},t_{l+1}|p_{l+1},P_{l+1},t_{l+1}>$
$p_{l+1},P_{l+1},t_{l+1}|x_{l},X_{l},t_{l}>dp_{l}dP_{l}.$
From the expression of the evolution at small $\delta t_{l}=t_{l+1}-t_{l}$ given in terms of the system‘s Hamiltonian by
$p_{l+1},P_{l+1},t_{l+1}|x_{l},X_{l},t_{l}>=<p_{l+1},P_{l+1},t_{l}|e^{-iH\delta t}|x_{l},X_{l},t_{l}>$
$\frac{1}{(2\pi)^{m}}\exp\{-iH(p_{l},x_{l},P_{l},X_{l})\delta t_{l}-i(p_{l}x_{l}+P_{l}X_{l})\},$
where $m$ is the number of degrees of freedom and $H(p_{l},x_{l},P_{l},X_{l})$ is the mean value
$H_{a}(p_{l},x_{l},P_{l},X_{l})=\frac{<p_{l},P_{l},t_{l}|H(p_{l},x_{l},P_{l},X_{l})|x_{l},X_{l},t_{l}>}{<p_{l},P_{l},t_{l}|x_{l},X_{l},t_{l}>},$
An important point is that the above mean value has to be constructed taking into account that the eigenvalues of $P_{i}$ and $X_{i}$ are imaginary. Therefore, at the classical level it corresponds to put imaginary values $iP_{l}$ and $iX_{l}$. The minus in the momentum is due to the fact that <P|<P| is on the left when calculating the mean value, and that $P$ is skew hermitian. Collecting all these results, it follows that
$x_{f},X_{f},t_{f}|x_{i},X_{i},t_{i}>=\int[Dx][DX][Dp][DP]\exp\bigg\{i\int_{t_{i}}^{t_{f}}dt[p\dot{x}+P\dot{X}-H_{a}(p,x,P,X)]\bigg\}$
The last expression imitates usual path integral in QM, however with the subtle detail that the Hamiltonian is continued to imaginary values of the ghost variables. This may generate real divergent exponentials, and the integral may not exist. This existence issue has to be analyzed model by model. Fortunately, in the present case, this unpleasant feature does not take place [1]-[2], [4]-[29].
In order to see that the path integral does not have the apparent divergencies problem, note that the Hamiltonian (3.27), which it written again here by convenience after choosing $\beta=\lambda=1$
$H=P_{1}Q_{2}-\frac{P_{2}^{2}}{18\gamma}+P_{2}\Delta Q_{1}-\frac{3}{2}Q_{2}^{2}+\frac{3}{2}(\nabla Q_{1})^{2}+V(Q_{1})$ (3.40)
only depends on $P_{1}$ by a linear term, which in the path integral may lead to a divergent exponential. However, under the prescription of continue the values of $Q_{2}\to iQ_{2}$ and $P_{2}\to-iP_{2}$ it becomes
$H=iP_{1}Q_{2}+\frac{P_{2}^{2}}{18\gamma}-iP_{2}\Delta Q_{1}+\frac{3}{2}Q_{2}^{2}+\frac{3}{2}(\nabla Q_{1})^{2}+V(Q_{1}).$ (3.41)
The dangerous term now is converted in a harmless imaginary term $iP_{1}Q_{2}$ whose exponentiation and integration does not lead to divergent results. To see this note that the euclidean path integral is
$Z_{e}(0)=\int DP^{1}D\phi_{1}DP^{2}D\phi_{2}\exp\bigg\{-\int[iP_{1}\dot{Q}_{1}+iP_{2}\dot{Q}_{2}-H]dVd\tau\bigg\}$
$\int DP^{1}D\phi_{1}DP^{2}D\phi_{2}$
$\exp\bigg\{-\int[iP_{1}\dot{\phi}_{1}+iP_{2}\dot{\phi}_{2}-iP_{1}Q_{2}-\frac{P_{2}^{2}}{18\gamma}+iP_{2}\Delta Q_{1}-\frac{3}{2}Q_{2}^{2}-\frac{3}{2}(\nabla Q_{1})^{2}]dVd\tau\bigg\}$
The integration over $P_{1}$ leads to a Dirac delta that enforces $Q_{2}=-Q^{\prime}_{1}$, where the prime indicates derivative with respect to the euclidean time. The path integral now is related to $D\dot{Q}_{1}$ due to this identification, but a change to the variable $DQ_{1}$ may be done by multiplying the integration measure by $\prod_{x_{E^{\prime}}}\Delta\tau^{-1}$. The integral over $P_{2}$ is a Gaussian integral after setting $Q_{2}=-Q^{\prime}_{1}$. The general formula
$\int d^{N}xe^{x^{T}Ax+bx}=\frac{\pi^{\frac{N}{2}}}{\sqrt{\det A}}e^{-\frac{1}{4}b^{T}Ab},$
can be employed to make this integral. After some calculation, it is found that the result with currents $J$ turned on given by
$Z_{e}(J)=\int D\phi\exp\bigg\{-\int L_{e}dVd\tau+\int J\phi dVd\tau\bigg\}.$ (3.42)
Here $L_{e}$ is the euclidean version of the model corresponding to the action
$S=\int d\tau\int d^{3}x[\frac{3}{2}\partial_{\mu}\phi\partial_{\mu}\phi+\frac{9\gamma}{2}(\partial_{\mu}\partial_{\mu}\phi)^{2}+V(\phi)].$
Therefore, the apparent divergence which can invalidate the model is avoided with this approach. The correlation functions can be calculated in the euclidean formalism and analitically continued to the Lorenzian setting, by going from imaginary $\tau$ to real $t=i\tau$ times.
A further comment about the literature about the subject is mandatory. The prescription given above, which requires to continue the values of the coordinates and momentum to imaginary values, is inspired in some deep works, for instance [7]-[30]. I believe that at the end it is equivalent to the prescription of [1]. Although the procedure may look naive, it has it roots in works of Dirac and Pauli, and was extensively studied also in [13]. The main difference with these approaches is to put the imaginary unit $i$ not in the creation and annihilation operators, but in the states themselves. In other words, these references work with modified states $|n>$ whose norm is not insured to be positive due to the inclusion of $i$ unit. The norm of these states $\eta_{nm}=<n|m>$ is has negative entries. However, the mixed index metric $\eta_{n}^{m}=\delta_{n}^{m}$ is positive definite. These references distinguish between covariant, contravariant and mixed Hilbert space metrics. The mixed one is adapted to describe probabilities [13], [4]-[29], [1]. The mean values can be given by taking into account the covariant or contravariant metric. This is the formalism employed in [42]. Our conjecture is that the prescription presented here will lead, at least in these models, the same or analogous results than those pioneer references.
4. LSZ reduction for ghost variables
Given the previous discussion, consider again the $S$ matrix for a given process involving the scalar $\phi$. The action on the creation operator for a state $b_{in}^{\dagger}(k)$ of momentum $k$ is defined by
$S^{\dagger}b^{\dagger}_{in}(k)S=(\sum_{\text{all states}}|r,\text{out}><r,\text{in}|)(\sum_{\text{all states}}c_{k}\sqrt{n(k)+1}|h+k,\text{in}><h,\text{out}|)$
$\sum_{\text{all states}}c_{k}\sqrt{n(k)+1}|h+k,\text{out}><h,\text{out}|.$ (4.43)
Here $n(k)$ are the number of quanta of momentum $k$ that were already present in the state $h$, and $h+k$ denotes the state with one of such quanta added. A constant $c_{k}$ is included, in order to anticipate that these operators may not be standard.
The following analysis repeats the standard one adapted to a theory of second order to one of fourth order, with ghost included. The idea is to understand in which part of the deduction the fact that the creation or annihilation operators are non standard may affect the calculations. Therefore, although the analysis may look straightforward, it is necessary to work out all the details in order to be sure what are the consequences, if any, of the non standard creation and annihilation nature of the operators.
Clearly, the last expression represents $b^{\dagger}_{out}(k)$, which leads to the conclusion that
$S^{\dagger}b^{\dagger}_{in}(k)S=b^{\dagger}_{out}(k),\qquad S^{\dagger}b_{in}(k)S=b_{out}(k),$
the second relation is completely analogous to the first one. The value of the constant $c_{k}$ parameterizing the non standard nature of the operator is irrelevant in order to obtain this conclusion. The same relation follows for $c(k)$ and $c^{\dagger}(k)$. Therefore, it is deduced that
$S^{\dagger}\phi_{in}(k)S=\phi_{out}(k),$ (4.44)
regardless the non standard commutation of the creation and annihilation operators.
The use of the standard Fock space ensures that, at this point, the matrix $S$ in (4.43) is unitary.
Before deducing the LSZ formula, it is mandatory to deduce some classical relations for the fields in the model. The first is related to the equations of motion of $\phi(x)$ is given by
$\hat{O}\phi_{1}=c\frac{\partial L_{int}}{\partial\phi},$ (4.45)
where $L_{int}$ contains the infinite vertices involving the interaction of $\phi_{1}$ with the rest of the fields. The operator $\hat{O}$, corresponding to the equations of motion of the scalar field of the previous section, is a quartic operator
$\hat{O}=\square^{2}-\frac{1}{3\gamma}\square,$
up to a constant. That is the reason for adding a constant $c$ in the interaction term. The following discussion follows closely the methods of [158]. The kernel of the last operator, that is, a Green function for $\hat{O}$ is given by
$\hat{O}_{y}G(y-x)=\delta(y-x).$
It is clear that
$G(x-y)=\frac{1}{\hat{O}_{y}}=\frac{1}{\square^{2}-\frac{1}{3\gamma}\square}=\bigg(\frac{1}{\square}-\frac{1}{\square-\frac{1}{3\gamma}}\bigg)\frac{1}{3\gamma}.$
Therefore the Green function G(xy)G(x-y) is given by
$G(x-y)=\frac{1}{3\gamma}[G^{m=0}(x-y)-G^{m}(x-y)],$
being $G^{m}(x-y)$ any of the standard Green functions for a massive scalar field. There are plenty of such Green functions since the addition of a homogeneous solution of the wave operator leads to a new Green function. In terms of any of these Geen kernel, the following integral relation for $\phi_{1}$
$\int[\phi_{1}(y)\hat{O}_{y}G(y-x)-G(y-x)\hat{O}_{y}\phi_{1}(y)]d^{4}y=\phi_{1}(x)-c\int G(y-x)\frac{\partial L_{int}}{\partial\phi(y)}d^{4}y,$ (4.46)
is found. By assuming that $\phi_{1}$ vanish fast enough at the spatial infinite, it is concluded that the spatial part of the left hand side is related to two types of integrals. The first is related to the laplacian $\Delta_{y}$ in three dimensions, appearing in $\square_{y}$,. This contribution is
$\int[\phi_{1}(y)\Delta_{y}G(y-x)-G(y-x)\Delta_{y}\phi_{1}(y)]d^{3}ydy_{0}$
$\int[\phi_{1}(y)\nabla_{y}G(y-x)-G(y-x)\nabla_{y}\phi_{1}(y)]dS_{y}dy_{0}.$ (4.47)
The right hand side is obtained from the left by noticing that the left integrand is the gradient of $\phi_{1}(y)\nabla_{y}G(y-x)-G(y-x)\nabla_{y}\phi_{1}(y)$. If everything vanish in the infinite surface $S_{y}$ this gives no contributions. The other spatial contribution, due to the terms with $\square_{y}^{2}$, are given by
$\int[\phi_{1}(y)\Delta_{y}\Delta_{y}G(y-x)-G(y-x)\Delta_{y}\Delta_{y}\phi_{1}(y)]d^{3}ydy_{0}=0,$
after integration by parts. There are also terms in $\square^{2}_{y}$ mixing spatial and time derivatives. These are of the form
$\int[\phi_{1}(y)\Delta_{y}\partial^{2}_{y_{0}}G(y-x)-G(y-x)\Delta_{y}\partial^{2}_{y_{0}}\phi_{1}(y)]d^{3}ydy_{0}$
$\int\partial_{y_{0}}[\phi_{1}(y)\Delta_{y}\partial_{y_{0}}G(y-x)-G(y-x)\Delta_{y}\partial_{y_{0}}\phi_{1}(y)]d^{3}ydy_{0}$
$\int[\partial_{y_{0}}\phi_{1}(y)\Delta_{y}\partial_{y_{0}}G(y-x)-\partial_{y_{0}}G(y-x)\Delta_{y}\partial_{y_{0}}\phi_{1}(y)]d^{3}ydy_{0}.$
The last two terms cancel by a same argument as (4.47). Therefore the last equality can be worked out as follows
$\int[\phi_{1}(y)\Delta_{y}\partial^{2}_{y_{0}}G(y-x)-G(y-x)\Delta_{y}\partial^{2}_{y_{0}}\phi_{1}(y)]d^{3}ydy_{0}$
$\int\partial_{y_{0}}[\phi_{1}(y)\Delta_{y}\partial_{y_{0}}G(y-x)-G(y-x)\Delta_{y}\partial_{y_{0}}\phi_{1}(y)]d^{3}ydy_{0}$
$\int\partial_{y_{0}}[\Delta_{y}\phi_{1}(y)\partial_{y_{0}}G(y-x)-G(y-x)\Delta_{y}\partial_{y_{0}}\phi_{1}(y)]d^{3}ydy_{0}.$ (4.48)
The unique spatial contribution is then (4.48). The pure time components that come from $\square_{y}$ are of the form
$\int[\phi_{1}(y)\partial^{2}_{y_{0}}G(y-x)-G(y-x)\partial^{2}_{y_{0}}\phi_{1}(y)]d^{4}y$
$\int\partial_{y_{0}}[\phi_{1}(y)\partial_{y_{0}}G(y-x)-G(y-x)\partial_{y_{0}}\phi_{1}(y)]d^{4}y$
$\bigg(\int_{y^{-}_{0}}-\int_{y^{+}_{0}}\bigg)G(y-x)\overleftrightarrow{\partial}_{y_{0}}\phi_{1}(y)d^{3}y.$ (4.49)
The first two integrals are three dimensional, the first is done at an spatial surface corresponding to the time $y_{0}^{-}$at the past of $x_{0}$ and the second at another time $y_{0}^{+}$ at the future of $x_{0}$. The last integral is four dimensional, and is performed on the $4$-volume delimited by these two three dimensional surfaces. The last pure time contribution is due to the time part of $\square_{y}^{2}$. It is, and can be worked out as
$\int[\phi_{1}(y)\partial^{4}_{y_{0}}G(y-x)-G(y-x)\partial^{4}_{y_{0}}\phi_{1}(y)]d^{4}y$
$\int\partial_{y_{0}}[G(y-x)\partial^{3}_{y_{0}}\phi_{1}(y)-\phi_{1}(y)\partial^{3}_{y_{0}}G(y-x)]d^{4}y$
$\int\partial_{y_{0}}[\partial^{2}_{y_{0}}G(y-x)\partial_{y_{0}}\phi_{1}(y)-\partial^{2}_{y_{0}}\phi_{1}(y)\partial_{y_{0}}G(y-x)]d^{4}y$
$\bigg(\int_{y^{-}_{0}}-\int_{y^{+}_{0}}\bigg)[G(y-x)\overleftrightarrow{\partial}_{y_{0}}\partial_{y_{0}}^{2}\phi_{1}(y)+\partial_{y_{0}}^{2}G(y-x)\overleftrightarrow{\partial}_{y_{0}}\phi_{1}(y)]d^{3}y.$ (4.50)
By collecting all the above formulas (4.47)-(4.50), the expression (4.46)t becomes
$\phi_{1}(x)=-\bigg(\int_{y^{-}_{0}}-\int_{y^{+}_{0}}\bigg)G(y-x)\overleftrightarrow{\partial}_{y_{0}}\phi_{1}(y)d^{3}y$
$\frac{1}{3\gamma}\bigg(\int_{y^{-}_{0}}-\int_{y^{+}_{0}}\bigg)[G(y-x)\overleftrightarrow{\partial}_{y_{0}}\partial_{y_{0}}^{2}\phi_{1}(y)+\partial_{y_{0}}^{2}G(y-x)\overleftrightarrow{\partial}_{y_{0}}\phi_{1}(y)]d^{3}y$
$c\int_{y_{0}^{-}}^{y_{0}^{+}}G(y-x)\frac{\partial L_{int}}{\partial\phi(y)}d^{4}y.$
At the present point G(yx)G(y-x) can be any Green function, defined by any possible boundary condition. Particularly useful examples are the advanced $\Delta_{a}(x)$ or retarded $\Delta_{r}(x)$ Green functions, which vanish for $x_{0}>0$ and $x_{0}<0$, respectively, and vanish outside the light cone. If these functions are employed, then the two spatial integrals in the last formulas are reduced into one. For instance
$\phi_{1}(x)=\int_{y^{-}_{0}}\Delta_{r}(y-x)\overleftrightarrow{\partial}_{y_{0}}\phi_{1}(y)d^{3}y$
$c\int_{y^{-}_{0}}[\Delta_{r}(y-x)\overleftrightarrow{\partial}_{y_{0}}\partial_{y_{0}}^{2}\phi_{1}(y)+\partial_{y_{0}}^{2}\Delta_{r}(y-x)\overleftrightarrow{\partial}_{y_{0}}\phi_{1}(y)]d^{3}y+c\int\Delta_{r}(y-x)\frac{\partial L_{int}}{\partial\phi(y)}d^{4}y.$
The first term in the last formula has a simple interpretation. This term is such that
$\hat{O}_{x}\int_{y^{-}_{0}}\Delta_{r}(y-x)\overleftrightarrow{\partial}_{y_{0}}\phi_{1}(y)d^{3}y-\hat{O}_{x}\int_{y^{-}_{0}}[\Delta_{r}(y-x)\overleftrightarrow{\partial}_{y_{0}}\partial_{y_{0}}^{2}\phi_{1}(y)+\partial_{y_{0}}^{2}\Delta_{r}(y-x)\overleftrightarrow{\partial}_{y_{0}}\phi_{1}(y)]d^{3}y$
$\int_{y^{-}_{0}}\delta(y-x)\overleftrightarrow{\partial}_{y_{0}}\phi_{1}(y)d^{3}y-\int_{y^{-}_{0}}[\delta(y-x)\overleftrightarrow{\partial}_{y_{0}}\partial_{y_{0}}^{2}\phi_{1}(y)+\partial_{y_{0}}^{2}\delta(y-x)\overleftrightarrow{\partial}_{y_{0}}\phi_{1}(y)]d^{3}y$
In the limit $y_{0}^{-}\to$-$\infty$ the last integral vanishes if $x_{0}$ is fixed, since this point will be always at the future of $y_{0}$. Therefore this particular quantity is killed by the action of $\hat{O}_{x}$ and may be identified as a free field, the incoming field. The previous formula, by taking into account the equations of motion (4.45), may be expressed after taking this limit as
$\phi_{1}(x)=\phi_{1in}(x)+c\int\Delta_{r}(y-x)\frac{\partial L_{int}}{\partial\phi(y)}d^{4}y,$
or by use of the equations of motion
$\phi_{1}(x)=\phi_{1in}(x)+c\int\Delta_{r}(x-y)\hat{O}_{y}\phi_{1}(y)d^{4}y.$ (4.51)
Here, the second integral can be done in the full space time volume, as the past time is taken to -$\infty$ and the retarded Green function vanishes at the future.
The last two formulas make sense, as this field is determined by time derivatives up to order three, consistent with a four order equation of motion.
Analogously
$\phi_{1}(x)=\phi_{1iut}(x)+c\int\Delta_{a}(y-x)\frac{\partial L_{int}}{\partial\phi(y)}d^{4}y,$
or
$\phi_{1}(x)=\phi_{1out}(x)+c\int\Delta_{a}(x-y)\hat{O}_{y}\phi_{1}(y)d^{4}y.$ (4.52)
All the above discussion is purely classical, at the moment this solution has not been prompted to operators.
Consider now the operator version of these classical identities. From the mode expansion
$\phi_{in}=\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}\omega(k)}[b(k)e^{i\omega(k)t-ik\cdot x}+b^{\ast}(k)e^{-i\omega(k)t+ik\cdot x}]$
$\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}\omega_{2}(k)}[c(k)e^{i\omega_{2}(k)t-ik\cdot x}+c^{\ast}(k)e^{-i\omega_{2}(k)t+ik\cdot x}],$
it is seen that, due to the non standard oscillator algebra of $b(k)$ and $b^{\dagger}(k)$
$\phi_{1in}(x),\phi_{1in}(y)]=\int\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}\omega_{1}(k)}\frac{d^{3}l}{(2\pi)^{\frac{3}{2}}\omega_{1}(l)}\bigg[[b(k),b^{\dagger}(l)]e^{-i\omega(k)x_{0}+ik\cdot x}e^{i\omega(l)y_{0}-il\cdot y}$
$b(l),b^{\dagger}(k)]e^{i\omega(k)x_{0}-ik\cdot x}e^{-i\omega(l)y_{0}+il\cdot y}\bigg].$
$\int\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}\omega_{2}(k)}\frac{d^{3}l}{(2\pi)^{\frac{3}{2}}\omega_{2}(l)}\bigg[[c(k),c^{\dagger}(l)]e^{-i\omega(k)x_{0}+ik\cdot x}e^{i\omega(l)y_{0}-il\cdot y}$
$c(l),c^{\dagger}(k)]e^{i\omega(k)x_{0}-ik\cdot x}e^{-i\omega(l)y_{0}+il\cdot y}\bigg].$
$i\int\frac{d^{3}k}{(\pi)^{\frac{3}{2}}\omega_{1}(k)}e^{ik\cdot(x-y)}\sin(\omega_{1}(k)(x_{0}-y_{0}))-i\int\frac{d^{3}k}{(\pi)^{\frac{3}{2}}\omega_{2}(k)}e^{ik\cdot(x-y)}\sin(\omega_{2}(k)(x_{0}-y_{0})).$
This is a linear combination of two propagators for a massless and massive particle, but one of them has a non standard minus sign in front. Each of these propagators are known to be written as a difference of a retarded and advanced Green functions
$\Delta^{m}(x-y)=\Delta^{m}_{a}(x-y)-\Delta^{m}_{r}(x-y),$ (4.53)
and the analogous formula holds for the massless case, with dispersion $\omega_{1}(k)$. In other words, by denoting the last commutator as $i\Delta(x-y)$ it is clear that
$\phi_{1in}(x),\phi_{1in}(y)]=<0|[\phi_{1in}(x),\phi_{1in}(y)]|0>=i\Delta(x-y)=i\Delta_{a}(x-y)-i\Delta_{r}(x-y)$
$i\Delta^{m=0}_{a}(x-y)-i\Delta^{m=0}_{r}(x-y)-i\Delta^{m}_{a}(x-y)+i\Delta^{m}_{r}(x-y),$
two propagators are due to the massless mode and two due to the massive mode. The first identity is due to the fact that the commutator is a c-number. The last expression allows to infer that
$0|[\phi_{1in}(x),\phi_{1in}(y)]|0>=<0|[\phi_{1}(x),\phi_{1}(y)]|0>.$ (4.54)
This is seen as follows. The last identity, deduced from (4.51) is true up to a term proportional to
$0|[\phi_{in}(x),\hat{O}_{y}\phi(y)]|0>.$
Therefore, the identity will be valid if this term is zero. It is, since
$0|\hat{O}_{y}\phi_{1}(y)|p>=\hat{O}_{y}e^{ipy}<0|\phi(0)|p>=p^{2}(1-(3\gamma)^{-1}p^{2})e^{ipy}<0|\phi(0)|p>=0.$
The last identity is due to the fact that the modes satisfy the dispersion relation $p^{2}(1-(3\gamma)^{-1}p^{2})=0$. Due to this property the insertion of the identity $\int dp|p><p|$ in this extra piece gives vanishing result and the desired identity is proved. In addition, the value of $\phi_{1}(x),\phi_{1}(y)]$ is known to be a c-number [158] and therefore (4.54) may be replaced by
$\phi_{1in}(x),\phi_{1in}(y)]=[\phi_{1}(x),\phi_{1}(y)]=i\Delta(x-y),$ (4.55)
in other words, the vacuum expectation value may be deleted.
By collecting all the above information, consider the fundamental operator for scattering in QFT namely
$Z(J)=<0|I(J)|0>,\qquad I(J)=Te^{i\int\phi(x)J(x)dx},$
the last quantity satisfies
$i\frac{\delta I(J)}{\delta J(x)}=T(\phi(x)I(J)).$
The last formula combined with formulas (4.51) and (4.52) leads to the following formulas
$i\frac{\delta I(J)}{\delta J(x)}=I(x)\phi_{1in}(x)+i\int\Delta_{r}(x-y)\hat{O}_{y}\frac{\delta I(J)}{\delta J(y)}d^{4}y,$ (4.56)
$i\frac{\delta I(J)}{\delta J(x)}=\phi_{1out}(x)I(x)+i\int\Delta_{a}(x-y)\hat{O}_{y}\frac{\delta I(J)}{\delta J(y)}d^{4}y.$ (4.57)
From the two equations (4.56) and (4.57) it is arrived to
$\phi_{1out}(x)I(x)-\phi_{1in}(x)I(x)=-i\int\Delta(x-y)\hat{O}_{y}\frac{\delta I(J)}{\delta J(y)}d^{4}y,$
and this, together with the action of the $S$ matrix defined in (4.44) leads to the following commutator
$\phi_{1in}(x),SI(J)]=i\int\Delta(x-y)\hat{O}_{y}\frac{\delta SI(J)}{\delta J(y)}d^{4}y.$ (4.58)
In addition, it has been shown in (4.55) that the commutator of the fields is a c-number. By applying the following consequence of the Bakker-Campbell formula, which is true only for commutators which are c-numbers,
$A,e^{B}]=[A,B]e^{B},$
it can be seen that the general solution of the last equation (4.58) is
$SI(J)=e^{\int\phi_{1in}(z)\hat{O}_{z}\frac{\delta}{\delta J(z)}d^{4}z}F(J),$
with $F(J)$ arbitrary. The proof is found from the version of the Baker Campbell formula given above. To visualize this statement, note that formula and the last trial function imply
$\phi_{1in}(x),SI(J)]=\bigg[\int[\phi_{1in}(x),\phi_{1in}(z)]\hat{O}_{z}\frac{\delta}{\delta J(z)}d^{4}z\bigg]e^{-\int\phi_{1in}(y)\hat{O}_{y}\frac{\delta}{\delta J(y)}d^{4}y}F(J)$
$ie^{-\int\phi_{1in}(y)\hat{O}_{y}\frac{\delta}{\delta J(y)}d^{4}y}\bigg[\int\Delta(x-z)\hat{O}_{z}\frac{\delta F(J)}{\delta J(z)}d^{4}z\bigg].$
This characterizes the left hand side of the equation (4.58) The right hand side is
$i\int\Delta(x-y)\hat{O}_{y}\frac{\delta SI(J)}{\delta J(y)}d^{4}y=-i\int\Delta(x-y)\hat{O}_{y}e^{-\int\phi_{1in}(z)\hat{O}_{z}\frac{\delta}{\delta J(z)}d^{4}z}\frac{\delta F(J)}{\delta J(y)}d^{4}y$
The last two formulas are equal, and this shows that (4.58) is satisfied for every $F(J)$. This unknown function is determined by taking into account that, for flat Minkowski space, $S|0>=|0>$, and that
$0|e^{A}|0>=<0|:e^{A}:|0>=1,$
for any operator $A$. In these terms
<0|IS(J)|0>=<0|F(J)|0>=<0|I(J)|0>=Z(J).<0|IS(J)|0>=<0|F(J)|0>=<0|I(J)|0>=Z(J).
In the last step it was assumed that there are no external fields, and that this implies $S|0>=|0>$. In presence of such fields, this hypothesis has to be revised.
Therefore, it is concluded that the scattering operator is given by
$S=:e^{\int\phi_{1in}(z)\hat{O}\frac{\delta}{\delta J(z)}d^{4}z}:Z(J)\bigg|_{J=0},$ (4.59)
The LSZ reduction formula is then as in standard QFT, but the wave operator is replaced by the quartic wave one $\hat{O}$. The propagator $\Delta(x-y)$ has disappeared in the left hands side of the final formula, although it is inside the path integral $Z(J)$. As in (3.42), there is not a modification of $\phi$ in all this approach. All the calculation for scattering amplitudes is consistent with the field $\phi$ unmodified. However, if this reduction formula is expanded, it will lead to non standard results due to the non standard commutation relation.
It is needed to clarify how to operate with the last formula. However, this will be discussed in detail in the next sections, in a model which is more interesting namely, Quadratic Gravity.
5. The full Stelle action in synchronous gauge
The previous ideas will be applied first to Stelle gravity [31]-[32] in the synchronous gauge. One of the main results of reference [1] is that the quantization presented above also cure the behavior of $Z(J)$ for this theory. Here the main points are reviewed, further details can be found in the original reference.
$S=-\frac{1}{2\kappa}\int\sqrt{-g}d^{4}x\bigg[R+\frac{\gamma}{2}R^{2}-\frac{\alpha}{2}C_{\mu\nu\alpha\beta}C^{\mu\nu\alpha\beta}\bigg].$
The Stelle lagrangian is
$S=\int[-\frac{2}{\kappa^{2}}R+bR^{2}+\frac{a}{2}W_{\mu\nu}W^{\mu\nu}]\sqrt{-g}d^{4}x,$
with the first term is the Einstein-Hilbert lagrangian and
$W_{\mu\nu}W^{\mu\nu}=\frac{1}{2}R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}-R_{\mu\nu}R^{\mu\nu}+\frac{R^{2}}{6}.$
The metric in the synchronous gauge is
$ds^{2}=-dt^{2}+g_{ij}(x,t)dx^{i}dx^{j}.$ (5.60)
In these coordinates the Christofell symbols read as follows
$\Gamma_{\mu\nu}^{\rho}=\frac{1}{2}g^{\rho\sigma}(\partial_{\mu}g_{\sigma\nu}+\partial_{\nu}g_{\sigma\mu}-\partial_{\sigma}g_{\mu\nu})$ (5.61)
$\Gamma^{0}_{ij}=\Gamma^{0}_{ji}=\frac{1}{2}\dot{g}_{ij}\equiv-K_{ij},\quad\Gamma^{l}_{ij},\quad\Gamma^{l}_{0j}=\Gamma^{l}_{j0}=\frac{1}{2}g^{lm}\dot{g}_{mj}\equiv-K^{l}_{~j},$ (5.62)
where a dot represents a derivative with respect to $t$. The curvature terms in the synchronous gauge is decomposed as [1]
$R=R^{3}-2g^{ij}\dot{K}_{ij}-3K_{ij}K^{ij}+K^{2},$
$R_{\mu\nu}R^{\mu\nu}=\left[g^{ij}(\dot{K}_{ij}+K_{il}K^{l}_{~j})\right]^{2}-2(D_{i}K-D^{j}K_{ji})(D^{i}K-D^{l}K_{l}^{~i})$
$(~R^{3}_{ij}-\dot{K}_{ij}-2K_{il}K^{l}_{~j}+KK_{ij})g^{il}g^{jm}(~R^{3}_{lm}-\dot{K}_{lm}-2K_{lp}K^{p}_{~m}+KK_{lm}),$
$R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}=4(\dot{K}_{ij}+K_{il}K^{l}_{~j})g^{il}g^{jm}(\dot{K}_{lm}+K_{lp}K^{p}_{~m})$
$4(D^{j}K_{il}-D_{l}K^{j}_{~i})(D_{j}K^{il}-D^{l}K_{j}^{~i})$
$(~R^{3}_{ijlm}+K_{il}K_{jm}-K_{jl}K_{im})(~R^{\hskip-11.38092pt3\hskip 8.5359ptijlm}+K^{il}K^{jm}-K^{jl}K^{im}).$ (5.63)
Here  $R^{3}_{ijlm}$,  $R^{3}_{ij}$,  $R^{3}$ and $D_{i}$ are, respectively, the three-dimensional Riemann tensor, Ricci tensor, Ricci scalar and covariant derivative built with the three-dimensional metric $g_{ij}$.the inverse three-dimensional metric is represented as usual by $g^{ij}$, and the following quantity $K\equiv K_{i}^{~i}$ was also introduced in those formulas.
The quantity $\dot{K}_{ij}$ is time-reversal invariant and appears linearly, while $K_{ij}\equiv-\dot{g}_{ij}/2$ is odd but appears quadratically. The full lagrangian is time reversal.
As done in previous section by following the Ostrogradsky method, the variables $g_{ij}$, $\dot{g}_{ij}$ or, equivalently, of $g_{ij}$, $K_{ij}$ can be chosen. In this formalism $g_{ij}$ and $K_{ij}$ are independent canonical coordinates. The second choice will employed leading to the following conjugate momentum densities
$\pi^{ij}\equiv\frac{\partial L}{\partial\dot{g}_{ij}}-\frac{d}{dt}\frac{\partial L}{\partial\ddot{g}_{ij}},\quad P^{ij}\equiv\frac{\partial L}{\partial\dot{K}_{ij}}-\frac{d}{dt}\frac{\partial L}{\partial\ddot{K}_{ij}}$ (5.64)
Since $L_{s}$ is independent of $\ddot{K}_{ij}$,
$P^{ij}=\frac{\partial L}{\partial\dot{K}_{ij}}.$
This leads to
$P^{ij}=-2\sqrt{-g}[G^{ijlm}\dot{K}_{lm}+aK^{ij}K-ag^{ij}K_{lm}K^{lm}+\alpha R^{ij}$
$g^{ij}[(\frac{\alpha}{3}+2\gamma)(~R^{3}-3K_{lm}K^{lm}+K^{2})],$ (5.65)
where the quantity
$G^{ijlm}=a\,\frac{g^{il}g^{jm}+g^{im}g^{jl}}{2}+\left(4\gamma-\frac{\alpha}{3}\right)g^{ij}g^{lm},$
has been introduced. Note that, for non vanishing $\alpha$ and $\beta$ the quantity
$\Gamma_{pqij}\equiv\frac{1}{a}\frac{g_{pi}g_{qj}+g_{pj}g_{qi}}{2}-\frac{4\gamma-\alpha/3}{12\gamma\alpha}g_{pq}g_{ij}$ (5.66)
is the inverse of $G^{ijlm}$ , with the definition of inverse given by
$\Gamma_{pqij}G^{ijlm}=\frac{1}{2}(\delta_{p}^{l}\delta_{q}^{m}+\delta_{p}^{m}\delta_{q}^{l}).$
The crucial point is the following. The last expressions show that $\dot{K}_{ij}$ can be expressed in terms of $P^{ij}$, $g_{ij}$ and $K_{ij}$. Inserting this expression in $L_{s}$ the resulting functional $g_{ij}$, $K_{ij}$ and $P^{ij}$ with no dependence on $\pi^{ij}$. Therefore the Ostrogradsky Hamiltonian
$H=\pi^{ij}\dot{g}_{ij}+P^{ij}\dot{K}_{ij}-L_{s}.$ (5.67)
As the Stelle lagrangian $L_{s}$ does not depend on $\pi_{ij}$ it is clear that the only dependence is given by
$\pi^{ij}\dot{g}_{ij}=-2\pi^{ij}K_{ij}.$ (5.68)
This is a linear term in $\pi_{ij}$ and, if the path integral formalism is applied, the resulting path integral $Z(J)$ will involve a linear exponential in $\pi_{ij}$ whose integration will diverge. Therefore the path integral would not exist.
The above non existence argument does not hold if the quantization presented in previous sections is applied. This remarkable observation was done in [1]. The point is that $K_{ij}$ and its conjugate momentum $\pi_{ij}$ are ghost like variables, since they change the sign under time reversal, and in this quantization the Hamiltonian must be continued to imaginary values of these variables. This will convert the divergent exponential into an imaginary one, which is harmless.
To see this statement more explicitly, consider the quantization of Stelle gravity in the synchronous gauge. From (5.65) it is obtained that
$\dot{K}_{ab}=\Gamma_{abij}\bigg\{\frac{P^{ij}}{2\sqrt{-g}}+\alpha K^{ij}K-\alpha g^{ij}K_{lm}K^{lm}+a_{ij}$
$g^{ij}\left[\left(\frac{\alpha}{3}+2\gamma\right)(~R^{3}-3K_{lm}K^{lm}+K^{2}\right]\bigg\},$
By use of this formula replace $\dot{K}_{ij}$ in all the curvatures given in (5.63). The result will depend on $P^{ij}$, $g_{ij}$ and $K_{ij}$ but not on $\pi_{ij}$, as the last expression (5.65) does not depend on this quantity. The Hamiltoian is then an expression of the form
$H=\pi^{ij}\dot{g}_{ij}+P^{ij}\dot{K}_{ij}-L_{s}(g_{ij},P^{ij},K_{ij}).$
It is needed to analitically continue the classical Hamiltonian to imaginary values of the variables: $K_{ij}\to iK_{ij}$, $P^{ij}\to-iP^{ij}$ and remembering that $\dot{g}_{ij}=-2K_{ij}$
$H=-2i\pi^{ij}K_{ij}-iP^{ij}\dot{K}_{ij}-L_{s}(g_{ij},-iP^{ij},iK_{ij}).$
In the last expression it is understood that $K_{ij}$ should be written as (5.65). The path integral, in its euclidean version, where the metric is
$ds^{2}_{E}=d\tau^{2}+g_{ij}(x)dx^{i}dx^{j}.$
is given by
$Z_{e}(0)=\int D\pi^{ij}DK_{ij}DP^{ij}Dg_{ij}\exp\bigg\{-\int[i\pi^{ij}\dot{g}_{ij}+P^{ij}\dot{K}_{ij}-H(g_{ij},\pi_{ij},-iP^{ij},iK_{ij})]d\tau\bigg\}$
$\int D\pi^{ij}DK_{ij}DP^{ij}Dg_{ij}$
$\exp\bigg\{-\int[i\pi^{ij}\dot{g}_{ij}+iP^{ij}\dot{K}_{ij}-2i\pi^{ij}K_{ij}-iP^{ij}\dot{K}_{ij}-L_{s}(g_{ij},-iP^{ij},iK_{ij})]d\tau\bigg\}$
The integration over $\pi^{ij}$ leads to a Dirac delta that enforces $K_{ij}=-\frac{1}{2}g^{\prime}_{ij}$, where the sign indicates derivative with respect to the euclidean time. The path integral now is related to $D\dot{g}_{ij}$ due to this identification, but a change of variable to $Dg_{ij}$ can be done by multiplying the integration measure by $\prod_{x_{E}}\Delta\tau^{-6}$. The power of $6$ is due to the fact that $\pi^{ij}$, $g_{ij}$, $K_{ij}$ and $P^{ij}$ have a total of $6$ independent components at each spacetime point.
The integral over $P^{ij}$ is a Gaussian integral after setting $K_{ij}=-\frac{1}{2}g_{ij}^{\prime}$ , this follows after inspection of (5.63). The general formula
$\int d^{N}xe^{x^{T}Ax+bx}=\frac{\pi^{\frac{N}{2}}}{\sqrt{\det A}}e^{-\frac{1}{4}b^{T}Ab},$
can be employed to make this integral. Alternatively, they can be performed with a standard stationary point method. The stationary point satisfies now the equation
$iK_{ij}^{\prime}=\frac{\partial{\cal H}}{\partial P^{ij}},$
results in
$\frac{1}{2}g_{ij}^{\prime\prime}=\frac{\partial{\cal H}}{\partial(-iP^{ij})}.$
The last equation allows to express $P^{ij}$ in terms of $g_{ij}$, $\dot{g}_{ij}$ and $\ddot{g}_{ij}$ in classical physics but with $P^{ij}$ substituted by $iP_{ij}$. Regardless the method employed, the calculation is not difficult (it is just gaussian integration) the intermediate steps are a bit cumbersome. The final result is remarkable [1]-[2]. For finite initial and final euclidean times, after turning on a current $J^{ij}$, it is
$Z_{e}(J,\tau_{i},\tau_{f})=\int^{q(\tau_{f})=q_{f}}_{q(\tau_{i})=q_{i}}\,CDg_{ij}\,\exp\left(-\frac{S_{E}}{\hbar}+\int_{\tau_{i}}^{\tau_{f}}d\tau\int d^{3}x\,J^{ij}g_{ij}\right),$ (5.69)
where $S_{E}$ is the Euclidean version of the Stelle action
$S_{E}=\int_{\tau_{i}}^{\tau_{f}}d\tau\int d^{3}x\sqrt{g}\left(\frac{\alpha}{2}W_{E}^{2}+\gamma R_{E}^{2}+\frac{2}{\kappa^{2}}R_{E}\right).$ (5.70)
This result is important for the following reason. The theory that is constructed with this quantization is going to be unitary. However, it was not warranted that it will coincide with the euclidean Stelle gravity version. The Stelle theory is known to be renormalizable, and this property would be broken if another lagrangian was to be found. The fact that it coincides means that all the good properties of the Stelle model will be preserved, in particular renormalizability, while keeping unitarity [1]. The extra factor in the last path integral is
$C\sim\prod_{x_{E}^{\prime}}\frac{1}{\sqrt{g(x_{E})}}.$
This factor follows from the Gaussian integration formula directly. It is well known that it may introduce $\delta^{4}(0)$ singularities. However, if perturbation around flat spaces $\eta_{\mu\nu}$ are considered and dimensional regularization is employed, the well known Veltmann identities throw these terms to zero. The quantity $J^{ij}$ is the external “current” corresponding to $g_{ij}$ in the generating functional.
The above path integral is done in the synchronous gauge. Other gauges may be of interest, such as the De Donder one. The change to a general gauge leads to
$Z(J)=N\int\,{\cal D}g\,\left(\det\frac{\delta f}{\delta\xi}\right)\,\delta(f)\exp\left(-\frac{S_{E}}{\hbar}+\int d^{4}x_{E}\,J^{\mu\nu}g_{\mu\nu}\right),$ (5.71)
where the normalization factor is
$N=\frac{1}{\int\,{\cal D}g\,\left(\det\frac{\delta f}{\delta\xi}\right)\,\delta(f)\exp\left(-\frac{S_{E}}{\hbar}\right)}$
The determinant $\det\frac{\delta f}{\delta\xi}$ is the Faddeev-Popov determinant that usually appears when fixing a gauge in such theories.
After having the euclidean version of the model, the Lorentzian Green’s functions may be found in the following way. Replace the euclidean currents $J^{\mu\nu}(\tau)$ with a Lorentzian currents
$J^{kl}(\tau)\to J^{kl}(it)/\hbar,\qquad J^{k4}(\tau)\to iJ^{k0}(it)/\hbar,\qquad J^{44}(\tau)\to-J^{00}(it)/\hbar,$
and also make the replacement
$g_{kl}(\tau)\to g_{kl}(it),\qquad g_{k4}(\tau)\to-ig_{k0}(it),\qquad g_{44}(\tau)\to-g_{00}(it).$
These substitutions lead to
$Z(J)=N\int\,Dg\,\left(\det\frac{\delta f}{\delta\xi}\right)\,\delta(f)\exp\left(\frac{iS}{\hbar}+i\int d^{4}x\,J^{\mu\nu}g_{\mu\nu}/\hbar\right),$ (5.72)
where ${\cal Z}(J)$ is the generating functional of the Lorentzian Green’s functions. It is defined only as an analytic continuation of the Euclidean one.
The next task is to characterize the second quantization and the LSZ rules of the model.
6. The LSZ formula for Stelle gravity
6.1 The operator expression of the S matrix
By collecting all the information given above, the LSZ rules for the model may be written. Forgetting for one moment the ghost consider the rules that appear if one consider the standard Stelle theory. The analogy with (4.59) will lead to
$\hat{S}=:\exp\bigg\{\frac{2}{\kappa^{2}}\int h^{\mu\nu}_{in}(x)\square\bigg[\frac{1}{2}\bigg(1-\frac{\kappa^{2}\alpha}{4}\square\bigg)P^{2}-\bigg(1+\frac{\kappa^{2}(\alpha-3\beta)}{2}\square\bigg)P^{0-s}$
$\bigg(\frac{1}{2}+\frac{\kappa^{2}(4\beta-3\alpha)\square}{8}\bigg)P^{0-w}+\frac{\kappa^{4}}{8\lambda}\square[P^{1}-P^{0-w}]\bigg]_{\alpha\beta,\mu\nu}\frac{\delta}{\delta J_{\alpha\beta}(x)}d^{4}x.\bigg\}:Z(J)\bigg|_{J=0}.$ (6.73)
Note that the operator on the left corresponds to the Stelle lagrangian with a replacement $h^{\alpha\beta}\to\frac{\delta}{\delta J_{\alpha\beta}}$. In order to be consistent with the procedure of taking $P_{2}\to iP_{2}$ and $Q_{2}\to iQ_{2}$ one must write this lagrangian in terms of the variables $P_{1}$, $P_{2}$, $Q_{1}$ and $Q_{2}$. Then make the continuation. The results of the previous section show that the result is the free euclidean Stelle gravity. The expansion of the exponential in (6.73) leads to terms related to correlation functions of the following form
$\prod_{i=1}^{n}\hat{h}^{in}_{\gamma\delta}(x^{i})\hat{O}^{\gamma\delta\mu\nu}_{x^{i}}<\hat{T}h_{\mu\nu}(x^{1})...h_{\alpha\beta}(x^{n})>,$ (6.74)
where $\hat{O}$ is the operator defined in the exponential in (6.73). This operator removes external legs. The full element (6.74) is to be sandwiched between states which a definite momentum, considered as the standard ones.
However, the above result must indicate which are the free fields $h^{in}_{\mu\nu}$. In particular, how they act on the vacuum. In the following, a simpler gauge will be employed to indicate the structure constants of the creation and annihilation operator, assuming that these constants are gauge independent.
6.2 Second quantization of the model
Consider a small perturbation around the Minkowski metric
[…]
Inspired by the methods of [47]-[48], decompose the perturbation as
$\widetilde{h}_{\mu\nu}=\Psi_{\mu\nu}+\frac{\eta_{\mu\nu}}{2}\bigg(\phi+\frac{\Psi}{3}\bigg).$
The component $\Psi_{\mu\nu}$ is of spin two and the field $\phi$ is a scalar. The last expression for the perturbation assumes that the trace of the perturbation is not necessarily zero, and its value is encoded in $\phi$. The Stelle lagrangian, written in terms of the Weyl tensor. may be linealized and it becomes
$L=\frac{1}{2\kappa}\bigg[\frac{1}{4}\partial_{\mu}\Psi^{\alpha\beta}\partial^{\mu}\Psi_{\alpha\beta}-\frac{3}{2}\partial_{\mu}\phi\partial^{\mu}\phi+\partial_{\beta}\Psi^{\alpha\beta}\partial_{\alpha}\phi+\frac{1}{2}\partial_{\alpha}\phi\partial^{\alpha}\Psi\bigg]$
$\frac{\gamma}{2\kappa}\bigg[-\frac{9}{2}\left(\square\phi\right)^{2}+3\partial_{\alpha}\partial_{\beta}\Psi^{\alpha\beta}\square\phi-\frac{3}{2}\square\Psi\square\phi\bigg]+\frac{\alpha}{2\kappa}\left[\frac{1}{4}\square\Psi_{\mu\beta}\square\Psi^{\mu\beta}\right]+L_{int},$ (6.75)
up to interaction terms collected in $L_{int}$ that appear due to the higher order expansion of the curvature and the metric determinant. The equation of motion for $\phi$ is
$3\square\phi-\partial_{\alpha\beta}\Psi^{\alpha\beta}+\frac{1}{2}\square\Psi=\frac{\gamma}{2}\bigg[18\square^{2}\phi-6\square\partial_{\mu}\partial_{\nu}\Psi^{\mu\nu}+3\square^{2}\Psi\bigg],$ (6.76)
while for the spin two perturbation $\Psi_{\mu\nu}$ the equations become
$\square\Psi_{\mu\nu}+2\partial_{\mu}\partial_{\nu}\phi-\eta_{\mu\nu}\square\phi-6\gamma\square\partial_{\mu}\partial_{\nu}\phi+3\gamma\eta_{\mu\nu}\square^{2}\phi-\alpha\square^{2}\Psi_{\mu\nu}=0.$ (6.77)
The choice $\partial_{\mu}\Psi^{\mu\nu}=0$ leads to
$3\square\phi+\frac{1}{2}\square\Psi=\frac{\gamma}{2}\bigg[18\square^{2}\phi+3\square^{2}\Psi\bigg],$
$\square\partial_{\mu}\phi-3\gamma\square^{2}\partial_{\mu}\phi=0.$ (6.78)
This represents $\partial_{\mu}\phi$ as a massless mode plus a massive mode with mass
$m^{2}=\frac{1}{3\gamma}.$
This together with (6.78) means that
$\frac{1}{2}\square\partial_{\mu}\Psi=\frac{3\gamma}{2}\square^{2}\partial_{\mu}\Psi.$ (6.79)
The trace of the equation for $\Psi_{\mu\nu}$ leads to
$\square\Psi+2\square\phi-4\square\phi-6\gamma\square^{2}\phi+12\gamma\square^{2}\phi-\alpha\square^{2}\Psi=0.$
$\square\partial_{\mu}\Psi-\alpha\square^{2}\partial_{\mu}\Psi=0.$ (6.80)
Therefore $\partial_{\mu}\Psi$ contains only a massless mode, as the last two equations (6.79) and (6.80) are non compatible in other case. The equation (6.77) is now
$\square\Psi_{\mu\nu}+2\partial_{\mu}\partial_{\nu}\phi-6\gamma\square\partial_{\mu}\partial_{\nu}\phi-\alpha\square^{2}\Psi_{\mu\nu}=0.$
This, together with (6.78) shows that $\Psi_{\mu\nu}$ is composed by two modes, plus a third one dictated by the massless part of $\phi$. Then
$\square\Psi_{\mu\nu}+2\partial_{\mu}\partial_{\nu}\phi_{m=0}-\alpha\square^{2}\Psi_{\mu\nu}=0.$
Here the index $m=0$ indicates that only the massless mode enters in consideration. One may postulate that $\Psi_{\mu\nu}$ is composed by a mode $\psi_{2\mu\nu}$ with mass $M^{2}=\frac{1}{\alpha}$ plus a further contribution $\Psi_{a\mu\nu}$. This contribution will be such that
$\square\Psi_{a\mu\nu}-\alpha\square^{2}\Psi_{a\mu\nu}=-2\partial_{\mu}\partial_{\nu}\phi_{m=0}.$
As the left is massless, the expression on the right is also massles. Then both vanish, the left by the action of $\square$ and the right consequently. Therefore, the mode for $\phi$ is only massive.
Also (6.80) implies that the trace $\Psi$ is not zero only for the massless mode.
For the scalar mode $\phi$
$\phi(x)=\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}\omega_{3}(k)}[c(k)e^{i\omega_{2}(k)t-ik\cdot x}+c^{\ast}(k)e^{-i\omega_{3}(k)t+ik\cdot x}],$
where the new energy is given by $\omega_{3}(k)=\sqrt{k^{2}+m^{2}}$. For the spin two perturbation the mode expansion becomes
$\Psi_{\mu\nu}(k)=\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}\omega(k)}[a_{\mu\nu}(k)e^{i\omega(k)t-ik\cdot x}+a_{\mu\nu}^{\ast}(k)e^{-i\omega(k)t+ik\cdot x}]$
$\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}\omega(k)}[b_{\mu\nu}(k)e^{i\omega_{2}(k)t-ik\cdot x}+b^{\ast}_{\mu\nu}(k)e^{-i\omega_{2}(k)t+ik\cdot x}],$
where $\omega(k)=|k|$ and $\omega_{2}(k)=\sqrt{k^{2}+M^{2}}$. Also the gauge and the above found trace conditions lead to
$b_{\mu\nu}(k)\eta^{\mu\nu}=0,\qquad k^{i}a_{\mu i}(k)=\omega a_{\mu 0}(k).\qquad k^{i}b_{\mu i}(k)=\omega_{3}b_{\mu 0}(k).$
The quantum version of these fields are the operators
$\phi(x)=\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}\omega_{3}(k)}[c(k)e^{i\omega_{2}(k)t-ik\cdot x}+c^{\ast}(k)e^{-i\omega_{3}(k)t+ik\cdot x}],$
$\Psi_{\mu\nu}(k)=\sum_{p}\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}\omega(k)}[a_{\mu\nu}a_{p}(k)e^{i\omega(k)t-ik\cdot x}+a_{\mu\nu}^{\ast}a_{p}^{c}(k)e^{-i\omega(k)t+ik\cdot x}]$
$\sum_{p}\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}\omega_{2}(k)}[b_{\mu\nu}(k)b_{p}(k)e^{i\omega_{2}(k)t-ik\cdot x}+b^{\ast}_{\mu\nu}(k)b_{p}^{c}(k)e^{-i\omega_{2}(k)t+ik\cdot x}].$
The Gauss-Ostrogradsky method, applied to the lagrangian (6.75) leads to the following momentum
$P_{2}=\frac{\gamma}{2\kappa}\bigg[9\square\phi-3\partial_{\alpha}\partial_{\beta}\Psi^{\alpha\beta}+\frac{3}{2}\square\Psi\bigg]$
$P_{1}=\frac{1}{2\kappa}\bigg[3\dot{\phi}+\partial_{\beta}\Psi^{0\beta}+\frac{1}{2}\partial^{0}\Psi\bigg]-\frac{\gamma}{2\kappa}\bigg[9\square\dot{\phi}-3\partial_{\alpha}\partial_{\beta}\dot{\Psi}^{\alpha\beta}+\frac{3}{2}\square\dot{\Psi}\bigg]$
Some terms do not contribute due to the gauge fixing, by employing the Gupta-Bleuler method and taking mean values with the physical states. Taking into account that $Q_{2}=\dot{\phi}$ and that $\phi$ only contains a massive mode, it is impossible that
$P_{2},Q_{2}]=i,\qquad[Q_{1},Q_{2}]=0.$
as the first implies that $\phi,\dot{\phi}]\neq 0$ while the second is $\phi,\dot{\phi}]=0$. It looks like the third mode is gone [^2] This fact confused the author and it was the reason for writing a version that was erroneous.. This alone does not allows to follow the Gauss-Ostrogradsky procedure. However, the error is to apply the Gauss-Ostrogradsky method for this mode alone. The procedure should be applied to the full $h^{\mu\nu}$ momentum.
Therefore, one should keep going on. For $\Psi_{\alpha\beta}$ it follows that
$P_{2\alpha\beta}=-\frac{\alpha}{2\kappa}\frac{1}{2}\square\Psi_{\mu\beta}+\frac{\gamma}{2\kappa}\bigg[3\delta_{0\alpha}\delta_{\beta 0}\square\phi+\frac{3}{2}\eta_{\alpha\beta}\square\phi\bigg]$
$P_{1\alpha\beta}=\frac{1}{2\kappa}\bigg[-\frac{1}{2}\dot{\Psi}_{\alpha\beta}+\delta_{0(\beta}\partial_{\alpha)}\phi-\frac{1}{2}\eta_{\alpha\beta}\dot{\phi}\bigg]+\frac{\alpha}{2\kappa}\frac{1}{2}\square\dot{\Psi}_{\mu\beta}-\frac{\gamma}{2\kappa}\bigg[3\delta_{0\alpha}\delta_{\beta 0}\square\dot{\phi}+\frac{3}{2}\eta_{\alpha\beta}\square\dot{\phi}\bigg]$
Due to the equations of motions given above for $\Psi_{\mu\nu}$, it is seen the massless mode only contributes to $P_{1\alpha\beta}$ while the massive contributes to $P_{2\alpha\beta}$. It is clear that
$Q_{1\alpha\beta},Q_{2\mu\nu}]=[\Psi_{\alpha\beta},\dot{\Psi}_{\mu\nu}]=[\Psi^{1}_{\alpha\beta},\dot{\Psi}^{1}_{\mu\nu}]+[\Psi^{2}_{\alpha\beta},\dot{\Psi}^{2}_{\mu\nu}].$
$P_{1\alpha\beta},Q_{1\mu\nu}]=-\frac{1}{4\kappa}[\dot{\Psi}^{1}_{\alpha\beta},\Psi^{1}_{\mu\nu}],\qquad[P_{2\alpha\beta},Q_{2\mu\nu}]=-\frac{1}{4\kappa}[\Psi^{2}_{\alpha\beta},\dot{\Psi}^{2}_{\mu\nu}]$
$P_{1\alpha\beta},P_{2\mu\nu}]=0.$
All the above conclusions should be interpreted ”up to a gauge terms” and neglecting terms which will vanish when mean values are taken. The conditions for $Q_{1\alpha\beta},Q_{2\mu\nu}]=0$ and that $P_{1\alpha\beta},Q_{1\mu\nu}]$ is the identity up to gauge terms are
$a_{p}(k),a_{q}^{c}(k^{\prime})]=4\kappa\delta_{pq}\omega_{1}(k)\delta(k-k^{\prime}),\qquad[b_{p}(k),b_{q}^{c}(k^{\prime})]=-4\kappa\delta_{pq}\omega_{2}(k)\delta(k-k^{\prime}),$ (6.81)
The full contribution has to take into account $\phi$. However, this field will lead to terms proportional to $\eta_{\alpha\beta}\eta_{\gamma\delta}$ or $k_{\alpha}k_{\beta}k_{\gamma}k_{\delta}$ which are not proportional to the identity $I=\eta_{\alpha(\mu}\eta_{\beta)\nu}$. Therefore, the algebra of creation and annihilation operators of this mode will not make any difference, and it may be assumed that they are standard
$c_{p}(k),c_{q}^{c}(k^{\prime})]=4\kappa\delta_{pq}\omega_{3}(k)\delta(k-k^{\prime}).$ (6.82)
Only $b_{p}$ has a wrong sign. The prescription $P_{2}\to iP_{2}$ and $Q_{2}\to iQ_{2}$ has the effect analogous to the one explained in the section about the Pais-Uhlenbeck oscillator. It makes the algebra (6.81) standard.
6.3 Universality properties of the effective action
The above discussion was related to free graviton fields $h_{in\mu\nu}$, which are relevant for studying scattering as they enter in the LSZ formula. The full formula is
$\hat{S}=:\exp\bigg\{\frac{2}{\kappa^{2}}\int h^{\mu\nu}_{in}(x)\square\bigg[\frac{1}{2}\bigg(1-\frac{\kappa^{2}\alpha}{4}\square\bigg)P^{2}-\bigg(1+\frac{\kappa^{2}(\alpha-3\beta)}{2}\square\bigg)P^{0-s}$
$\bigg(\frac{1}{2}+\frac{\kappa^{2}(4\beta-3\alpha)\square}{8}\bigg)P^{0-w}+\frac{\kappa^{4}}{8\lambda}\square[P^{1}-P^{0-w}]\bigg]_{\alpha\beta,\mu\nu}\frac{\delta}{\delta J_{\alpha\beta}(x)}d^{4}x.\bigg\}:Z(J)\bigg|_{J=0},$ (6.83)
with
$h_{in}^{\mu\nu}=\sum_{p}\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}\omega(k)}[a_{\mu\nu}a_{p}(k)e^{i\omega(k)t-ik\cdot x}+a_{\mu\nu}^{\ast}a_{p}^{c}(k)e^{-i\omega(k)t+ik\cdot x}]$
$\sum_{p}\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}\omega_{2}(k)}[b_{\mu\nu}b_{p}(k)e^{i\omega_{2}(k)t-ik\cdot x}+b^{\ast}_{\mu\nu}b_{p}^{c}(k)e^{-i\omega_{2}(k)t+ik\cdot x}].$
$\sum_{p}\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}\omega_{3}(k)}[c_{\mu\nu}c_{p}(k)e^{i\omega_{3}(k)t-ik\cdot x}+c^{\ast}_{\mu\nu}c_{p}^{c}(k)e^{-i\omega_{3}(k)t+ik\cdot x}].$
The sum is over all the polarizations but the space of physical states $|\Psi>$ eliminates some of them. These states have to be consistent with the Stelle gauge. The algebra of the creation and annihilation is the standard, the wrong (6.81)-(6.82) is changed. The prescription requires to employ the euclidean version of the theory, and then go to the Lorenzian counterpart by continuing to imaginary times.
One may employ other gauges than the Stelle one. The advantage of this gauge is that several universality properties are known for the effective action. The partition function in this gauge is
$Z(\xi,\overline{\xi},K,L,J)=e^{iG(\xi,\overline{\xi},K,L,J)}=\int Dh^{\mu\nu}D\eta^{a}D\overline{\eta}^{b}\exp\{i\int d^{4}x[\overline{\xi}_{\alpha}\eta^{\alpha}+\xi_{\alpha}\overline{\eta}^{\alpha}+\kappa J_{\mu\nu}h^{\mu\nu}]\}$
$\exp\{i\int d^{4}x[L_{s}-\frac{\kappa^{2}}{2\lambda}F_{\tau}\square F^{\tau}+\overline{\eta}_{\tau}F_{\mu\nu}^{\tau}D^{\mu\nu}_{\alpha}\eta^{\alpha}+\kappa K_{\mu\nu}D^{\mu\nu}_{\alpha}\eta^{\alpha}+\kappa^{2}L_{\alpha}\partial_{\beta}\eta^{\alpha}\eta^{\beta}]\}.$ (6.84)
Here the current coupling is $J_{\mu\nu}h^{\mu\nu}$ and $J_{\mu\nu}$ may be identified as the energy momentum tensor $J_{\mu\nu}=T_{\mu\nu}$ for an external source. The quantity $\lambda$ is simple a constant parameter in the gauge fixing term, as discussed in the previous section. As in those theories, the sources $K_{\mu\nu}$ and $L_{\alpha}$ are coupled to the nilpotent part of the BRST invariance. The $\eta_{a}$ and $\overline{\eta}_{a}$ fields are the Faddeev-Popov ghosts and anti-ghost respectively. The standard transition amplitude is
$Z(\xi,\overline{\xi},J)=Z(\xi,\overline{\xi},J,K,L)|_{K=L=0}$
From the expression of the above functional, it follows that the gauge ghost action for the gravitational perturbation can be expressed as follows
$S_{gf}=\int d^{4}x[L_{s}-\frac{\kappa^{2}}{2\lambda}F_{\tau}\square F^{\tau}+\overline{\eta}_{\tau}F_{\mu\nu}^{\tau}D^{\mu\nu}_{\alpha}\eta^{\alpha}+\kappa K_{\mu\nu}D^{\mu\nu}_{\alpha}\eta^{\alpha}+\kappa^{2}L_{\alpha}\partial_{\beta}\eta^{\alpha}\eta^{\beta}].$
The operator related to the Stelle gauge choice is $F^{\tau}=\partial_{\nu}h^{\tau\nu}$. Furthermore $F^{\tau}_{\mu\nu}=\delta^{\tau}_{\mu}\partial_{\nu}$. The mass dimensions of the fields are directly seen from this action
$h^{\mu\nu}]=0,\qquad[\eta^{a}]=[\overline{\eta}^{a}]=1,\qquad[K_{\mu\nu}]=[L_{\alpha}]=3.$ (6.85)
Here, as before, the values for the ghost and anti-ghost are non uniquely defined. The choice above is however valid, and is the one that is usually employed. The operator $D_{\alpha}^{\beta\gamma}$ is defined by
$D_{\alpha}^{\mu\nu}\xi^{\alpha}=\partial^{\mu}\xi^{\nu}+\partial^{\nu}\xi^{\mu}-\eta^{\mu\nu}\partial_{\alpha}\xi^{\alpha}+h^{\alpha\mu}\partial_{\alpha}\xi^{\nu}+h^{\alpha\nu}\partial_{\alpha}\xi^{\mu}-\xi^{\alpha}\partial_{\alpha}h^{\mu\nu}-h^{\nu\nu}\partial_{\alpha}\xi^{\alpha}.$ (6.86)
As is well known, there is a deep relation between the functional $G$ given above and the effective action $\Gamma$. The relation is
$e^{iG(J,\xi,\overline{\xi},K,L)}=\int_{\text{C.T.D}}Dh^{\mu\nu}D\eta^{a}D\overline{\eta}^{a}\exp\{i\int\;d^{4}x\;[\Gamma(h,\eta,\overline{\eta},K,L)+$
$\overline{\xi}_{\alpha}\eta^{\alpha}+\xi_{\alpha}\overline{\eta}^{\alpha}+\kappa J_{\mu\nu}h^{\mu\nu}]\},$ (6.87)
where the notation C.T.D. stands for connected three diagrams. The physical meaning of the last formula is the following. If functional form of the effective action $\Gamma(h,\eta,\overline{\eta},K,L)$ has been obtained exactly, the functional $G(J,\xi,\overline{\xi},K,L)$ can be found by using the effective action in place of the original action and by calculating Feynman diagrams only at tree level, without loops.
An important point is that, with the gauge described in the above sections, the Stelle effective action has the universal form
$\Gamma_{n}=\int d^{4}x\bigg[A(h_{\mu\nu})+\frac{\delta S_{0}}{\delta h^{\mu\nu}}P^{\mu\nu}+(\kappa K_{\rho\sigma}-\overline{\eta}_{\tau}\overleftarrow{F}^{\tau}_{\rho\sigma})\bigg(\frac{\delta D^{\rho\sigma}_{\alpha}}{\delta h^{\mu\nu}}\eta^{\alpha}P^{\mu\nu}$
$\frac{\partial P^{\rho\sigma}}{\delta h^{\mu\nu}}D^{\mu\nu}_{\alpha}\eta^{\alpha}-D^{\rho\sigma}_{\alpha}(Q^{\alpha}_{\tau}\eta^{\tau})\bigg)$
$\kappa L_{\sigma}\bigg(\kappa Q^{\sigma}_{\tau}\partial_{\beta}\eta^{\tau}-\kappa\partial_{\beta}(Q^{\sigma}_{\tau}\eta^{\tau})-\kappa(\partial_{\tau}\eta^{\sigma})Q^{\tau}_{\beta}-\frac{\delta Q^{\sigma}_{\tau}}{\delta h^{\mu\nu}}\eta^{\tau}D_{\beta}^{\mu\nu}\bigg)\eta^{\beta}\bigg].$ (6.88)
This is exactly the formula (6.17) of the original reference [32]. Here the following quantity
$A(h_{\mu\nu})=\sqrt{-g}\bigg[-\frac{2}{\kappa^{2}+\delta\kappa^{2}}R+(\beta+\delta\beta)R^{2}-(\alpha+\delta\alpha)R_{\mu\nu}R^{\mu\nu}\bigg],$
has been introduced and $S_{0}$ is the classical lagrangian of Stelle gravity, in other words, the limit of $A(h_{\mu\nu})$ with the $\delta$ quantities taken to zero. The quantities $P^{\mu\nu}(h^{\alpha\beta})$ and $Q_{\nu}^{\mu}(h^{\alpha\beta})$ are functions of $h^{\alpha\beta}$ solely, and are not determined without a proper loop analysis for the effective action. Other gauges can be employed, this is allowed, however this universal form is not ensured. Its new expression should have to be worked out independently.
7. Final comments and possible future research programs
Before discussing future research programs, some comments about the prescription described between (6.73) and (6.74) are in order. Although the corresponding line of reasoning was described throughout the text, a further justification comes from Chapter 9 of the well known book [159]. In this chapter, these authors develop the path integral formalism starting with creation $a$ and annihilation operators $a^{\dagger}$ and take into account coherent states. The full path integral $Z(J)$, or the effective action $\Gamma$ depend on the mode, even after the prescription that reference [1]. This may raise a doubt about the consistency of the throwing away procedure of the third mode. More precisely, in Chapter 9, the book [159] develop formulas such as (9-87) or (9.990), which state that the free path integral is
$Z_{0}(J)=\exp[\frac{i}{2}\int d^{4}xd^{4}yJ_{0}(x)\Delta(x-y)J_{0}(y)],$
while the interacting one is
$Z(J)=\exp[-i\int d^{4}xV(-\frac{i\delta}{\delta J})]Z_{0}(J).$
Here $V$ is the interaction term. These authors also derive the scattering LSZ in formula (9-88). These formulas are of course taught in every QFT course. But the particularity of this book presentation is that these formulas are deduced starting with the second quantized theory. In the present context, the interaction $V$ may be interpreted as all the interaction vertices that appear when expanding the Stelle action. The quantity $\Delta(x-y)$ would correspond to the Stelle propagator. This quantity depends on the three mass scales $m_{i}^{2}$ (2.22) with $i=1,2,3$ for each graviton mode. The formalism of [1], which includes the analytic continuation $P_{2}\to iP_{2}$ and $Q_{2}\to iQ_{2}$ rediscovers, in the euclidean setting, the euclidean version of Stelle gravity.
Another issue is that the $S$ matrix presented in this work is unitary. The explanation is that the Fock states are taken as standard with positive norm, and the usually norm problems are attributed to bad creation and annihilation operators instead to bad norm states. In author´s opinion, the formalism presented here is equivalent to the pioneer ones, in particular [6], [7],[1], [2], [4]-[29]. The use of these states ensures that the $S$ matrix (4.43)
$S=\sum_{\text{all states}}|r,\text{out}><r,\text{in}|,$
is unitary. The evolution of the states is dictated by all the above procedure. The inconsistency is mapped now to the non bounded from below energies or observables mean values. However, the mapping $P_{2}\to iP_{2}$ and $Q_{2}\to iQ_{2}$ acting on bra’s $|>$ corrects the signs leading to this unfortunate situation.
The theory is renormalizable. The path integral can be given in terms of an effective action that can be renormalized with a finite number of counter terms [31] which, in some specific term, leads to the general form (6.88). One then may employ tree diagrams approximation for calculating correlations. The Slavnov-Taylor identities [37]-[38] ensure order by order gauge invariance. A review is [39], further details can be found in the textbooks [160]-[165].
Recently, a reference [36] appeared which apparently presents a consistent quantization of the Pais-Uhlenbeck oscillator. It is of interest to understand how the methods developed there apply for Quadratic Gravity. A work that seems to have started this program is [43], which is of clear interest.
The formulation of the present model in different gauges is of course relevant. In this context, the quantization of [49]-[50] can be of special interest, see also [51]-[52]. I believe that the disappearance of the third mode is gage independent, but the study of different gauges is interesting to check this.
There are several potential applications related to the present formalism. A first one is related to Bremsstrahlung in gravity theories, and classical limits of gravity. A reference that considered this issue in detail is [84]. The calculation of this effect requires to take the mean value of the waveform $h_{\mu\nu}>$. This calculation, for Quadratic Gravity, involves the prescription presented here, in particular the mapping of the oscillator structure constants $c_{1}$, $c_{2}$ to $d_{1}$$d_{2}$ or, equivalently, $P_{2}\to iP_{2}$ and $Q_{2}\to iQ_{2}$. There are several works related to this topic, restricted to GR. The work [83] for instance, considers a worldline formalism which may be useful to be generalized to the present context. Other references related to this topic are [53]-[58], [138]-[139]. The works [59]-[62] present interesting results about wave forms, bound states and Hawking radiation which are worthy to be reconsidered in the present context.
The use of coherent states in order to study the classical limit of gravity and scattering process, which was employed in [63]-[64] may be of usefl in this context. In special, regarding the discussion about the consistency of the oscillator prescription described in the above paragraphs.
Not less important, it may be of clear interest to reformulate the issues regarded to the eikonal limit of gravity, in view of the quantization presented here. An extensive but perhaps incomplete list of works dealing with such topics are [65]-[152]. There are computational codes [153]-[156] and new formalisms [157] that may facilitate the calculations required when studying these subjects.
The renormalization of the model presented here is related to perturbations along flat spaces. It may seem that applications related to curved backgrounds such as [122]-[123] are outside of the present formalism. However, already in 1973 it was found in reference [77] a connection between quantum Feynman diagrams and the perturbation series of the classical Schwarzschild background. This topic was explored further in [78]-[87], related work is [94]-[100]. These connections between classical and quantum aspects of gravity are worthy to be generalized to the present context.
In addition, the renormalization group of the model, studied with the present techniques may be of clear interest [171]-[173]. Further applications may be found in [174].
Also, it may be interesting to study the Lyra version of the Stelle model [175]-[181] and consider processes which involve the interactions between graviton, matter and the scale fields. Finally, there are apparent super renormalizable generalizations of the Stelle model [166]-[170]. A treatment of their ghost structure is a desirable analysis for the future.
Acknowledgements
The author is supported by CONICET, Argentina and by the Grant PICT 2020-02181. I am indebted to Diego Buccio, who pointed out an error in a previous version and whose help was very useful.